### v2

# Movable Antenna Enhanced Networked Full-Duplex Integrated Sensing and Communication System

Abstract
Integrated sensing and communication (ISAC) is envisioned as a key technology for future sixth-generation (6G) networks. Classical ISAC system considering monostatic and/or bistatic settings will inevitably degrade both communication and sensing performance due to the limited service coverage and easily blocked transmission paths. Besides, existing ISAC studies usually focus on downlink (DL) or uplink (UL) communication demands and unable to achieve the systematic DL and UL communication tasks. These challenges can be overcome by networked FD ISAC framework. Moreover, ISAC generally considers the trade-off between communication and sensing, unavoidably leading to a loss in communication performance. This shortcoming can be solved by the emerging movable antenna (MA) technology. In this paper, we utilize the MA to promote communication capability with guaranteed sensing performance via jointly designing beamforming, power allocation, receiving filters and MA configuration towards maximizing sum rate. The optimization problem is highly difficult due to the unique channel model deriving from the MA. To resolve this challenge, via leveraging the cutting-the-edge majorization-minimization (MM) method, we develop an efficient solution that optimizes all variables via convex optimization techniques. Extensive simulation results verify the effectiveness of our proposed algorithms and demonstrate the substantial performance promotion by deploying MA in the networked FD ISAC system.
Index Terms:
networked integrated sensing and communication (ISAC), full-duplex (FD), movable antenna (MA), majorization-minimization (MM) algorithm.
I Introduction
Recently, with the rapid increase in devices requiring precision sensing and efficient communication, integrated sensing and communication (ISAC) [1]--[2] technologies have attracted great interest from both industry and academia. In this context, ISAC is envisioned as a promising solution, which aiming to deploy both communication and sensing functionalities on one unified hardware platform which allows for the sharing of frequency spectrums, hardware facilities, and signal-processing units. Many recent advancements in the joint design of radar sensing and communication are extensively documented in [1]--[2] and their associated references.
However, the conventional fixed-position antennas (FPAs) at the base station (BS) and/or mobile user in the ISAC system generally fail to fully exploit the degrees of freedom (DoFs) in the continuous spatial domain, leading to performance loss in both sensing and communication tasks. A novel movable antenna (MA) technology [3]--[4] is expected to overcome the this limitation. The MA is considered as a viable approach for enhancing wireless network performance, linked to a radio frequency (RF) chain through a flexible cable. It can reconstruct channel conditions by flexibly adjusting its position in a given spatial area with the aid of a driver component or by other means. The various aspects of wireless systems with applying MA have been extensively studied recently, see [5]--[9] and their references.
I-A Related Works
[…]
I-B Motivations and Contributions
As shown previously, although extensive literature has explored the joint beamforming optimization for MA-aided ISAC systems, most studies have focused on monostatic and/or bistatic settings. These settings have limited service coverage and insufficient both sensing and communication performance in complex environments. In contrast, the networked ISAC is a promising solution to overcome the above shortcomings [1]--[2]. Moreover, the recent literature [10]--[18] only considers the DL communication in MA-aided ISAC networks. Note the waveform design in networked FD ISAC system remains largely unexplored, apart from the recent study [19]. More importantly, all the existing works [10]--[19] have not taken into account the integration of MA into the networked FD ISAC system. Therefore, we are motivated to study an MA-aided networked FD ISAC system to fully enhance its communication and sensing capabilities. Specifically, the contributions of this paper are elaborated as follows:
  • This paper considers the joint beamforming and position design in a networked FD ISAC system enabled by MA to promote simultaneous DL and uplink (UL) communication and target sensing. This novel framework can i) effectively enlarge the line-of-sight (LoS) region for sensing; ii) provide concurrent DL and UL communication transmission; iii) reconstruct the wireless channel condition by the MA technology.
  • We investigate the maximization of the sum-rate for both DL and UL users while guaranteeing radar sensing quality via jointly optimizing the BSs’ beamforming, UL users power allocation, receive filter and MAs’ positions. To the best of our knowledge, this problem has not been considered in the existing literature, e.g., [10]--[19].
  • The joint beamforming and MA configuration design to improve sum-rate of both DL and UL users is a highly non-convex problem due to the position coefficients of MA. Based on the cutting-the-edge majorization-minimization (MM) transformation [20], we successfully develop an effective algorithm that updates the position coefficients of MA by convex optimization techniques.
  • Extensive numerical results are provided to verify the effectiveness of our proposed solution. Meanwhile, experiment results validate that the deployment of MA can significantly benefit both DL and UL communication, promoting the sum-rate by up to 20.0% compared to the FPA benchmark, in the networked FD ISAC scheme.
II System Model and Problem Formulation
II-A System Model
Figure 1: A MA-aided networked FD ISAC system. (a): System model (b): Illustrations of the MAs at the BS and users, respectively.
As shown in Fig. 1, we consider an MA-enhanced networked FD ISAC system that comprises $M_{t}$ transmit BSs (TBSs), $M_{r}$ receive BSs (RBSs), $K_{d}$ single MA DL mobile users, $K_{u}$ single MA UL mobile users, one point-like sensing target, and $K_{t}$ clutter sources. Moreover, TBSs and RBSs are respectively equipped with $N_{t}$ and $N_{r}$ MAs. Let $\mathcal{M}_{t}\triangleq\{1,\cdots,M_{t}\}$, $\mathcal{M}_{r}\triangleq\{1,\cdots,M_{r}\}$, $\mathcal{N}_{t}\triangleq\{1,\cdots,N_{t}\}$, $\mathcal{N}_{r}\triangleq\{1,\cdots,N_{r}\}$, $\mathcal{K}_{d}\triangleq\{1,\cdots,K_{d}\}$, $\mathcal{K}_{u}\triangleq\{1,\cdots,K_{u}\}$ and $\mathcal{K}_{t}\triangleq\{1,\cdots,K_{t}\}$ denoted the sets of TBSs, RBSs, TBSs’ MAs, RBSs’ MAs, DL users, UL users and clutter sources, respectively.
In this network, TBSs, with aid of MA, service DL users and send probing waveform to detect the target, and RBSs, with the assistance of MA, receive the information from UL users and echo signal from both the target and clutters. Besides, we assumed that all TBSs and RBSs achieve perfect synchronization and joint signal processing facilitated by the central controller [21].
Moreover, the MAs are connected to RF chains by flexible cables, and can move within local region in real time [3], [4]. The position of the $n$-th MA of $m$-th BS is given by $\mathbf{t}^{\iota}_{m,n}=[x^{\iota}_{m,n},y^{\iota}_{m,n}]^{T}\in\mathcal{C}$, where $\iota\in\{0,1\}$, index 00 represents $\forall n\in\mathcal{N}_{t}$ and $\forall m\in\mathcal{M}_{t}$, and index $1$ denotes $\forall n\in\mathcal{N}_{r}$ and $\forall m\in\mathcal{M}_{r}$. The positions of single MA at the $k$-th DL user and $l$-th UL user are respectively written as $\mathbf{r}_{0,k}=[x^{d}_{k},y^{d}_{k}]^{T}\in\mathcal{C}$ and $\mathbf{r}_{1,l}=[x^{u}_{l},y^{u}_{l}]^{T}\in\mathcal{C}$, where $\mathcal{C}$ denotes the given two-dimensional (2D) moving square region with a size of $A\times A$ for both the BSs’ and users’ MAs.
Since the MA’s moving area is usually much smaller than the signal propagation distance, the links between the BSs/users and the users are assumed to follow the far-field wireless channel model [3], [4]. Therefore, changing the positions of MAs has no effect on the angle of departure (AoD), the angle of arrival (AoA), and the amplitude of the complex path coefficient, while only the phase of the complex path coefficient varies with changes in MAs’ position.
Let $L^{t,\iota}_{m,k}$ and $L^{r,\iota}_{m,k}$ denote the total number of transmit and receive channel paths at the $m$-th BS from $k$-th user, respectively. And then, $L^{t,2}_{k,l}$ and $L^{r,2}_{k,l}$ are respectively denoted as the total number of transmit and receive channel paths between DL user $k$ and UL user $l$. The elevation and azimuth AoDs of $i$-th transmit path between BS $m$ and user $k$ are respectively denoted as $\theta^{t,\iota}_{m,k,i}\in[-\frac{\pi}{2},\frac{\pi}{2}]$ and $\phi^{t,\iota}_{m,k,i}\in[-\frac{\pi}{2},\frac{\pi}{2}]$. Similarly, the elevation and azimuth AoAs of $i$-th receive path between BS $m$ and user $k$ are represented as $\theta^{r,\iota}_{m,k,i}\in[-\frac{\pi}{2},\frac{\pi}{2}]$ and $\phi^{r,\iota}_{m,k,i}\in[-\frac{\pi}{2},\frac{\pi}{2}]$, respectively. Besides, denote the elevation and azimuth AoDs of $i$-th transmit path between UL user $l$ and DL user $k$ as $\theta^{t,2}_{l,k,i}\in[-\frac{\pi}{2},\frac{\pi}{2}]$ and $\phi^{t,2}_{l,k,i}\in[-\frac{\pi}{2},\frac{\pi}{2}]$, respectively, and denote separately the elevation and azimuth AoAs of $i$-th receive path between UL user $l$ and DL user $k$ as $\theta^{r,2}_{l,k,i}\in[-\frac{\pi}{2},\frac{\pi}{2}]$ and $\phi^{r,2}_{l,k,i}\in[-\frac{\pi}{2},\frac{\pi}{2}]$. Next, the difference in signal propagation for the $i$-th transmit path of $k$-th user between the position of $n$-th MA of $m$-th BS and its reference point, i.e., $\mathbf{t}^{\iota}_{m,0}=[0,0]^{T}$, can be given as
$\rho^{t,\iota}_{m,k,n,i}(\mathbf{t}^{\iota}_{m,n})=({\mathbf{t}_{m,n}^{{\iota}}})^{T}\mathbf{a}^{t,\iota}_{m,k,i}$ (1)
$=x^{\iota}_{m,n}\cos(\theta^{t,\iota}_{m,k,i})\sin(\phi^{t,\iota}_{m,k,i})+y^{\iota}_{m,n}\sin(\theta^{t,\iota}_{m,k,i}),$
and its phase difference can be obtained as $2\pi\rho^{t,\iota}_{m,k,n,i}(\mathbf{t}^{\iota}_{m,n})/\lambda$, where $\lambda$ denotes the carrier wavelength. Thus, the field-response vector (FRV) of the transmit channel paths between $k$-th user and $m$-th BS’ $n$-th MA is given by
$\mathbf{h}^{\iota}_{m,k,n}\!\!=\!\![e^{j\frac{2\pi}{\lambda}\rho^{t,\iota}_{m,k,n,1}},e^{j\frac{2\pi}{\lambda}\rho^{t,\iota}_{m,k,n,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{t,\iota}_{m,k,n,L^{t,\iota}_{m,k}}}]^{T},$ (2)
and then the field response matrixs (FRMs) of all $N_{t}$ MAs for $m$-th TBS and all $N_{r}$ MAs for $p$-th TBS can be written as
$\mathbf{H}^{0}_{m,k}(\mathbf{t}^{0}_{m})=[\mathbf{h}^{0}_{m,k,1},\mathbf{h}^{0}_{m,k,2},\cdots,\mathbf{h}^{0}_{m,k,N_{t}}]\in\mathbb{C}^{L^{t,0}_{m,k}\times N_{t}},$ (3)
$\mathbf{H}^{1}_{p,l}({\mathbf{t}^{1}_{p}})=[\mathbf{h}^{1}_{p,l,1},\mathbf{h}^{1}_{p,l,2},\cdots,\mathbf{h}^{1}_{p,l,N_{r}}]\in\mathbb{C}^{L^{t,1}_{p,l}\times N_{r}},$ (4)
respectively, where $\mathbf{t}^{0}_{m}\triangleq\{\mathbf{t}^{0}_{m,n}\}$ and $\mathbf{t}^{1}_{p}\triangleq\{\mathbf{t}^{1}_{p,i}\}$. Moreover, the FRVs of the receive channel paths between $k$-th DL user and $m$-th TBS and between $l$-th UL user and $p$-th RBS are respectively defined as
$\mathbf{h}^{0,0}_{m,k}(\mathbf{r}_{0,k})=[e^{j\frac{2\pi}{\lambda}\rho^{r,0}_{m,k,1}},e^{j\frac{2\pi}{\lambda}\rho^{r,0}_{m,k,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{r,0}_{m,k,L^{r,0}_{m,k}}}],$ (5)
$\mathbf{h}^{0,1}_{p,l}(\mathbf{r}_{1,l})=[e^{j\frac{2\pi}{\lambda}\rho^{r,1}_{p,l,1}},e^{j\frac{2\pi}{\lambda}\rho^{r,1}_{p,l,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{r,1}_{p,l,L^{r,1}_{p,l}}}],$ (6)
where $\rho^{r,0}_{m,k,i}(\mathbf{r}_{0,k})\!\triangleq\!\mathbf{r}_{0,k}^{T}\mathbf{a}^{r,0}_{m,k,i}\!=\!x^{d}_{k}\cos(\theta^{r,0}_{m,k,i})\sin(\phi^{r,0}_{m,k,i})+y^{d}_{k}\sin(\theta^{r,0}_{m,k,i})$ and $\rho^{r,1}_{p,l,j}(\mathbf{r}_{1,l})\!\!\triangleq\!\!\mathbf{r}_{1,l}^{T}\mathbf{a}^{r,1}_{p,l,j}\!=\!y^{u}_{l}\sin(\theta^{r,1}_{p,l,j})+x^{u}_{l}\cos(\theta^{r,1}_{p,l,j})\sin(\phi^{r,1}_{p,l,j})$.
Besides, the FRVs of the transmit and receive channel paths between $k$-th DL user and $l$-th UL user are respectively expressed as
$\mathbf{h}^{2}_{k,l}(\mathbf{r}_{1,l})=[e^{j\frac{2\pi}{\lambda}\rho^{t,2}_{k,l,1}},e^{j\frac{2\pi}{\lambda}\rho^{t,2}_{k,l,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{t,2}_{k,l,L^{t,2}_{k,l}}}],$ (7)
$\mathbf{h}^{3}_{k,l}(\mathbf{r}_{0,k})=[e^{j\frac{2\pi}{\lambda}\rho^{r,2}_{k,l,1}},e^{j\frac{2\pi}{\lambda}\rho^{r,2}_{k,l,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{r,2}_{k,l,L^{r,2}_{k,l}}}],$ (8)
where $\rho^{t,2}_{k,l,i}(\mathbf{r}_{1,l})\triangleq\mathbf{r}_{1,l}^{T}\mathbf{a}^{t,2}_{k,l,i}=x^{u}_{l}\cos(\theta^{t,2}_{k,l,i})\sin(\phi^{t,2}_{k,l,i})+y^{u}_{l}\sin(\theta^{t,2}_{k,l,i})$ and $\rho^{r,2}_{k,l,i}(\mathbf{r}_{0,k})\triangleq\mathbf{r}_{0,k}^{T}\mathbf{a}^{r,2}_{k,l,i}=y^{d}_{k}\sin(\theta^{r,2}_{k,l,i})+x^{d}_{k}\cos(\theta^{r,2}_{k,l,i})\sin(\phi^{r,2}_{k,l,i})$. Furthermore, the matrixes $\mathbf{\Sigma}^{0}_{m,k}$, $\mathbf{\Sigma}^{1}_{p,l}$, and $\mathbf{\Sigma}^{2}_{k,l}$ are respectively denoted as the responses of all transmit and receive paths from the $m$-th TBS to the $k$-th DL user, from the $p$-th RBS to the $l$-th DL user and from the $l$-th UL user to the $k$-th DL user. Therefore, the channel vectors between the $m$-th TBS and the $k$-th DL user and between the $p$-th RBS and the $l$-th UL user are given by
$\mathbf{h}_{d,m,k}(\mathbf{t}^{0}_{m},\mathbf{r}_{0,k})^{H}=\mathbf{h}^{0,0}_{m,k}(\mathbf{r}_{0,k})^{H}\mathbf{\Sigma}^{0}_{m,k}\mathbf{H}^{0}_{m,k}(\mathbf{t}^{0}_{m})\in\mathbb{C}^{1\times N_{t}},$ (9)
$\mathbf{h}_{u,p,l}(\mathbf{t}^{1}_{p},\mathbf{r}_{0,l})^{H}=\mathbf{h}^{0,1}_{p,l}(\mathbf{r}_{0,l})^{H}\mathbf{\Sigma}^{1}_{p,l}\mathbf{H}^{1}_{p,l}(\mathbf{t}^{1}_{p})\in\mathbb{C}^{1\times N_{r}},$ (10)
And, the channel between the $k$-th DL user and the $l$-th UL user can be expressed by
${h}_{du,k,l}(\mathbf{r}_{0,k},\mathbf{r}_{1,l})^{*}=(\mathbf{h}^{3}_{k,l}(\mathbf{r}_{0,k}))^{H}\mathbf{\Sigma}^{2}_{k,l}\mathbf{h}^{2}_{k,l}(\mathbf{r}_{1,l}).$ (11)
Besides the steering vector of the $m$-th TBS towards the sensing target and/or clutters is represented by
$\mathbf{g}^{t}_{m,j}=[e^{j\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{m,1})^{T}\mathbf{a}^{4}_{m,j}},e^{j\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{m,2})^{T}\mathbf{a}^{4}_{m,j}},$ (12)
$\cdots,e^{j\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{m,N_{t}})^{T}\mathbf{a}^{4}_{m,j}}],\forall j\in\mathcal{J}\triangleq\{0\}\cup\mathcal{K}_{t},$
where index 00 represent the sensing target, $\mathbf{a}^{4}_{m,j}\triangleq[\cos(\theta^{t,4}_{m,j})\sin(\phi^{t,4}_{m,j});\sin(\theta^{t,4}_{m,j})]^{T}$, and $\theta^{t,4}_{m,j}$ and $\phi^{t,4}_{m,j}$ denote the elevation and azimuth AoDs between TBS $m$ and target/clutters. Similarly, the steering vector of the $p$-th RBS towards the sensing target/clutters also can be written as
$\mathbf{g}^{r}_{p,j}=[e^{j\frac{2\pi}{\lambda}(\mathbf{t}^{1}_{p,1})^{T}\mathbf{a}^{5}_{p,j}},e^{j\frac{2\pi}{\lambda}(\mathbf{t}^{1}_{p,2})^{T}\mathbf{a}^{5}_{p,j}},$ (13)
$\cdots,e^{j\frac{2\pi}{\lambda}(\mathbf{t}^{1}_{p,N_{r}})^{T}\mathbf{a}^{5}_{p,j}}],\forall j\in\mathcal{J},$
where $\mathbf{a}^{5}_{p,j}\triangleq[\cos(\theta^{r,4}_{p,j})\sin(\phi^{r,4}_{p,j});\sin(\theta^{r,4}_{p,j})]^{T}$, and $\theta^{r,4}_{p,j}$ and $\phi^{r,4}_{p,j}$ denote the elevation and azimuth AoAs between RBS $p$ and target/clutters. Therefore, the $m$-th TBS-target/clutters and the $p$-th RBS-target/clutters channels are respectively given as
$\mathbf{g}_{t,m,j}=\beta^{t}_{m,j}\mathbf{g}^{t}_{m,j},\ \mathbf{g}_{r,p,j}=\beta^{r}_{p,j}\mathbf{g}^{r}_{p,j},$ (14)
where $\{\beta^{t}_{m,j}\}$ and $\{\beta^{r}_{p,j}\}$ are complex fading coefficients and are assumed to be known.
Denote $\mathbf{s}\triangleq[s_{1},\cdots,s_{k},\cdots,s_{K_{d}}]^{T}\in\mathbb{C}^{K_{d}\times 1}$ and $\mathbf{s}_{r}\in\mathbb{C}^{N_{t}\times 1}$ as the communication symbol transmitted to DL users and radar probing signal from TBS, respectively. For simplicity, it is assumed that $\mathbb{E}\{\mathbf{s}\mathbf{s}^{H}\}=\mathbf{I}$, $\mathbb{E}\{\mathbf{s}_{r}\mathbf{s}_{r}^{H}\}=\mathbf{I}$ and $\mathbb{E}\{\mathbf{s}\mathbf{s}_{r}^{H}\}=\mathbf{0}$. The transmit signal of the $m$-th TBS can be written as
$\mathbf{x}_{m}={\sum}_{k=1}^{K_{d}}\mathbf{w}_{m,k}s_{k}+\mathbf{W}^{r}_{m}\mathbf{s}_{r},\forall m\in\mathcal{M}_{t},$ (15)
where $\mathbf{w}_{m,k}\in\mathbb{C}^{N_{t}\times 1}$ and $\mathbf{W}^{r}_{m}\in\mathbb{C}^{N_{t}\times N_{t}}$ represent the transmit beamformer of the $m$-th TBS for the $k$-th DL user’s communication signal and the probing signal, respectively.
The $l$-th UL user’s uplink signal can be written as
$x^{u}_{l}=\sqrt{q_{l}}s_{l}^{u},\ \forall l\in\mathcal{K}_{u},$ (16)
where $s_{l}^{u}$ and $q_{l}$ denote the information symbol and transmission power of $l$-th UL user, respectively, and $s_{l}^{u}$ are mutually uncorrelated and each has zero mean and unit variance.
The received signal at the DL user $k$ can be written as
$y_{D,k}=\underbrace{{\sum}_{m=1}^{M_{t}}\mathbf{h}_{d,m,k}^{H}\mathbf{x}_{m}}\limits_{\textrm{DL signals}}+\underbrace{{\sum}_{l=1}^{K_{u}}{h}_{du,k,l}x^{u}_{l}}\limits_{\textrm{UL signals}}+n_{d,k}$ (17)
$={\sum}_{i=1}^{K_{d}}\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{i}s_{i}+\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{W}}^{r}\mathbf{s}^{r}+{\sum}_{l=1}^{K_{u}}{h}_{du,k,l}x^{u}_{l}+n_{d,k},$
where $\hat{\mathbf{h}}_{d,k}\triangleq[\mathbf{h}_{d,1,k}^{T},\mathbf{h}_{d,2,k}^{T},\cdots,\mathbf{h}_{d,M_{t},k}^{T}]^{T}\in\mathbb{C}^{M_{t}N_{t}\times 1}$, $\hat{\mathbf{w}}_{k}\triangleq[\mathbf{w}_{1,k}^{T},\mathbf{w}_{2,k}^{T},\cdots,\mathbf{w}_{M_{t},k}^{T}]^{T}\in\mathbb{C}^{M_{t}N_{t}\times 1}$, $\bar{\mathbf{W}}_{r}\triangleq[(\mathbf{W}^{r}_{1})^{T},(\mathbf{W}^{r}_{2})^{T},\cdots,(\mathbf{W}^{r}_{M_{t}})^{T}]^{T}\in\mathbb{C}^{M_{t}N_{t}\times N_{t}}$, $\hat{\mathbf{W}}_{r}\triangleq\mathrm{blkdiag}(\bar{\mathbf{W}}_{r})$ and $n_{d,k}\sim\mathcal{CN}(0,\sigma_{d,k}^{2})$ is denoted as the complex additive white Gaussian noise (AWGN) at the $k$-th DL user.
The received signal at $p$-th RBS can be expressed as
$\mathbf{y}_{r,p}$ $=\underbrace{{\sum}_{j=0}^{J}\alpha_{t,j}\mathbf{g}_{r,p,j}({\sum}_{m=1}^{M_{t}}\mathbf{g}_{t,m,j}^{H}\mathbf{x}_{m})}\limits_{\textrm{sensing echoes}}$ (18)
$+\underbrace{{\sum}_{l=1}^{K_{u}}\mathbf{h}_{u,p,l}x^{u}_{l}}\limits_{\textrm{UL signals}}+\mathbf{n}_{r,p},$
where $\mathbf{n}_{r,p}\sim\mathcal{CN}(0,\sigma_{r}^{2}\mathbf{I})$ is the AWGN at the $p$-th RBS, $\alpha_{t,j}$ represents the target radar cross section (RCS) between TBS and RBS and $\alpha_{t,j}\sim\mathcal{CN}(0,\sigma_{t,j}^{2})$.
To obtain the information of UL users and target separately, RBS $p$ will adopt $K_{u}+1$ linear filters $\mathbf{u}_{p,i}\in\mathbb{C}^{M_{r}\times 1}$, $\forall i\in\mathcal{J}_{u}\triangleq\{0\}\cup\mathcal{K}_{u}$, to post-process the received signal, where $\mathbf{u}_{p,0}$ corresponds to the radar sensing filter. Therefore, the output information of $i$-th filter at $p$-th RBS can be written as
$\hat{y}_{r,p,i}=\mathbf{u}_{p,i}^{H}\mathbf{y}_{r,p},\forall i\in\mathcal{J}_{u},$ (19)
and then the information from all $M_{r}$ RBS collected at the central controller is given by
$\hat{y}^{u}_{r,i}={\sum}_{p=1}^{M_{r}}\hat{y}_{r,p,i}=\sum_{l=1}^{K_{u}}\hat{\mathbf{u}}_{i}^{H}\hat{\mathbf{h}}_{u,l}x^{u}_{l}$ (20)
$+{\sum}_{j=0}^{K_{t}}\alpha_{j}\hat{\mathbf{u}}_{i}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}({\sum}_{k=1}^{K_{d}}\hat{\mathbf{w}}_{k}s_{k}+\hat{\mathbf{W}}_{r}\hat{\mathbf{s}}^{r})+\hat{\mathbf{u}}_{i}^{H}\mathbf{n}_{r},$
where
$\hat{\mathbf{u}}_{i}\triangleq[\mathbf{u}_{1,i}^{T},\mathbf{u}_{2,i}^{T},\cdots,\mathbf{u}_{M_{r},i}^{T}]^{T}\in\mathbb{C}^{M_{r}N_{r}\times 1},$ (21)
$\hat{\mathbf{h}}_{u,l}\triangleq[\mathbf{h}_{u,1,l}^{T},\mathbf{h}_{u,2,l}^{T},\cdots,\mathbf{h}_{u,M_{r},l}^{T}]^{T}\in\mathbb{C}^{M_{r}N_{r}\times 1},$
$\hat{\mathbf{g}}_{t,j}\triangleq[\mathbf{g}_{t,1,j}^{T},\mathbf{g}_{t,2,j}^{T},\cdots,\mathbf{g}_{t,M_{t},j}^{T}]^{T}\in\mathbb{C}^{M_{t}N_{t}\times 1},$
$\hat{\mathbf{g}}_{r,j}\triangleq[\mathbf{g}_{r,1,j}^{T},\mathbf{g}_{r,2,j}^{T},\cdots,\mathbf{g}_{r,M_{r},j}^{T}]^{T}\in\mathbb{C}^{M_{r}N_{r}\times 1},$
$\mathbf{n}_{r}\triangleq[\mathbf{n}_{r,1}^{T},\mathbf{n}_{r,2}^{T},\cdots,\mathbf{n}_{r,M_{r}}^{T}]^{T}\in\mathbb{C}^{M_{r}N_{r}\times 1}.$
The achievable rate of $k$-th DL user and $l$-th UL user can be obtained as
$\textrm{R}_{d,k}=\textrm{log}(1+\textrm{SINR}_{d,k}),\forall k\in\mathcal{K}_{d},$ (22)
$\textrm{R}_{u,l}=\textrm{log}(1+\textrm{SINR}_{u,l}),\forall l\in\mathcal{K}_{u},$ (23)
respectively, where
$\textrm{SINR}_{d,k}=$ (24)
$\frac{|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{k}|^{2}}{{\sum}_{i\neq k}^{K_{d}}|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{i}|^{2}+\|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{W}}^{r}\|^{2}_{F}+{\sum}_{l=1}^{K_{u}}q_{l}|{h}_{du,k,l}|^{2}+\sigma_{d,k}^{2}},$
and the SINR for each UL user can be written as (25), as shown on the top of this page.
$\textrm{SINR}_{u,l}=\frac{q_{l}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,l}|^{2}}{\sum_{i\neq l}^{K_{u}}q_{i}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,i}|^{2}+\sum_{j=0}^{K_{t}}\sum_{k=1}^{K_{d}}\sigma_{t,j}^{2}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}|^{2}+\sum_{j=0}^{K_{t}}\sigma_{t,j}^{2}\|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{F}+\sigma_{r}^{2}\|\hat{\mathbf{u}}_{l}^{H}\|^{2}_{2}}$ (25)
The radar SINR [22] for sensing can be derived as (26), as shown on the top of next page.
$\textrm{SINR}_{t}=\frac{\sum_{k=1}^{K_{d}}\sigma_{t,0}^{2}|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,0}\hat{\mathbf{g}}_{t,0}^{H}\hat{\mathbf{w}}_{k}|^{2}+\sigma_{t,0}^{2}\|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,0}\hat{\mathbf{g}}_{t,0}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{F}}{\sum_{i=1}^{K_{u}}q_{i}|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{h}}_{u,i}|^{2}+\sum_{j=1}^{K_{t}}\sum_{k=1}^{K_{d}}\sigma_{t,j}^{2}|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}|^{2}+\sum_{j=1}^{K_{t}}\sigma_{t,j}^{2}\|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{F}+\sigma_{r}^{2}\|\hat{\mathbf{u}}_{0}^{H}\|^{2}_{2}}$ (26)
II-B Problem Formulation
Our goal is to maximize the sum-rate of all DL and UL users via jointly optimizing the transmit beamformer $\{\mathbf{w}_{m,k},\mathbf{W}^{r}_{m}\}$, the linear filters $\{\mathbf{u}_{p,i}\}$, the DL users’ uplink transmit power $\{q_{l}\}$, the positions of all MAs $\{\mathbf{t}^{0}_{m,n},\mathbf{t}^{1}_{p,j},\mathbf{r}_{0,k},\mathbf{r}_{1,l}\}$. The sum-rate optimization problem can be formulated as
[…]
s.t. $\textrm{SINR}_{t}\geq\Gamma_{r},$ (27b)
${\sum}_{k=1}^{K_{d}}\|\mathbf{w}_{m,k}\|_{2}^{2}+\|\mathbf{W}^{r}_{m}\|_{F}^{2}\leq\textrm{P}^{BS}_{m},\forall m\in\mathcal{M}_{t},$ (27c)
$0\leq q_{l}\leq P_{u,l},\forall l\in\mathcal{K}_{u},$ (27d)
$\|\mathbf{t}^{0}_{m,n}\!-\!\mathbf{t}^{0}_{m,i}\|_{2}\!\geq\!D_{t},\forall n,i\!\in\!\mathcal{N}_{t},n\!\neq\!i,\forall m\!\in\!\mathcal{M}_{t},$ (27e)
$\|\mathbf{t}^{1}_{p,j}\!-\!\mathbf{t}^{1}_{p,s}\|_{2}\!\geq\!D_{t},\forall j,s\in\mathcal{N}_{r},j\neq s,\forall p\in\mathcal{M}_{r},$ (27f)
$\mathbf{t}^{0}_{m,n}\in\mathcal{C},\ \mathbf{t}^{1}_{p,j}\in\mathcal{C},\ \mathbf{r}_{0,k}\in\mathcal{C},\ \mathbf{r}_{1,l}\in\mathcal{C},$ (27g)
where $\Gamma_{r}$ denotes the minimum target sensing SINR threshold, $P^{BS}_{m}$ and $P_{u,l}$ represent the maximum transmission power of the $m$-th TBS and the $l$-th UL user, respectively, and $D_{t}$ is the minimum distance between MAs of BSs to avoid the coupling effect. Obviously, due to the non-convex objective function and constraints, the optimization problem (P0) is highly challenging.
III Algorithm Design
III-A Problem Reformulation
To make the problem (P0) more tractable, we first employ the weighted minimum mean squared error (WMMSE) method [23] to equivalently transform the objective functions $\textrm{R}_{d,k}$ and $\textrm{R}_{u,l}$ into (29) and (30) , respectively, at the top of next page, where $\{\omega_{d,k}\}$, $\{\beta_{d,k}\}$, $\{\omega_{u,l}\}$ and $\{\beta_{u,l}\}$ are the introduced auxiliary variables. Therefore, the original problem (P0) can be reexpressed as
[…]
s.t. $\textrm{SINR}_{t}\geq\Gamma_{r},$ (28b)
${\sum}_{k=1}^{K_{d}}\|\mathbf{w}_{m,k}\|_{2}^{2}+\|\mathbf{W}^{r}_{m}\|_{F}^{2}\leq\textrm{P}^{BS}_{m},\forall m\in\mathcal{M}_{t},$ (28c)
$0\leq q_{l}\leq P_{u,l},\forall l\in\mathcal{K}_{u},$ (28d)
$\|\mathbf{t}^{0}_{m,n}\!-\!\mathbf{t}^{0}_{m,i}\|_{2}\!\geq\!D_{t},\forall n,i\!\in\!\mathcal{N}_{t},n\!\neq\!i,\forall m\!\in\!\mathcal{M}_{t},$ (28e)
$\|\mathbf{t}^{1}_{p,j}\!-\!\mathbf{t}^{1}_{p,s}\|_{2}\!\geq\!D_{t},\forall j,s\in\mathcal{N}_{r},j\neq s,\forall p\in\mathcal{M}_{r},$ (28f)
$\mathbf{t}^{0}_{m,n}\in\mathcal{C},\ \mathbf{t}^{1}_{p,j}\in\mathcal{C},\ \mathbf{r}_{0,k}\in\mathcal{C},\ \mathbf{r}_{1,l}\in\mathcal{C},$ (28g)
Next, we will adopt block coordinate ascent (BCA) method [24] to efficiently solve problem (P1) by updating different blocks of variables.
$\tilde{\mathrm{R}}_{d,k}\triangleq\textrm{log}(\omega_{d,k})-\omega_{d,k}\big{(}1-2\textrm{Re}\{\beta_{d,k}^{\ast}\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{k}\}+|\beta_{d,k}|^{2}({\sum}_{i=1}^{K_{d}}|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{i}|^{2}+\|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{W}}^{r}\|^{2}_{F}+{\sum}_{l=1}^{K_{u}}q_{l}|{h}_{du,k,l}|^{2}+\sigma_{d,k}^{2})\big{)}+1$ (29)
$\tilde{\mathrm{R}}_{u,l}\triangleq\textrm{log}(\omega_{u,l})-\omega_{u,l}\bigg{(}1-2\textrm{Re}\{\beta_{u,l}^{\ast}\sqrt{q_{l}}\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,l}\}$ (30)
$+|\beta_{u,l}|^{2}\big{(}{\sum}_{i=1}^{K_{u}}q_{i}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,i}|^{2}+{\sum}_{j=0}^{K_{t}}\sigma_{t,j}^{2}({\sum}_{k=1}^{K_{d}}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}|^{2}+\|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{F})+\sigma_{r}^{2}\|\hat{\mathbf{u}}_{l}^{H}\|^{2}_{2}\big{)}\bigg{)}+1$
III-B Optimizing auxiliary variables
With other variables being fixed, the analytical solutions of the auxiliary variables, i.e., $\{\omega_{d,k}\}$, $\{\beta_{d,k}\}$, $\{\omega_{u,l}\}$ and $\{\beta_{u,l}\}$, can be obtained from the derivation of WMMSE transformation proven in [23], as shown in (31)-(34).
$\beta^{\star}_{d,k}={\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{k}}/({{\sum}_{i=1}^{K_{d}}|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{i}|^{2}+\|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{W}}^{r}\|^{2}_{F}+{\sum}_{l=1}^{K_{u}}q_{l}|{h}_{du,k,l}|^{2}+\sigma_{d,k}^{2}}),$ (31)
$\omega^{\star}_{d,k}=\big{(}|1-\beta^{\ast}_{d,k}\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{k}|^{2}+|\beta^{\ast}_{d,k}|^{2}({\sum}_{i\neq k}^{K_{d}}|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{i}|^{2}+\|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{W}}^{r}\|^{2}_{F}+{\sum}_{l=1}^{K_{u}}q_{l}|{h}_{du,k,l}|^{2}+\sigma_{d,k}^{2})\big{)}^{-1},$ (32)
$\beta^{\star}_{u,l}={\sqrt{q_{l}}\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,l}}/({\sum_{i=1}^{K_{u}}q_{i}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,i}|^{2}+\sum_{j=0}^{K_{t}}\sigma_{t,j}^{2}(\sum_{k=1}^{K_{d}}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}|^{2}+\|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{F})+\sigma_{r}^{2}\|\hat{\mathbf{u}}_{l}^{H}\|^{2}_{2}}),$ (33)
$\omega^{\star}_{u,l}=\bigg{(}|1-\beta^{\ast}_{u,l}\sqrt{q_{l}}\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,l}|^{2}+|\beta^{\ast}_{u,l}|^{2}\big{(}{\sum}_{j=0}^{K_{t}}\sigma_{t,j}^{2}({\sum}_{k=1}^{K_{d}}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}|^{2}+\|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{F})+\sigma_{r}^{2}\|\hat{\mathbf{u}}_{l}^{H}\|^{2}_{2}\big{)}\bigg{)}^{-1}$ (34)
III-C Updating The TBS Beamformer
In this subsection, we investigate the optimization of all $M_{t}$ TBS beamformer $\{\mathbf{w}_{m,k},\mathbf{W}^{r}_{m}\}$ with other variables being given. Firstly, the functions $\tilde{\mathrm{R}}_{d,k}$ and $\tilde{\mathrm{R}}_{u,l}$, which are respectively defined in (29) and (30), are rewritten as
$-\!\sum_{k=1}^{K_{d}}\!\tilde{\mathrm{R}}_{\!d\!,k}\!\!=\!\!\!\sum_{k=1}^{K_{d}}\!(\!\hat{\mathbf{w}}^{H}_{k}\!\!\mathbf{A}_{1}\!\hat{\mathbf{w}}_{k}\!\!-\!\!2\textrm{Re}\{\!\mathbf{a}^{H}_{1,k}\hat{\mathbf{w}}_{k}\!\}\!)\!\!+\!\!\!\sum_{i=1}^{M_{t}}\!(\!\bar{\mathbf{w}}^{r}_{i}\!)^{\!H\!}\!\mathbf{A}_{\!2\!,i}\!\bar{\mathbf{w}}^{r}_{i}\!\!-\!\!c_{\!1\!,\!1},$ (35)
$-\sum_{l=1}^{K_{u}}\tilde{\mathrm{R}}_{u,l}\!=\!\sum_{k=1}^{K_{d}}\!\hat{\mathbf{w}}^{H}_{k}\mathbf{A}_{3}\hat{\mathbf{w}}_{k}\!+\!\sum_{i=1}^{M_{t}}\!(\bar{\mathbf{w}}^{r}_{i})^{H}\mathbf{A}_{4,i}\bar{\mathbf{w}}^{r}_{i}\!-\!c_{1,2},$ (36)
where $c_{1,1}$ and $c_{1,2}$ are constant terms and the newly introduced parameters in (35) and (36) being defined as follows
$\mathbf{a}_{1,k}\triangleq\omega_{d,k}\beta_{d,k}\hat{\mathbf{h}}_{d,k},\bar{\mathbf{w}}^{r}_{i}\triangleq\textrm{vec}({\mathbf{W}^{r}_{i}}),$ (37)
$\mathbf{A}_{1}\triangleq{\sum}_{k=1}^{K_{d}}\omega_{d,k}|\beta_{d,k}|^{2}\hat{\mathbf{h}}_{d,k}\hat{\mathbf{h}}_{d,k}^{H},\mathbf{A}_{2,i}\triangleq\mathbf{I}\otimes{\mathbf{h}}_{d,i,k}{\mathbf{h}}_{d,i,k}^{H},$
$\mathbf{A}_{3}\triangleq{\sum}_{l=1}^{K_{u}}{\sum}_{j=0}^{K_{t}}\beta_{u,l}|\beta_{u,l}|^{2}\sigma_{t,j}^{2}\hat{\mathbf{g}}_{t,j}\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{l}\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H},$
$\mathbf{A}_{4,i}\!\triangleq\!\!{\sum}_{l=1}^{K_{u}}{\sum}_{j=0}^{K_{t}}\beta_{u,l}|\beta_{u,l}|^{2}\big{(}\sigma_{t,j}^{2}\mathbf{I}\!\otimes\!({\mathbf{g}}_{t,i,j}\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{l}\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}{\mathbf{g}}_{t,i,j}^{H})\big{)}.$
Similarly, the constraint (28b) can also be rewritten as
$\textrm{SINR}_{t}\geq\Gamma_{r}\Longleftrightarrow{\sum}_{k=1}^{K_{d}}\hat{\mathbf{w}}_{k}^{H}\mathbf{A}_{11}\hat{\mathbf{w}}_{k}+{\sum}_{i=1}^{M_{t}}(\bar{\mathbf{w}}^{r}_{i})^{H}\mathbf{A}_{12,i}\bar{\mathbf{w}}^{r}_{i}$
$+c_{1,3}-{\sum}_{k=1}^{K_{d}}\hat{\mathbf{w}}_{k}^{H}\mathbf{A}_{9}\hat{\mathbf{w}}_{k}-{\sum}_{i=1}^{M_{t}}(\bar{\mathbf{w}}^{r}_{i})^{H}\mathbf{A}_{10,i}\bar{\mathbf{w}}^{r}_{i}\leq 0,$ (38)
where $c_{1,3}$ is constant and the above newly introduced coefficients are defined as
$\mathbf{A}_{9}\triangleq\sigma^{2}_{t,0}\hat{\mathbf{g}}_{t,0}\hat{\mathbf{g}}_{r,0}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,0}\hat{\mathbf{g}}_{t,0}^{H}/\Gamma_{r},$ (39)
$\mathbf{A}_{10,i}\triangleq\mathbf{I}\otimes\sigma^{2}_{t,0}{\mathbf{g}}_{t,i,0}\hat{\mathbf{g}}_{r,0}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,0}{\mathbf{g}}_{t,i,0}^{H}/\Gamma_{r},$
$\mathbf{A}_{11}\triangleq{\sum}_{j=1}^{K_{t}}\sigma^{2}_{t,j}\hat{\mathbf{g}}_{t,j}\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H},$
$\mathbf{A}_{12,i}\triangleq{\sum}_{j=1}^{K_{t}}\mathbf{I}\otimes\sigma^{2}_{t,j}{\mathbf{g}}_{t,i,j}\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,j}{\mathbf{g}}_{t,i,j}^{H}.$
Based on the above equivalent transformation, the problem of updating $\{\hat{\mathbf{w}}_{k},\bar{\mathbf{w}}^{r}_{i}\}$ can be formulated as
[…]
$+{\sum}_{i=1}^{M_{t}}(\bar{\mathbf{w}}^{r}_{i})^{H}\mathbf{A}_{6,i}\bar{\mathbf{w}}^{r}_{i}-\hat{c}_{1,2}$
s.t. ${\sum}_{k=1}^{K_{d}}\hat{\mathbf{w}}_{k}^{H}\mathbf{A}_{11}\hat{\mathbf{w}}_{k}\!+\!\!{\sum}_{i=1}^{M_{t}}(\bar{\mathbf{w}}^{r}_{i})^{H}\!\mathbf{A}_{12,i}\bar{\mathbf{w}}^{r}_{i}\!+\!c_{1,3}$ (40b)
-- ${\sum}_{k=1}^{K_{d}}\hat{\mathbf{w}}_{k}^{H}\mathbf{A}_{9}\hat{\mathbf{w}}_{k}-{\sum}_{i=1}^{M_{t}}(\bar{\mathbf{w}}^{r}_{i})^{H}\mathbf{A}_{10,i}\bar{\mathbf{w}}^{r}_{i}\leq 0,$
${\sum}_{k=1}^{K_{d}}\|\hat{\mathbf{w}}_{k}\mathbf{R}_{i}\|_{2}^{2}+\|\bar{\mathbf{w}}^{r}_{i}\|_{2}^{2}\leq\textrm{P}^{BS}_{i},\forall i\in\mathcal{M}_{t},$ (40c)
where $\mathbf{A}_{5}\triangleq\mathbf{A}_{1}+\mathbf{A}_{3}$, $\mathbf{A}_{6,i}\triangleq\mathbf{A}_{2,i}+\mathbf{A}_{4,i}$, $\hat{c}_{1,2}\triangleq c_{1,2}+c_{1,1}$, $\mathbf{R}_{i}\triangleq\textrm{blkdiag}([\mathbf{0},\cdots,\mathbf{0}_{i-1,i-1},\mathbf{I}_{i,i},\mathbf{0}_{i+1,i+1},\cdots,\mathbf{0}])\in\mathbb{C}^{M_{t}N_{t}\times M_{t}N_{t}}$, $\mathbf{0}\in\mathbb{C}^{N_{t}\times N_{t}}$, $\mathbf{I}\in\mathbb{C}^{N_{t}\times N_{t}}$.
It is obviously that the non-convex constraint (43b) makes the problem (P2) intractable. Inspired by the MM framework [20], we respectively linearize the quadratic terms $\hat{\mathbf{w}}_{k}^{H}\mathbf{A}_{9}\hat{\mathbf{w}}_{k}$ and $\bar{\mathbf{w}}^{H}_{r,i}\mathbf{A}_{10,i}\bar{\mathbf{w}}_{r,i}$ as follows
$\hat{\mathbf{w}}_{k}^{H}\mathbf{A}_{9}\hat{\mathbf{w}}_{k}\geq 2\textrm{Re}\{\hat{\mathbf{w}}_{k,0}^{H}\mathbf{A}_{9}(\hat{\mathbf{w}}_{k}-\hat{\mathbf{w}}_{k,0})\}+\hat{\mathbf{w}}_{k,0}^{H}\mathbf{A}_{9}\hat{\mathbf{w}}_{k,0}$
$=2\textrm{Re}\{\hat{\mathbf{w}}_{k,0}^{H}\mathbf{A}_{9}\hat{\mathbf{w}}_{k}\}-(\hat{\mathbf{w}}_{k,0}^{H}\mathbf{A}_{9}\hat{\mathbf{w}}_{k,0})^{\ast},$ (41)
$(\bar{\mathbf{w}}^{r}_{i})^{H}\mathbf{A}_{10,i}\bar{\mathbf{w}}^{r}_{i}\!\!\geq\!\!2\textrm{Re}\{\!(\bar{\mathbf{w}}^{r}_{i,0})^{H}\!\mathbf{A}_{10,i}(\!\bar{\mathbf{w}}^{r}_{i}\!\!-\!\!\bar{\mathbf{w}}^{r}_{i,0}\!)\!\}\!\!+\!\!(\bar{\mathbf{w}}^{r}_{i,0})^{H}\!\mathbf{A}_{10,i}\bar{\mathbf{w}}^{r}_{i,0}$
$=2\textrm{Re}\{(\bar{\mathbf{w}}^{r}_{i,0})^{H}\mathbf{A}_{10,i}\bar{\mathbf{w}}^{r}_{i}\}-((\bar{\mathbf{w}}^{r}_{i,0})^{H}\mathbf{A}_{10,i}\bar{\mathbf{w}}^{r}_{i,0})^{\ast},$ (42)
where $\hat{\mathbf{w}}_{k,0}$ and $\bar{\mathbf{w}}^{r}_{i,0}$ are feasible solutions obtained in the last iteration. Therefore, by replace the non-convex terms of constraint (43b) by (41) and (42), the problem (P2) can be rewritten as
[…]
$+{\sum}_{i=1}^{M_{t}}(\bar{\mathbf{w}}^{r}_{i})^{H}\mathbf{A}_{6,i}\bar{\mathbf{w}}^{r}_{i}-\hat{c}_{1,2}$
s.t. ${\sum}_{k=1}^{K_{d}}\hat{\mathbf{w}}_{k}^{H}\mathbf{A}_{11}\hat{\mathbf{w}}_{k}\!+\!\!{\sum}_{i=1}^{M_{t}}(\bar{\mathbf{w}}^{r}_{i})^{H}\!\mathbf{A}_{12,i}\bar{\mathbf{w}}^{r}_{i}\!+\!\hat{c}_{1,3}$ (43b)
-- $\!{\sum}_{k=1}^{K_{d}}2\textrm{Re}\{\mathbf{a}_{2,k}^{H}\hat{\mathbf{w}}_{k}\}\!-\!{\sum}_{i=1}^{M_{t}}2\textrm{Re}\{\mathbf{a}_{3,i}^{H}\bar{\mathbf{w}}^{r}_{i}\}\!\leq\!0,$
${\sum}_{k=1}^{K_{d}}\|\hat{\mathbf{w}}_{k}\mathbf{R}_{i}\|_{2}^{2}+\|\bar{\mathbf{w}}^{r}_{i}\|_{2}^{2}\leq\textrm{P}^{BS}_{i},\forall i\in\mathcal{M}_{t},$ (43c)
where $\mathbf{a}_{2,k}\triangleq\mathbf{A}_{9}^{H}\hat{\mathbf{w}}_{k,0}$, $\mathbf{a}_{3,i}\triangleq\mathbf{A}_{10,i}^{H}\bar{\mathbf{w}}^{r}_{i,0}$, $\hat{c}_{1,3}\triangleq\!-\!{\sum}_{k=1}^{K_{d}}\!(\hat{\mathbf{w}}_{k,0}^{H}\mathbf{A}_{9}\hat{\mathbf{w}}_{k,0})^{\ast}\!-\!{\sum}_{i=1}^{M_{t}}\!((\bar{\mathbf{w}}^{r}_{i,0})^{H}\mathbf{A}_{10,i}\bar{\mathbf{w}}^{r}_{i,0})^{\ast}+{c}_{1,3}$. The problem (P3) is a typical second order cone program (SOCP) and can be efficiently solved by standard numerical solver, such as CVX [25].
III-D Optimizing The UL User Transmission Power
In this subsection, we present the method for optimizing the UL users’ transmission power allocation. With other variables being fixed, the optimization problem of updating $\{q_{l}\}$ is formulated as
[…]
s.t. ${\sum}_{l=1}^{K_{u}}b_{3,l}q_{l}-c_{2,2}\leq 0,$ (44b)
$0\leq q_{l}\leq P_{u,l},\forall l\in\mathcal{K}_{u},$ (44c)
where $c_{2,1}$ and $c_{2,2}$ are constant terms, and the new parameters in problem (P4) are defined as
$b_{2,l}\triangleq 2\textrm{Re}\{\omega_{u,l}\beta_{u,l}^{\ast}\hat{\mathbf{u}}_{i}^{H}\hat{\mathbf{h}}_{u,l}\},b_{3,l}\triangleq|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{h}}_{u,l}|,$ (45)
$b_{1,l}\!\triangleq\!\!{\sum}_{i=1}^{K_{u}}\omega_{u,l}|\beta_{u,l}|^{2}|\hat{\mathbf{u}}_{i}^{H}\hat{\mathbf{h}}_{u,l}|^{2}\!\!+\!\!\!{\sum}_{k=1}^{K_{d}}\omega_{d,k}|\beta_{d,k}|^{2}|{h}_{du,k,l}|^{2}.$
The problem (P4) is an SOCP and can be directly solved by CVX.
III-E Optimizing The UL User Receiver Filter
With given other variables, the UL user receive filters $\{\hat{\mathbf{u}}_{l}\}$ optimization reduces to solving the following problem
[…]
where the above newly introduced coefficients are given as
$\mathbf{d}_{1,l}\triangleq\omega_{u,l}\beta_{u,l}^{\ast}\sqrt{q_{l}}\hat{\mathbf{h}}_{u,l},\mathbf{D}_{1}\triangleq\omega_{u,l}|\beta_{u,l}|^{2}\big{(}{\sum}_{i=1}^{K_{u}}q_{i}\hat{\mathbf{h}}_{u,i}\hat{\mathbf{h}}_{u,i}^{H}$ (47)
$\!+\!\!\!\!{\sum}_{j\!=\!0}^{K_{t}}\!\!\sigma_{t,j}^{2}\!(\!{\sum}_{k\!=\!1}^{K_{d}}\!\hat{\mathbf{g}}_{\!r,j}\!\hat{\mathbf{g}}_{\!t,j}^{H}\!\hat{\mathbf{w}}_{k}\!\hat{\mathbf{w}}_{k}^{H}\!\hat{\mathbf{g}}_{\!t,j}\hat{\mathbf{g}}_{\!r,j}^{H}\!\!+\!\!\hat{\mathbf{g}}_{\!r,j}\!\hat{\mathbf{g}}_{\!t,j}^{H}\!\hat{\mathbf{W}}_{\!r}\!\hat{\mathbf{W}}_{\!r}^{H}\!\hat{\mathbf{g}}_{\!t,j}\hat{\mathbf{g}}_{\!r,j}^{H}\!\!+\!\!\sigma_{r}^{2}\mathbf{I}\!\big{)}.$
Since the variables $\{\hat{\mathbf{u}}_{l}\}$ are not intricately coupled in (46a), the problem (P5) can be decomposed into $K_{u}$ independent sub-problems, which each subproblem can be written as
[…]
The subproblem (P6l) is a typical unconstrained convex quadratic problem. And then, by setting problem (P6l)’s derivative to zero, the optimal solution can be obtained as
$\hat{\mathbf{u}}_{l}^{\star}=\mathbf{D}_{l}^{-1}\mathbf{d}_{1,l},\forall l\in\mathcal{K}_{u}.$ (49)
III-F Optimizing The Sensing Filter
Given fixed other variables, the problem with respect to (w.r.t.) $\hat{\mathbf{u}}_{0}$ is reduced to a typical feasibility check problem, i.e., Phase-I problem [26], which is given as
[…]
where the newly defined coefficients are given as
$\mathbf{E}_{1}\!\!\triangleq\!\!\sigma_{t,0}^{2}(\!{\sum}_{k=1}^{K_{d}}\!\hat{\mathbf{g}}_{r,0}\hat{\mathbf{g}}_{t,0}^{H}\!\hat{\mathbf{w}}_{k}\!\hat{\mathbf{w}}_{k}^{H}\!\hat{\mathbf{g}}_{t,0}\hat{\mathbf{g}}_{r,0}^{H}\!\!+\!\!\hat{\mathbf{g}}_{r,0}\hat{\mathbf{g}}_{t,0}^{H}\hat{\mathbf{W}}_{r}\!\hat{\mathbf{W}}_{r}^{H}\hat{\mathbf{g}}_{t,0}\hat{\mathbf{g}}_{r,0}^{H}),$
$\mathbf{E}_{2}\!\triangleq\!\!{\sum}_{i=1}^{K_{u}}q_{i}\hat{\mathbf{h}}_{u,i}\hat{\mathbf{h}}_{u,i}^{H}\!\!+\!\!{\sum}_{j=1}^{K_{t}}\!\sigma_{t,j}^{2}\!(\!{\sum}_{k=1}^{K_{d}}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}\hat{\mathbf{g}}_{t,j}\hat{\mathbf{g}}_{r,j}^{H}$
$+\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\!\hat{\mathbf{W}}_{r}^{H}\hat{\mathbf{g}}_{t,j}\hat{\mathbf{g}}_{r,j}^{H})+\sigma_{r}^{2}\mathbf{I}.$ (51)
Since the variable $\hat{\mathbf{u}}_{0}$ only appears in the constraint (50b), we can turn to consider a sensing SINR maximization problem, which is formulated as
[…]
The above problem (P8) is typical Rayleigh quotient maximization. By applying the Rayleigh-Rits theorem, the optimal $\hat{\mathbf{u}}_{0}$ can be obtained by aligning with the eigenvector associated with the largest eigenvalue of the matrix $\mathbf{E}_{2}^{-1}\mathbf{E}_{1}$ [27].
Note that although maximizing (P8) does not directly impact the objective value, it will find a more “feasible” variable $\hat{\mathbf{u}}_{0}$, which can provide a larger margin to satisfy the constraint (50b) and hence benefit the optimization of other variables.
III-G Optimizing The TBS MAs’ Position
In this subsection, we discuss updating the position of $m$-th TBS’ $n$-th MA, i.e., $\mathbf{t}^{0}_{m,n}$, when other variables are given. Firstly, the function $\tilde{\mathrm{R}}_{d,k}$, which is defined in (29), can be rewritten as
$-\tilde{\mathrm{R}}_{d,k}=(\mathbf{h}^{0}_{m,k,n})^{H}\mathbf{F}_{5,m,k,n,n}\mathbf{h}^{0}_{m,k,n}$ (53)
$+\textrm{Re}\{(\mathbf{f}^{7}_{m,k,n})^{H}\mathbf{h}^{0}_{m,k,n}\}+c_{4,1,m,k,n},$
where $c_{4,1,m,k,n}$ is a constant and the above newly introduced coefficients are given as
$\mathbf{F}^{1}_{m,n,k,i}\triangleq(\mathbf{w}_{m,i}\mathbf{w}_{n,i}^{H})^{T}\otimes((\mathbf{\Sigma}^{0}_{n,k})^{H}\mathbf{h}^{0,0}_{n,k}(\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{\Sigma}^{0}_{n,k}),$ (54)
$\tilde{\mathbf{h}}^{0}_{m,k}\triangleq\textrm{vec}({\mathbf{H}^{0}_{m,k}}),\mathbf{f}_{m,k,j}^{1}\triangleq{\sum}_{j\neq m}^{M_{t}}(\mathbf{F}^{1}_{m,j,k,i})^{H}\tilde{\mathbf{h}}^{0}_{j,k},$
$\mathbf{F}^{2}_{m,k}\triangleq(\bar{\mathbf{w}}^{r}_{m}(\bar{\mathbf{w}}^{r}_{m})^{H})^{T}\otimes((\mathbf{\Sigma}^{0}_{m,k})^{H}\mathbf{h}^{0,0}_{m,k}(\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{\Sigma}^{0}_{n,k}),$
$\mathbf{f}^{2}_{m,k}\!\!\triangleq\!\!2\omega_{d,k}\beta_{d,k}(\mathbf{w}_{m,k}^{T}\!\!\otimes\!\!(\!(\mathbf{h}^{0,0}_{m,k})^{H}\!\mathbf{\Sigma}^{0}_{m,k})\!)^{H},\mathbf{F}^{4}_{m,k}\!\!\triangleq\!\omega_{d,k}\!|\!\beta_{d,k}\!|^{2}\!\mathbf{F}^{2}_{m,k},$
$\mathbf{F}^{3}_{m,m,k}\!\!\triangleq\!\!\!{\sum}_{i=1}^{K_{d}}\omega_{d,k}|\!\beta_{d,k}\!|^{2}\mathbf{F}^{1}_{m,m,k,i},\mathbf{f}^{4}_{m,k}\!\!\triangleq\!\!\!{\sum}_{i=1}^{K_{d}}\!2\omega_{d,k}|\!\beta_{d,k}\!|^{2}\!\mathbf{f}_{m,k,i}^{1},$
$\mathbf{F}^{5}_{m,k}\triangleq\mathbf{F}^{3}_{m,k}+\mathbf{F}^{4}_{m,k},\mathbf{f}^{6}_{m,k,n}\triangleq 2({\sum}_{i\neq n}^{N_{t}}(\bar{\mathbf{F}}^{5}_{m,k,i,n})^{H}\acute{\mathbf{h}}^{0}_{m,k,i}),$
$\mathbf{f}^{5}_{m,k}\triangleq\mathbf{f}^{4}_{m,k}-\mathbf{f}^{3}_{m,k},\acute{\mathbf{h}}^{0}_{m,k,n}\triangleq\tilde{\mathbf{h}}^{0}_{m,k}[(n-1)L^{t,0}_{m,k}\!+\!1:nL^{t,0}_{m,k}],$
$\bar{\mathbf{F}}^{5}_{m,k,i,j}\!\triangleq\!\mathbf{F}^{5}_{m,k}[(i-1)L^{t,0}_{m,k}\!+\!1\!:\!iL^{t,0}_{m,k},(j-1)L^{t,0}_{m,k}\!+\!1\!:\!jL^{t,0}_{m,k}],$
$\mathbf{f}^{7}_{m,k,n}\triangleq\mathbf{f}^{6}_{m,k,n}+\mathbf{f}^{5}_{m,k}[(n-1)L^{t,0}_{m,k}+1:nL^{t,0}_{m,k}].$
Furthermore, the function $\tilde{\mathrm{R}}_{u,l}$ defined in (30) is also rewritten as follows
${\sum}_{l=1}^{K_{u}}\tilde{\mathrm{R}}_{u,l}\!=\!\!\!{\sum}_{j=0}^{K_{t}}\!\!-\!\!\textrm{Re}\{(\mathbf{f}^{10}_{m,j,n})^{\ast}\mathbf{g}_{t,m,j}[n]\}\!\!+\!c_{4,2,m,n},$ (55)
where $c_{4,2,m,n}$ is a constant term and the new parameters in (55) being defined as follows
$\mathbf{F}^{7}_{l,j}\triangleq(\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{l}\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j})^{T}\otimes(\hat{\mathbf{W}}_{r}\hat{\mathbf{W}}_{r}^{H}),$ (56)
$\mathbf{F}^{6}_{l,j,k}\!\!\triangleq\!\!(\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{l}\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\!)^{T}\!\!\!\otimes\!\!(\!\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}\!),\mathbf{f}^{8}_{m,j}\!\triangleq\!\!\!{\sum}_{i\neq m}^{M_{t}}\!\!2(\!\bar{\mathbf{F}}^{9}_{j,m,n}\!)^{H}\!\mathbf{g}_{t,i,j},$
$\mathbf{F}^{8}_{l,j}\triangleq{\sum}_{k=1}^{K_{d}}\mathbf{F}^{6}_{l,j,k}+\mathbf{F}^{7}_{l,j},\mathbf{F}^{9}_{j}\triangleq{\sum}_{l=1}^{K_{u}}\omega_{u,l}|\beta_{u,l}|^{2}\sigma_{t,j}^{2}\mathbf{F}^{8}_{l,j},$
$\bar{\mathbf{F}}^{9}_{j,m,n}\triangleq\mathbf{F}^{9}_{j}[(m-1)N_{t}+1:mN_{t},(n-1)N_{t}+1:nN_{t}],$
$f^{9}_{m,j,n}\!\!\triangleq\!\!\!{\sum}_{i\neq n}^{N_{t}}\!(\bar{\mathbf{F}}^{9}_{j,m,m}[i,n])^{\ast}\mathbf{g}_{t,m,j}[i],f^{10}_{m,j,n}\!\triangleq\!f^{9}_{m,j,n}\!+\!\mathbf{f}^{8}_{m,j}[n].$
Similarly, the sensing SINR constraint (28b) is rewritten as
$\textrm{SINR}_{t}\geq\Gamma_{r}\Longleftrightarrow{\sum}_{i=1}^{K_{t}}\textrm{Re}\{(f^{13}_{i,m,n})^{\ast}\mathbf{g}_{t,m,i}[n]\}$ (57)
$-\textrm{Re}\{(f^{16}_{m,n})^{\ast}\mathbf{g}_{t,m,0}[n]\}+c_{4,3,m,n}\leq 0,$
where $c_{4,3,m,n}$ represents a constant term and the above newly introduced coefficients defined in (58).
$\mathbf{F}^{10}_{k}\triangleq\sigma_{t,0}^{2}(\hat{\mathbf{g}}_{r,0}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,0})^{T}\otimes(\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}),\mathbf{F}^{11}\triangleq\sigma_{t,0}^{2}(\hat{\mathbf{g}}_{r,0}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,0})^{T}\otimes(\hat{\mathbf{W}}_{r}\hat{\mathbf{W}}_{r}^{H}),\mathbf{F}^{12}\triangleq({\sum}_{k=1}^{K_{d}}\mathbf{F}^{10}_{k}+\mathbf{F}^{11})/\Gamma,$ (58)
$\mathbf{F}^{13}_{k,j}\triangleq\sigma_{t,j}^{2}(\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,j})^{T}\otimes(\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}),\mathbf{F}^{14}_{j}\triangleq\sigma_{t,j}^{2}(\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,j})^{T}\otimes(\hat{\mathbf{W}}_{r}\hat{\mathbf{W}}_{r}^{H}),\mathbf{F}^{15}_{j}\triangleq{\sum}_{k=1}^{K_{d}}\mathbf{F}^{13}_{k,j}+\mathbf{F}^{14}_{j},$
$\bar{\mathbf{F}}^{15}_{j,m,n}\triangleq\mathbf{F}^{15}_{j}[(m-1)N_{t}+1:mN_{t},(n-1)N_{t}+1:nN_{t}],f^{12}_{j,m,n}\triangleq{\sum}_{z\neq n}^{N_{t}}2(\bar{\mathbf{F}}^{15}_{j,m,m}[z,n])^{\ast}\mathbf{g}_{t,m,j}[z],$
$\bar{\mathbf{F}}^{12}_{m,n}\triangleq\mathbf{F}^{12}[(m-1)N_{t}+1:mN_{t},(n-1)N_{t}+1:nN_{t}],{f}^{15}_{m,n}\triangleq{\sum}_{i\neq n}^{N_{t}}2(\bar{\mathbf{F}}^{12}_{m,m}[i,n])^{\ast}\mathbf{g}_{t,m,0}[i],$
$\mathbf{f}^{11}_{j,m}\triangleq{\sum}_{i\neq m}^{M_{t}}2(\bar{\mathbf{F}}^{15}_{j,i,m})^{H}\mathbf{g}_{t,i,j},\mathbf{f}^{14}_{m}\triangleq{\sum}_{n\neq m}^{M_{t}}2(\bar{\mathbf{F}}^{12}_{m,n})^{H}\mathbf{g}_{t,n,0},f^{13}_{j,m,n}\triangleq f^{12}_{j,m,n}+\mathbf{f}^{11}_{j,m}[n],{f}^{16}_{m,n}\triangleq{f}^{15}_{m,n}+\mathbf{f}^{14}_{m}[n].$
Based on the above equivalent transformation, the subproblem w.r.t $\mathbf{t}^{0}_{m,n}$ is formulated as
[…]
$+\textrm{Re}\{(\mathbf{f}^{7}_{m,k,n})^{H}\mathbf{h}^{0}_{m,k,n}(\mathbf{t}^{0}_{m,n})\}\big{)}+\hat{c}_{4,2,m,k,n}$
$+{\sum}_{j=0}^{K_{t}}\textrm{Re}\{({f}^{10}_{m,j,n})^{\ast}\mathbf{g}_{t,m,j}[n](\mathbf{t}^{0}_{m,n})\}$ (59a)
s.t. ${\sum}_{i=1}^{K_{t}}\textrm{Re}\{(f^{13}_{i,m,n})^{\ast}\mathbf{g}_{t,m,i}[n](\mathbf{t}^{0}_{m,n})\}$ (59b)
$-\textrm{Re}\{(f^{16}_{m,n})^{\ast}\mathbf{g}_{t,m,0}[n](\mathbf{t}^{0}_{m,n})\}+c_{4,20,m,n}\leq 0,$
$\|\mathbf{t}^{0}_{m,n}-\mathbf{t}^{0}_{m,i}\|_{2}\geq D_{t},\forall n,i\in\mathcal{N}_{t},n\neq i,$ (59c)
$\mathbf{t}^{0}_{m,n}\in\mathcal{C},$ (59d)
where $\hat{c}_{4,2,m,k,n}$ is a constant. Obviously, the current form of problem (P9) is difficult to obtain the optimal $\mathbf{t}^{0}_{m,n}$ due to the non-convex objective function and constraints. In the following, we adopt the MM methodology [20] to resolve the above difficulty.
Firstly, for any positive semidefinite matrix $\mathbf{F}$, we have the following inequality
$\mathbf{h}^{H}\mathbf{F}\mathbf{h}$ (60)
$=(\mathbf{h}\!-\!\mathbf{h}_{0})^{H}\mathbf{F}(\mathbf{h}\!-\!\mathbf{h}_{0})+2\mathrm{Re}\{\mathbf{h}_{0}^{H}\mathbf{F}(\mathbf{h}\!-\!\mathbf{h}_{0})\}+\mathbf{h}_{0}^{H}\mathbf{F}\mathbf{h}_{0}$
$\leq\lambda_{max}(\mathbf{F})\|\mathbf{h}-\mathbf{h}_{0}\|^{2}_{2}+2\mathrm{Re}\{\mathbf{h}_{0}^{H}\mathbf{F}(\mathbf{h}-\mathbf{h}_{0})\}+\mathbf{h}_{0}^{H}\mathbf{F}\mathbf{h}_{0}$
$=\lambda_{max}(\mathbf{F})\|\mathbf{h}\|^{2}_{2}+2\mathrm{Re}\{(\mathbf{F}\mathbf{h}_{0}-\lambda_{max}(\mathbf{F})\mathbf{h}_{0})^{H}\mathbf{h}\}+c_{F},$
with $\lambda_{max}(\mathbf{F})$ being the maximal eigenvalue of $\mathbf{F}$, $c_{F}$ is a constant, $\mathbf{h}_{0}\triangleq\mathbf{h}^{0}_{m,k,n,0}(\mathbf{t}^{0}_{m,n,0})$, $\mathbf{t}^{0}_{m,n,0}=[x^{0}_{m,n,0},y^{0}_{m,n,0}]^{T}$ is the latest value of $\mathbf{t}^{0}_{m,n}$. Next, by leveraging the arguments presented in (60), we can construct a tight upper-bound for $(\mathbf{h}^{0}_{m,k,n}(\mathbf{t}^{0}_{m,n}))^{H}\mathbf{F}_{m,k,n}\mathbf{h}^{0}_{m,k,n}(\mathbf{t}^{0}_{m,n})$, which can be formulated in (61), where $\|\mathbf{h}^{0}_{m,k,n}\|^{2}_{2}=\|\mathbf{h}^{0}_{m,k,n,0}\|^{2}_{2}=L^{t,0}_{m,k}$, $c_{4,4,m,k,n}\triangleq 2\lambda_{max}(\mathbf{F}_{5,m,k,n,n})L^{t,0}_{m,k}-((\mathbf{h}^{0}_{m,k,n,0})^{H}\mathbf{F}_{5,m,k,n,n}\mathbf{h}^{0}_{m,k,n,0})^{\ast}$, $\mathbf{f}^{17}_{m,k,n}\!\triangleq\!2((\mathbf{F}_{5,m,k,n,n})^{H}\mathbf{h}^{0}_{m,k,n,0}\!-\!\lambda_{max}(\mathbf{F}_{5,m,k,n,n})\mathbf{h}^{0}_{m,k,n,0})$ and $\mathbf{f}^{18}_{m,k,n}\!\triangleq\!\mathbf{f}^{17}_{m,k,n}\!+\!\mathbf{f}^{7}_{m,k,n}$.
$(\mathbf{h}^{0}_{m,k,n}(\mathbf{t}^{0}_{m,n}))^{H}\mathbf{F}_{5,m,k,n,n}\mathbf{h}^{0}_{m,k,n}(\mathbf{t}^{0}_{m,n})+\textrm{Re}\{(\mathbf{f}^{7}_{m,k,n})^{H}\mathbf{h}^{0}_{m,k,n}(\mathbf{t}^{0}_{m,n})\}$ (61)
$\leq\lambda_{max}(\mathbf{F}_{5,m,k,n,n})\|\mathbf{h}^{0}_{m,k,n}\|^{2}_{2}-2\textrm{Re}\{\lambda_{max}(\!\mathbf{F}_{5,m,k,n,n}\!)(\mathbf{h}^{0}_{m,k,n,0})^{H}\mathbf{h}^{0}_{m,k,n}\}\!+\!\!\lambda_{max}(\!\mathbf{F}_{5,m,k,n,n}\!)\|\mathbf{h}^{0}_{m,k,n,0}\|^{2}_{2}$
+2Re {($+\!\!2\textrm{Re }\!\!\{(\mathbf{h}^{0}_{m,k,n,0})^{H}\!\mathbf{F}_{5,m,k,n,n}\mathbf{h}^{0}_{m,k,n}\}-\!\!2\textrm{Re }\!\!\{(\mathbf{h}^{0}_{m,k,n,0})^{H}\!\mathbf{F}_{5,m,k,n,n}\mathbf{h}^{0}_{m,k,n,0}\}\!\!+\!\!(\mathbf{h}^{0}_{m,k,n,0})^{H}\mathbf{F}_{5,m,k,n,n}\mathbf{h}^{0}_{m,k,n,0}$
$=\textrm{Re}\{(\mathbf{f}^{17}_{m,k,n})^{H}\mathbf{h}^{0}_{m,k,n}\}+c_{4,21,m,k,n}+\textrm{Re}\{(\mathbf{f}^{7}_{m,k,n})^{H}\mathbf{h}^{0}_{m,k,n}(\mathbf{t}^{0}_{m,n})\}=\textrm{Re}\{(\mathbf{f}^{18}_{m,k,n})^{H}\mathbf{h}^{0}_{m,k,n}\}+c_{4,4,m,k,n}$
Next, to tackle the challenging terms $\textrm{Re}\{(\mathbf{f}^{18}_{m,k,n})^{H}\mathbf{h}^{0}_{m,k,n}\}$, $\textrm{Re}\{({f}^{10}_{m,j,n})^{\ast}\mathbf{g}_{t,m,j}[n](\mathbf{t}^{0}_{m,n})\}$, $\textrm{Re}\{(f^{13}_{i,m,n})^{\ast}\mathbf{g}_{t,m,i}[n](\mathbf{t}^{0}_{m,n})\}$ and $\textrm{Re}\{(f^{16}_{m,n})^{\ast}\mathbf{g}_{t,m,0}[n](\mathbf{t}^{0}_{m,n})\}$, we introduce the following results that are proved in [3].
Lemma 1.
Assume that $\mathbf{f}\in\mathbb{C}^{L\times 1}$, $\mathbf{h}(\mathbf{t})\in\mathbb{C}^{L\times 1}$ and $\mathbf{t}=[x,y]^{T}$. The convex lower and upper bounds of the function $h_{t}(\mathbf{t})\triangleq\textrm{Re}\{\mathbf{f}^{H}\mathbf{h}(\mathbf{t})\}=\sum_{l=1}^{L}|\mathbf{f}[l]||\mathbf{h}(\mathbf{t})[l]|\cos(\angle\mathbf{h}(\mathbf{t})[l]-\angle\mathbf{f}[l])$ can be written as
$h_{t}(\mathbf{t})\!\geq\!h_{t}(\mathbf{t}_{0})\!+\!\nabla h_{t}^{T}(\mathbf{t}_{0})(\mathbf{t}\!-\!\mathbf{t}_{0})\!-\!\frac{\delta}{2}(\mathbf{t}\!-\!\mathbf{t}_{0})^{T}(\mathbf{t}\!-\!\mathbf{t}_{0}),$ (62)
$h_{t}(\mathbf{t})\!\leq\!h_{t}(\mathbf{t}_{0})\!+\!\nabla h_{t}^{T}(\mathbf{t}_{0})(\mathbf{t}\!-\!\mathbf{t}_{0})\!+\!\frac{\delta}{2}(\mathbf{t}\!-\!\mathbf{t}_{0})^{T}(\mathbf{t}\!-\!\mathbf{t}_{0}),$ (63)
respectively, where $\mathbf{t}_{0}$ is obtained from the last iteration, the symbols |||\ || | and $\angle$ denote the operations of obtaining the amplitude and phase, respectively, $\delta$ is a positive real number satisfying $\delta=\frac{8\pi^{2}}{\lambda^{2}}{\sum}_{l=1}^{L}|\mathbf{f}[l]|$, $\nabla h_{t}(\mathbf{t})=\big{[}\frac{\partial h_{t}(\mathbf{t})}{\partial x},\frac{\partial h_{t}(\mathbf{t})}{\partial y}\big{]}^{T}\in\mathbb{C}^{2\times 1}$ denotes the gradient vector of $h_{t}(\mathbf{t})$ over $\mathbf{t}$.
Firstly, by using the inequality (63) in Lemma 1, we can obtain convex upper-bound of $\textrm{Re}\{(\mathbf{f}^{18}_{m,k,n})^{H}\mathbf{h}^{0}_{m,k,n}\}$ as shown in (64),
$\textrm{Re}\{(\mathbf{f}^{18}_{m,k,n})^{H}\mathbf{h}^{0}_{m,k,n}\}={\sum}_{l=1}^{L^{t,0}_{m,k}}|\mathbf{f}^{18}_{m,k,n}[l]|\cos\big{(}\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{m,n})^{T}\mathbf{a}^{t,0}_{m,k,l}-\angle\mathbf{f}^{18}_{m,k,n}[l]\big{)}\triangleq h^{t0}_{,m,k,n}(\mathbf{t}^{0}_{m,n})$ (64)
$\leq h^{t0}_{m,k,n}(\mathbf{t}^{0}_{m,n,0})+\nabla^{T}h^{t0}_{m,k,n}(\mathbf{t}^{0}_{m,n,0})(\mathbf{t}^{0}_{m,n,0}-\mathbf{t}^{0}_{m,n})+\frac{\delta_{1,m,k,n}}{2}(\mathbf{t}^{0}_{m,n,0}-\mathbf{t}^{0}_{m,n})^{T}(\mathbf{t}^{0}_{m,n,0}-\mathbf{t}^{0}_{m,n})$
$=\frac{\delta_{1,m,k,n}}{2}\|\mathbf{t}^{0}_{m,n}\|^{2}_{2}+(\mathbf{f}^{19}_{m,k,n})^{T}\mathbf{t}^{0}_{m,n}+c_{4,5,m,k,n}$
where $c_{4,5,m,k,n}$ is a constant term, $\delta_{1,m,k,n}=\frac{8\pi^{2}}{\lambda^{2}}{\sum}_{l=1}^{L^{t,0}_{m,k}}|\mathbf{f}^{18}_{m,k,n}[l]|$, $\mathbf{f}^{19}_{m,k,n}\triangleq\nabla h^{t0}_{m,k,n}(\mathbf{t}^{0}_{m,n,0})-\delta_{1,m,k,n}\mathbf{t}^{0}_{m,n,0}$, the elements of $\nabla h^{t0}_{m,k,n}(\mathbf{t}^{0}_{m,n,0})=\big{[}\frac{\partial h^{t0}_{m,k,n}(\mathbf{t}^{0}_{m,n,0})}{\partial x^{0}_{m,n,0}},\frac{\partial h^{t0}_{m,k,n}(\mathbf{t}^{0}_{m,n,0})}{\partial y^{0}_{m,n,0}}\big{]}^{T}$ are respectively formulated as
$\frac{\partial h^{t0}_{m,k,n}(\mathbf{t}^{0}_{m,n,0})}{\partial x^{0}_{m,n,0}}=-{\sum}_{l=1}^{L^{t,0}_{m,k}}\frac{2\pi}{\lambda}|\mathbf{f}^{18}_{m,k,n}[l]|\cos(\theta^{t,0}_{m,k,l})$ (65)
$\sin(\phi^{t,0}_{m,k,l})\sin\big{(}\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{m,n})^{T}\mathbf{a}^{t,0}_{m,k,l}-\angle\mathbf{f}^{18}_{m,k,n}[l]\big{)},$
$\frac{\partial h^{t0}_{m,k,n}(\mathbf{t}^{0}_{m,n,0})}{\partial y^{0}_{m,n,0}}=-{\sum}_{l=1}^{L^{t,0}_{m,k}}\frac{2\pi}{\lambda}|\mathbf{f}^{18}_{m,k,n}[l]|\sin(\theta^{t,0}_{m,k,l})$ (66)
$\sin\big{(}\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{m,n})^{T}\mathbf{a}^{t,0}_{m,k,l}-\angle\mathbf{f}^{18}_{m,k,n}[l]\big{)}.$
Similarly, the terms $\textrm{Re}\{({f}^{10}_{m,j,n})^{\ast}\mathbf{g}_{t,m,j}[n](\mathbf{t}^{0}_{m,n})\}$ and $\textrm{Re}\{(f^{13}_{i,m,n})^{\ast}\mathbf{g}_{t,m,i}[n](\mathbf{t}^{0}_{m,n})\}$ can be respectively convexified by leveraging (63) as follows
$\textrm{Re}\{({f}^{10}_{m,j,n})^{\ast}\mathbf{g}_{t,m,j}[n](\mathbf{t}^{0}_{m,n})\}=|{f}^{10}_{m,j,n}||\mathbf{g}_{t,m,j}[n]|$ (67)
$\cos\big{(}\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{m,n})^{T}\mathbf{a}^{4}_{m,j}-\angle{f}^{10}_{m,j,n}\big{)}\triangleq g_{t1,m,j,n}(\mathbf{t}^{0}_{m,n})$
$\leq\frac{\delta_{2,m,j,n}}{2}\|\mathbf{t}^{0}_{m,n}\|^{2}_{2}+(\mathbf{f}^{20}_{m,j,n})^{T}\mathbf{t}^{0}_{m,n}+c_{4,6,m,j,n},$
$\textrm{Re}\{(f^{13}_{i,m,n})^{\ast}\mathbf{g}_{t,m,i}[n](\mathbf{t}^{0}_{m,n})\}=|f^{13}_{i,m,n}||\mathbf{g}_{t,m,i}[n]|$ (68)
$\cos\big{(}\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{m,n})^{T}\mathbf{a}^{4}_{m,i}-\angle f^{13}_{i,m,n}\big{)}\triangleq g_{t2,i,m,n}(\mathbf{t}^{0}_{m,n})$
$\leq\frac{\delta_{4,i,m,n}}{2}\|\mathbf{t}^{0}_{m,n}\|^{2}_{2}+(\mathbf{f}^{22}_{i,m,n})^{T}\mathbf{t}^{0}_{m,n}+c_{4,7,i,m,n},$
with $c_{4,6,m,j,n}$ and $c_{4,7,i,m,n}$ being constants, $\delta_{2,m,j,n}\triangleq\frac{8\pi^{2}}{\lambda^{2}}|{f}^{10}_{m,j,n}||\mathbf{g}_{t,m,j}[n]|$, $\delta_{4,i,m,n}\triangleq\frac{8\pi^{2}}{\lambda^{2}}|f^{13}_{i,m,n}||\mathbf{g}_{t,m,i}[n]|$, $\mathbf{f}^{20}_{m,j,n}\triangleq\nabla g_{t1,m,j,n}(\mathbf{t}^{0}_{m,n})-\delta_{2,m,j,n}\mathbf{t}^{0}_{m,n,0}$, $\mathbf{f}^{22}_{i,m,n}\triangleq\nabla g_{t2,i,m,n}(\mathbf{t}^{0}_{m,n})-\delta_{4,i,m,n}\mathbf{t}^{0}_{m,n,0}$, $\nabla g_{t1,m,j,n}(\mathbf{t}^{0}_{m,n})$ and $\nabla g_{t1,m,j,n}(\mathbf{t}^{0}_{m,n})$ are the gradient vectors, and the details of their derivations are similar to (65) and (66) and are omitted to avoid repetition.
And then, we proceed to cope with non-convex term $\textrm{Re}\{(f^{16}_{i,m,n})^{\ast}\mathbf{g}_{t,m,0}[n](\mathbf{t}^{0}_{m,n})\}$. Unlike the upper-bounding procedures presented in (64) to (68), this time we require a convex lower bound for $\textrm{Re}\{(f^{16}_{m,n})^{\ast}\mathbf{g}_{t,m,0}[n](\mathbf{t}^{0}_{m,n})\}$ by applying (62) in Lemma 1, which is given as follows
$\textrm{Re}\{(f^{16}_{m,n})^{\ast}\mathbf{g}_{t,m,0}[n](\mathbf{t}^{0}_{m,n})\}=|f^{16}_{m,n}||\mathbf{g}_{t,m,0}[n]|$ (69)
$\cos\big{(}\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{m,n})^{T}\mathbf{a}^{4}_{m,0}-\angle f^{16}_{m,n}\big{)}\triangleq g_{t3,m,n}(\mathbf{t}^{0}_{m,n})$
$\geq-\frac{\delta_{6,m,n}}{2}\|\mathbf{t}^{0}_{m,n}\|^{2}_{2}+(\mathbf{f}^{24}_{m,n})^{T}\mathbf{t}^{0}_{m,n}+c_{4,8,m,n},$
where $c_{4,8,m,n}$ is constant, $\delta_{6,m,n}\triangleq\frac{8\pi^{2}}{\lambda^{2}}|f^{16}_{m,n}||\mathbf{g}_{t,m,0}[n]|$, $\mathbf{f}^{24}_{m,n}\triangleq\nabla g_{t3,m,n}(\mathbf{t}^{0}_{m,n})+\delta_{6,m,n}\mathbf{t}^{0}_{m,n,0}$, $\nabla g_{t3,m,n}(\mathbf{t}^{0}_{m,n})$ is the the gradient vector, and the details of its derivation (refer to (65) and (66)) are omitted to avoid repetition.
To make (P9) tractable, we still need to convexify the constraint (59c). Inspired by the MM framework [20], the constraint (59c) can be linearized as follows
$\|\mathbf{t}^{0}_{m,n}-\mathbf{t}^{0}_{m,i}\|_{2}\geq\frac{(\mathbf{t}^{0}_{m,n,0}-\mathbf{t}^{0}_{m,i})^{T}}{\|\mathbf{t}^{0}_{m,n,0}-\mathbf{t}^{0}_{m,i}\|_{2}}(\mathbf{t}^{0}_{m,n}-\mathbf{t}^{0}_{m,i}).$ (70)
Based on the above MM transformation, the update of $\mathbf{t}^{0}_{m,n}$ can be performed by solving a convex surrogate of the original challenging problem (P9), as follows
[…]
s.t. $\delta_{7,m,n}\|\mathbf{t}^{0}_{m,n}\|^{2}_{2}\!+\!(\mathbf{f}^{25}_{m,n})^{T}\mathbf{t}^{0}_{m,n}\!+\!c_{4,10,m,n}\leq 0,$ (71b)
$\frac{(\mathbf{t}^{0}_{m,n,0}\!\!-\!\mathbf{t}^{0}_{m,i})^{T}}{\|\mathbf{t}^{0}_{m,n,0}\!\!-\!\mathbf{t}^{0}_{m,i}\|_{2}}\!(\mathbf{t}^{0}_{m,n}\!\!\!-\!\!\mathbf{t}^{0}_{m,i})\!\!\geq\!\!D_{t},\!\forall\!n,\!i\!\!\in\!\!\mathcal{N}_{t},n\!\!\neq\!\!i,$ (71c)
$\mathbf{t}^{0}_{m,n}\in\mathcal{C},$ (71d)
with $c_{4,9,m,n}$ and $c_{4,10,m,n}$ being constant terms, $\delta_{3,m,n}\triangleq\sum_{k=1}^{K_{d}}\frac{\delta_{1,m,k,n}}{2}+\sum_{j=0}^{K_{t}}\frac{\delta_{2,m,j,n}}{2}$, $\delta_{7,m,n}\triangleq\sum_{j=1}^{K_{t}}\frac{\delta_{4,j,m,n}}{2}+\frac{\delta_{6,m,n}}{2}$, $\mathbf{f}^{21}_{m,n}\triangleq\sum_{k=1}^{K_{d}}\mathbf{f}^{19}_{m,k,n}+\sum_{j=0}^{K_{t}}\mathbf{f}^{20}_{m,j,n}$ and $\mathbf{f}^{25}_{m,n}\triangleq\sum_{j=1}^{K_{t}}\mathbf{f}^{22}_{j,m,n}-\mathbf{f}^{24}_{m,n}$. Clearly, problem (P10) is convex and can be numerically solved.
III-H Optimizing The RBS MAs’ Position
In this subsection, with other variables being given, we investigate the update of the posotion of the $m$-th RBS’ $n$-th MA, i.e., $\mathbf{t}^{1}_{m,n}$. Firstly, the function $\tilde{\mathrm{R}}_{u,l}$ given in (30) is rewritten as
$\!-\!{\sum}_{l=1}^{K_{u}}\tilde{\mathrm{R}}_{u,l}\!=\!{\sum}_{l=1}^{K_{u}}\big{(}(\mathbf{h}^{1}_{m,l,n})^{H}\mathbf{D}^{7}_{m,l,n,n}\mathbf{h}^{1}_{m,l,n}$ (72)
$\!+\!\!\textrm{Re}\{(\mathbf{d}^{9}_{m,l,n})^{H}\mathbf{h}^{1}_{m,l,n}\}\big{)}\!\!+\!\!\!{\sum}_{j=0}^{K_{t}}\textrm{Re}\{(d^{11}_{j,m,n})^{\ast}\mathbf{g}_{r,m,j}[n]\}\!\!+\!\!c_{5,1,m,n},$
with $c_{5,1,m,n}$ being a constant and the parameters in (72) are defined as follows in (73).
$\mathbf{D}^{1}_{m,n,l,i}\triangleq\omega_{u,l}|\beta_{u,l}|^{2}(\mathbf{u}_{m,l}\mathbf{u}_{n,l}^{H})^{T}\otimes(q_{i}(\mathbf{\Sigma}^{1}_{n,i})^{H}\mathbf{h}^{0,1}_{n,i}\mathbf{h}^{0,1}_{m,i}\mathbf{\Sigma}^{1}_{m,i}),\hat{\mathbf{h}}^{1}_{m,l}\triangleq\textrm{vec}(\mathbf{H}^{1}_{m,l}(\mathbf{t}^{1}_{m})),\mathbf{d}^{1}_{m,l,i}\triangleq 2({\sum}_{j\neq m}^{M_{t}}(\mathbf{D}^{1}_{m,j,l,i})^{H}\hat{\mathbf{h}}^{1}_{j,i}),$ (73)
$\mathbf{d}^{2}_{m,l}\triangleq 2\omega_{u,l}\beta_{u,l}\sqrt{q_{l}}\mathbf{u}_{m,l}^{H}\otimes((\mathbf{h}^{0,1}_{m,l})^{H}\mathbf{\Sigma}^{1}_{m,l}),\mathbf{D}^{2}_{j,l}\triangleq{\sum}_{k=1}^{K_{d}}[(\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}\hat{\mathbf{g}}_{t,j})^{T}\otimes\sigma^{2}_{t,j}(\hat{\mathbf{u}}_{l}\hat{\mathbf{u}}_{l}^{H})],\mathbf{D}^{4}_{j,l}\triangleq\omega_{u,l}|\beta_{u,l}|^{2}(\mathbf{D}^{2}_{j,l}+\mathbf{D}^{3}_{j,l}),$
$\mathbf{D}^{3}_{j,l}\triangleq(\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}^{r}(\hat{\mathbf{W}}^{r})^{H}\hat{\mathbf{g}}_{t,j})^{T}\otimes\sigma^{2}_{t,j}(\hat{\mathbf{u}}_{l}\hat{\mathbf{u}}_{l}^{H}),\bar{\mathbf{D}}^{4}_{j,l,m,n}\triangleq\mathbf{D}^{4}_{j,l}[(m-1)N_{r}+1:mN_{r},(n-1)N_{r}+1:nN_{r}],$
$\mathbf{d}^{3}_{j,l,m}\triangleq 2({\sum}_{i\neq m}^{M_{r}}(\bar{\mathbf{D}}^{4}_{j,l,i,m})^{H}\mathbf{g}_{r,i,j}),\mathbf{D}^{5}_{m,i}\triangleq{\sum}_{l=1}^{K_{u}}\mathbf{D}^{1}_{m,m,l,i},\mathbf{d}^{4}_{m,i}\triangleq{\sum}_{l=1}^{K_{u}}\mathbf{d}^{1}_{m,l,i}-\mathbf{d}^{2}_{m,l},\mathbf{D}^{6}_{j,m}\triangleq{\sum}_{l=1}^{K_{u}}\bar{\mathbf{D}}^{4}_{j,l,m,m},$
$\mathbf{d}^{6}_{j,m}\triangleq{\sum}_{l=1}^{K_{u}}\mathbf{d}^{3}_{j,l,m},\mathbf{D}^{7}_{m,l,n,j}\triangleq\mathbf{D}^{5}_{m,l}[(n-1)L^{t,1}_{m,l}+1:nL^{t,1}_{m,l},(j-1)L^{t,1}_{m,l}+1:jL^{t,1}_{m,l}],\mathbf{d}^{8}_{m,l,n}\triangleq 2({\sum}_{j\neq n}^{N_{r}}(\mathbf{D}^{7}_{m,l,n,j})^{H}\mathbf{h}^{1}_{m,l,n}),$
$\mathbf{d}^{9}_{m,l,n}\triangleq\mathbf{d}^{8}_{m,l,n}+\mathbf{d}^{4}_{m,l}[(n-1)L^{t,1}_{m,l}+1:nL^{t,1}_{m,l}],d^{10}_{j,m,n}\triangleq 2({\sum}_{i\neq n}^{N_{r}}(\mathbf{D}^{6}_{j,m}[i,n])^{\ast}{\mathbf{g}}_{r,m,j}[i])+\mathbf{d}^{6}_{j,m}[n].$
Furthermore, the sensing SINR constraint (28b) can be rewritten as
$\textrm{SINR}_{t}\geq\Gamma_{r}\Longleftrightarrow{\sum}_{l=1}^{K_{u}}\big{(}(\mathbf{h}^{1}_{m,l,n})^{H}\mathbf{D}^{15}_{l,m,n,n}\mathbf{h}^{1}_{m,l,n}$ (74)
$\!+\!\!\textrm{Re}\{(\mathbf{d}^{16}_{l,m,n})^{H}\mathbf{h}^{1}_{m,l,n}\}\big{)}+{\sum}_{j=1}^{K_{t}}\big{(}\textrm{Re}\{(d^{17}_{j,m,n})^{\ast}\mathbf{g}_{r,m,j}[n]\}\big{)}$
$-\textrm{Re}\{(d^{19}_{m,n})^{\ast}\mathbf{g}_{r,m,0}[n]\}+c_{5,2,m,n}\leq 0,$
where $c_{5,2,m,n}$ is a constant term and the above newly introduced coefficients given in (75).
$\mathbf{D}^{10}\triangleq\big{(}{\sum}_{k=1}^{K_{d}}[(\hat{\mathbf{g}}_{t,0}^{H}\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}\hat{\mathbf{g}}_{t,0})^{T}\otimes\sigma^{2}_{t,0}(\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H})]+(\hat{\mathbf{g}}_{t,0}^{H}\hat{\mathbf{W}}^{r}(\hat{\mathbf{W}}^{r})^{H}\hat{\mathbf{g}}_{t,0})^{T}\otimes\sigma^{2}_{t,0}(\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H})\big{)}/\Gamma_{t},$ (75)
$\bar{\mathbf{D}}^{10}_{n,m}\triangleq\mathbf{D}^{10}[(n-1)N_{r}+1:nN_{r},(m-1)N_{r}+1:mN_{r}],\mathbf{D}^{11}_{l,m,n}\triangleq(\mathbf{u}_{m,0}\mathbf{u}_{n,0}^{H})^{T}\otimes(q_{l}(\mathbf{\Sigma}^{1}_{n,l})^{H}\mathbf{h}^{0,1}_{n,l}\mathbf{h}^{0,1}_{m,l}\Sigma^{1}_{m,l}),$
$\mathbf{d}^{12}_{l,m}\triangleq 2({\sum}_{n\neq m}^{M_{r}}(\mathbf{D}^{11}_{l,m,n})^{H}\hat{\mathbf{h}}^{1}_{n,l}),\mathbf{D}^{14}_{j}\triangleq{\sum}_{k=1}^{K_{d}}[(\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}\hat{\mathbf{g}}_{t,j})^{T}\otimes\sigma^{2}_{t,j}(\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H})]+(\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}^{r}(\hat{\mathbf{W}}^{r})^{H}\hat{\mathbf{g}}_{t,j})^{T}\otimes\sigma^{2}_{t,j}(\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}),$
$\bar{\mathbf{D}}^{14}_{j,i,z}\triangleq\mathbf{D}^{14}_{j}[(i-1)N_{r}+1:iN_{r},(z-1)N_{r}+1:zN_{r}],\mathbf{d}^{13}_{j,m}\triangleq 2({\sum}_{n\neq m}^{M_{r}}({\mathbf{g}}_{r,n,j}^{H}\bar{\mathbf{D}}^{14}_{j,n,m})),\mathbf{d}^{14}_{m}\triangleq 2({\sum}_{n\neq m}^{M_{r}}({\mathbf{g}}_{r,n,0}^{H}\bar{\mathbf{D}}^{10}_{n,m})),$
${\mathbf{D}}^{15}_{l,m,n,i}\triangleq\mathbf{D}^{11}_{l,m,m}[(n-1)L^{t,1}_{m,l}+1:nL^{t,1}_{m,l},(i-1)L^{t,1}_{m,l}+1:iL^{t,1}_{m,l}],\mathbf{d}^{15}_{l,m,n}\triangleq 2({\sum}_{j\neq n}^{N_{r}}({\mathbf{D}}^{15}_{l,m,j,n})^{H}\mathbf{h}^{1}_{m,l,j}),$
$\mathbf{d}^{16}_{l,m,n}\!\triangleq\!\mathbf{d}^{15}_{l,m,n}\!+\!\mathbf{d}^{12}_{l,m}[(n\!-\!1)L^{t,1}_{m,l}\!+\!1\!:\!nL^{t,1}_{m,l}],d^{17}_{j,m,n}\!\triangleq\!2({\sum}_{i\neq n}^{N_{r}}(\bar{\mathbf{D}}^{14}_{j,m,m}[i,n])^{\ast}{\mathbf{g}}_{r,m,j}[i])\!+\!\mathbf{d}^{13}_{j,m}[n],$
$d^{19}_{m,n}\!\triangleq\!2({\sum}_{i\neq n}^{N_{r}}(\bar{\mathbf{D}}^{10}_{m,m}[i,n])^{\ast}{\mathbf{g}}_{r,m,0}[i])\!+\!\mathbf{d}^{14}_{m}[n].$
Therefore, the variable $\mathbf{t}^{1}_{m,n}$ optimization reduces to solving the following problem
[…]
$+\textrm{Re}\{(\mathbf{d}^{9}_{m,l,n})^{H}\mathbf{h}^{1}_{m,l,n}\}\big{)}+c_{5,1,m,n}$
$+{\sum}_{j=0}^{K_{t}}\textrm{Re}\{(d^{11}_{j,m,n})^{\ast}\mathbf{g}_{r,m,j}[n]\}$
s.t. ${\sum}_{l=1}^{K_{u}}\big{(}(\mathbf{h}^{1}_{m,l,n})^{H}\mathbf{D}^{15}_{l,m,n,n}\mathbf{h}^{1}_{m,l,n}$ (76b)
$\!+\!\!\textrm{Re}\!\{\!(\mathbf{d}^{16}_{l,m,n})^{H}\!\mathbf{h}^{1}_{m,l,n}\}\!\big{)}\!\!+\!\!\!{\sum}_{j=1}^{K_{t}}\!\big{(}\!\textrm{Re}\{\!(d^{17}_{j,m,n})^{\ast}\!\mathbf{g}_{r,m,j}[n]\}\!\big{)}$
$-\textrm{Re}\{(d^{19}_{m,n})^{\ast}\mathbf{g}_{r,m,0}[n]\}+c_{5,2,m,n}\leq 0,$
$\|\mathbf{t}^{1}_{m,n}-\mathbf{t}^{1}_{m,i}\|_{2}\geq D_{t},\forall n,i\in\mathcal{N}_{t},n\neq i,$ (76c)
$\mathbf{t}^{1}_{m,n}\in\mathcal{C}.$ (76d)
It is obviously that the above problem is intractable. Next, we can use the MM framework to solve it. However, due to the structural similarity between problems (P9) and (P11), the convexity arguments utilized in (P9) are equally applicable to (P11), e.g., the convexification procedures for the objective function (76a) can be referenced in (61) and (64). Therefore, the specifics of these convexity operations, as detailed in Sec. III-G, are omitted here to avoid redundancy. And then, the optimization problem (P11) is rewritten as
[…]
s.t. $\tau_{2,m,n}\|\mathbf{t}^{1}_{m,n}\|^{2}_{2}\!+\!(\mathbf{d}^{21}_{m,n})^{T}\mathbf{t}^{1}_{m,n}\!+\!c_{5,4,m,n}\leq 0,$ (77b)
$\frac{(\mathbf{t}^{1}_{m,n,0}\!\!-\!\mathbf{t}^{1}_{m,i})^{T}}{\|\mathbf{t}^{1}_{m,n,0}\!\!-\!\mathbf{t}^{1}_{m,i}\|_{2}}\!(\mathbf{t}^{1}_{m,n}\!\!\!-\!\!\mathbf{t}^{1}_{m,i})\!\!\geq\!\!D_{t},\!\forall\!n,\!i\!\!\in\!\!\mathcal{N}_{r},n\!\!\neq\!\!i,$ (77c)
$\mathbf{t}^{1}_{m,n}\in\mathcal{C},$ (77d)
where $c_{5,3,m,n}$ and $c_{5,4,m,n}$ are constant terms, $\mathbf{t}^{1}_{m,n,0}$ is derived from the previous iteration, $\tau_{1,m,n}$ and $\tau_{2,m,n}$ are all positive real numbers and satisfy the condition in Lemma 1, and additional coefficients are omitted here for brevity (refer to Sec. III-G). The optimal solution of problem (P12) can be obtained by CVX.
III-I Optimizing DL User MA’s Position
In this subsection, we will discuss the optimization of the MA position $\mathbf{r}_{0,k}$ for the $k$-th DL user. With given other variables, the problem w.r.t. $\mathbf{r}_{0,k}$ is formulated as
[…]
$+{\sum}_{m=1}^{M_{t}}((\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{Q}^{2}_{k,m}\mathbf{h}^{0,0}_{m,k}\!-\!\!\textrm{ Re }\{(\mathbf{q}^{1}_{k,m})^{H}\mathbf{h}^{0,0}_{m,k}\})$
$+{\sum}_{l=1}^{K_{u}}(\mathbf{h}^{3}_{k,l})^{H}\mathbf{Q}^{3}_{k,l}\mathbf{h}^{3}_{k,l}+c_{6,1,k},$
with $c_{6,1,k}$ being constant and other new coefficients are given as follows
$\mathbf{q}^{1}_{k,m}\triangleq 2\omega_{d,k}\beta_{d,k}^{\ast}\mathbf{\Sigma}^{0}_{m,k}\mathbf{H}^{0}_{m,k}\mathbf{w}_{m,k},$ (79)
$\mathbf{Q}^{3}_{k,l}\triangleq\omega_{d,k}|\beta_{d,k}|^{2}q_{l}\mathbf{\Sigma}^{2}_{k,l}\mathbf{h}^{2}_{k,l}(\mathbf{h}^{2}_{k,l})^{H}(\mathbf{\Sigma}^{2}_{k,l})^{H},$
$\mathbf{Q}^{2}_{k,m}\triangleq\omega_{d,k}|\beta_{d,k}|^{2}\mathbf{\Sigma}^{0}_{m,k}\mathbf{H}^{0}_{m,k}\mathbf{W}^{r}_{m}\mathbf{W}^{r}_{m}(\mathbf{H}^{0}_{m,k})^{H}(\mathbf{\Sigma}^{0}_{m,k})^{H},$
$\mathbf{Q}^{1}_{k,m,n}\!\triangleq\!\!\!{\sum}_{i=1}^{K_{d}}\!\omega_{d,k}|\beta_{d,k}|^{2}\mathbf{\Sigma}^{0}_{m,k}\mathbf{H}^{0}_{m,k}\mathbf{w}_{m,i}\mathbf{w}_{n,i}(\mathbf{H}^{0}_{n,k})^{\!H}\!\!(\mathbf{\Sigma}^{0}_{n,k})^{H}.$
Obviously, obtaining the optimal solution to problem (P13) is challenging due to the non-convex nature of the objective function (78a). Therefore, we still employ the MM methodology to address it.
Firstly, the term $(\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{Q}^{1}_{k,m,n}\mathbf{h}^{0,0}_{n,k}$ in (78a) can be upper-bounded by invoking (63) in Lemma 1 as shown in (80), where $c_{6,2,k,m,n,i,j}$ is constant, $\mathbf{q}^{3}_{k,m,n,i,j}\triangleq\nabla h^{R1}_{k,m,n,i,j}-\psi_{1,k,m,n,i,j}\mathbf{r}_{0,k,0}$, $\mathbf{r}_{0,k,0}=[x^{d}_{k,0},y^{d}_{k,0}]^{T}$ is obtained form the last iteration, $\nabla h^{R1}_{k,m,n,i,j}$ is the the gradient vector of $h^{R1}_{k,m,n,i,j}$ over $\mathbf{r}_{0,k,0}$, $\psi_{1,k,m,n,i,j}=\frac{32\pi^{2}}{\lambda^{2}}|\mathbf{Q}^{1}_{k,m,n}[i,j]|$, and the derivations of $\nabla h^{R1}_{k,m,n,i,j}$ and $\psi_{1,k,m,n,i,j}$ are presented in Appendix A.
$(\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{Q}^{4}_{k,m,n}\mathbf{h}^{0,0}_{n,k}={\sum}_{i=1}^{L^{r,0}_{m,k}}{\sum}_{j=1}^{L^{r,0}_{n,k}}(\mathbf{h}^{0,0}_{m,k}[i])^{\ast}\mathbf{Q}^{4}_{k,m,n}[i,j]\mathbf{h}^{0,0}_{m,k}[j]$ (80)
$={\sum}_{i=1}^{L^{r,0}_{m,k}}{\sum}_{j=1}^{L^{r,0}_{n,k}}\underbrace{|\mathbf{Q}^{4}_{k,m,n}[i,j]|\cos(\frac{2\pi}{\lambda}(\mathbf{r}_{0,k})^{T}(-\mathbf{a}^{r,0}_{m,k,i}+\mathbf{a}^{r,0}_{n,k,j})+\angle\mathbf{Q}^{4}_{k,m,n}[i,j])}\limits_{h^{R1}_{k,m,n,i,j}}$
$\overset{(a)}{\leq}{\sum}_{i=1}^{L^{r,0}_{m,k}}{\sum}_{j=1}^{L^{r,0}_{n,k}}\frac{\psi_{1,k,m,n,i,j}}{2}\|\mathbf{r}_{0,k}\|^{2}_{2}+(\mathbf{q}^{3}_{k,m,n,i,j})^{T}\mathbf{r}_{0,k}+c_{6,2,k,m,n,i,j},$
Furthermore, the terms $(\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{Q}^{2}_{k,m}\mathbf{h}^{0,0}_{m,k}-\textrm{ Re }\{(\mathbf{q}^{1}_{k,m})^{H}\mathbf{h}^{0,0}_{m,k}\}$ and $(\mathbf{h}^{3}_{k,l})^{H}\mathbf{Q}^{3}_{k,l}\mathbf{h}^{3}_{k,l}$ all can be similarly convexified following the same arguments as in (61) and (64) in Sec III-G. Therefore, the upper-bounds of $(\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{Q}^{2}_{k,m}\mathbf{h}^{0,0}_{m,k}-\textrm{ Re }\{(\mathbf{q}^{1}_{k,m})^{H}\mathbf{h}^{0,0}_{m,k}\}$ and $(\mathbf{h}^{3}_{k,l})^{H}\mathbf{Q}^{3}_{k,l}\mathbf{h}^{3}_{k,l}$ are shown in (81) and (82), respectively, in the following
$(\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{Q}^{2}_{k,m}\mathbf{h}^{0,0}_{m,k}-\textrm{ Re }\{(\mathbf{q}^{1}_{k,m})^{H}\mathbf{h}^{0,0}_{m,k}\}$ (81)
$\leq\frac{\psi_{2,k,m}}{2}\|\mathbf{r}_{0,k}\|^{2}_{2}+(\mathbf{q}^{3}_{k,m})^{T}\mathbf{r}_{0,k}+c_{6,2,k,m},$
$(\mathbf{h}^{3}_{k,l})^{H}\mathbf{Q}^{3}_{k,l}\mathbf{h}^{3}_{k,l}\leq\frac{\psi_{3,k,l}}{2}\|\mathbf{r}_{0,k}\|^{2}_{2}+(\mathbf{q}^{4}_{k,l})^{T}\mathbf{r}_{0,k}+c_{6,3,k,l},$ (82)
with $c_{6,2,k,m}$ and $c_{6,3,k,l}$ being constants, the numbers $\psi_{2,k,m}$ and $\psi_{3,k,l}$ adhere to the condition outlined in Lemma 1.
In accordance with the aforementioned MM transformation, the update of $\mathbf{r}_{0,k}$ is achieved by solving a convex surrogate of the original problem (P13), as follows
[…]
where $c_{6,4,k}$ is constant, $\psi_{4,k}\!=\!0.5({\sum}_{m=1}^{M_{t}}\psi_{2,k,m}\!+\!\!{\sum}_{m=1}^{M_{t}}{\sum}_{n=1}^{M_{t}}{\sum}_{i=1}^{L^{r,0}_{m,k}}{\sum}_{j=1}^{L^{r,0}_{n,k}}\psi_{1,k,m,n,i,j}\!+\!\!{\sum}_{l=1}^{K_{u}}\psi_{3,k,l})$ and $\mathbf{q}^{5}_{k}\!=\!{\sum}_{m=1}^{M_{t}}{\sum}_{n=1}^{M_{t}}{\sum}_{i=1}^{L^{r,0}_{m,k}}{\sum}_{j=1}^{L^{r,0}_{n,k}}\mathbf{q}^{2}_{k,m,n,i,j}\!+\!\!{\sum}_{m=1}^{M_{t}}\mathbf{q}^{3}_{k,m}\!+\!\!{\sum}_{l=1}^{K_{u}}\mathbf{q}^{4}_{k,l}$. The problem (P14) is convex and can be solved by CVX.
Since the units of $\mathbf{r}_{0,k}$, i.e., $x^{d}_{k}$ and $y^{d}_{k}$, are not coupled, the problem w.r.t. $x^{d}_{k}$ can be written as
[…]
s.t. $-{A}/{2}\leq x^{d}_{k}\leq{A}/{2},$ (84b)
Since $\psi_{4,k}>0$, we can obtain the optimal solution $(x^{d}_{k})^{\star}$ by directly determining the position of the axis of symmetry of the function (84a), which are given as
$(x^{d}_{k})^{\star}\!=\!\left\{\begin{aligned} -{A}/{2},\ &\textrm{if}\ -{\mathbf{q}^{5}_{k}[1]}/({2\psi_{4,k}})<-{A}/{2},\\ -{\mathbf{q}^{5}_{k}[1]}/{2\psi_{4,k}},\ &\textrm{if}\ -{A}/{2}\leq-{\mathbf{q}^{5}_{k}[1]}/{2\psi_{4,k}}\leq{A}/{2},\\ {A}/{2},\ &\textrm{if}\ -{A}/{2}<-{\mathbf{q}^{5}_{k}[1]}/{2\psi_{4,k}}.\end{aligned}\right.$ (85)
Similarly, the optimal solution $(y^{d}_{k})^{\star}$ can be directly formulated as
$(y^{d}_{k})^{\star}\!=\!\left\{\begin{aligned} -{A}/{2},\ &\textrm{if}\ -{\mathbf{q}^{5}_{k}[2]}/({2\psi_{4,k}})<-{A}/{2},\\ -{\mathbf{q}^{5}_{k}[1]}/{2\psi_{4,k}},\ &\textrm{if}\ -{A}/{2}\leq-{\mathbf{q}^{5}_{k}[2]}/{2\psi_{4,k}}\leq{A}/{2},\\ {A}/{2},\ &\textrm{if}\ -{A}/{2}<-{\mathbf{q}^{5}_{k}[2]}/{2\psi_{4,k}}.\end{aligned}\right.$ (86)
III-J Optimizing UL User MA’s Position
When other variables are obtained, the optimization problem solving for the MA’s position ${\mathbf{r}_{1,l}}$ of the $l$-th UL user is formulated as
[…]
$+{\sum}_{m=1}^{M_{r}}{\sum}_{n=1}^{M_{r}}(\mathbf{h}^{0,1}_{m,l})^{H}\mathbf{Q}^{5}_{l,m,n}\mathbf{h}^{0,1}_{n,l}+c_{7,1,l}$ (87a)
s.t. ${\sum}_{m=1}^{M_{r}}{\sum}_{n=1}^{M_{r}}(\mathbf{h}^{0,1}_{m,l})^{H}\mathbf{Q}^{6}_{l,m,n}\mathbf{h}^{0,1}_{n,l}\!+\!c_{7,2,l}\leq 0,$ (87b)
$\mathbf{r}_{1,l}\in\mathcal{C},$ (87c)
with $c_{7,1,l}$ and $c_{7,2,l}$ being constant terms, and we define
$\mathbf{q}^{6}_{l,m}\triangleq 2\omega_{d,k}\beta_{d,k}\mathbf{\Sigma}^{1}_{m,l}\mathbf{H}^{1}_{m,i}\mathbf{u}_{m,l},$ (88)
$\mathbf{Q}^{4}_{k,l}\triangleq\omega_{d,k}|\beta_{d,k}|^{2}q_{l}(\mathbf{\Sigma}^{2}_{k,l})^{H}\mathbf{h}^{3}_{k,l}(\mathbf{h}^{3}_{k,l})^{H}\mathbf{\Sigma}^{2}_{k,l},$
$\mathbf{Q}^{6}_{l,m,n}\triangleq\mathbf{\Sigma}^{1}_{m,i}\mathbf{H}^{1}_{m,i}\mathbf{u}_{m,0}\mathbf{u}_{n,0}^{H}(\mathbf{H}^{1}_{n,i})^{H}(\mathbf{\Sigma}^{1}_{n,i})^{H},$
${\mathbf{Q}}^{5}_{i,m,n}\!\triangleq\!\!{\sum}_{l=1}^{K_{u}}\omega_{d,k}|\beta_{d,k}|^{2}\mathbf{\Sigma}^{1}_{m,i}\mathbf{H}^{1}_{m,i}\mathbf{u}_{m,l}\mathbf{u}_{n,l}^{H}(\mathbf{H}^{1}_{n,i})^{H}\!(\mathbf{\Sigma}^{1}_{n,i})^{H}.$
Unfortunately, this problem is still difficult to solve due to the non-convex objective (87a) and constraint (87b) w.r.t $\mathbf{r}_{1,l}$. To address these issues, we will construct the convex surrogate functions of (87a) and (87b) by applying MM method, and then recast the problem (P16) into an explicit form of $\mathbf{r}_{1,l}$.
Firstly, by using the same MM-based algorithmic framework in (80), the convex upper-bounds of the terms $(\mathbf{h}^{0,1}_{m,l})^{H}\mathbf{Q}^{5}_{l,m,n}\mathbf{h}^{0,1}_{n,l}$ and $(\mathbf{h}^{0,1}_{m,l})^{H}\mathbf{Q}^{6}_{l,m,n}\mathbf{h}^{0,1}_{n,l}$ are respectively provided as
$(\mathbf{h}^{0,1}_{m,l})^{H}\mathbf{Q}^{5}_{l,m,n}\mathbf{h}^{0,1}_{n,l}\leq{\sum}_{i=1}^{L^{r,1}_{m,l}}{\sum}_{j=1}^{L^{r,1}_{n,l}}\frac{\psi_{6,l,m,n,i,j}}{2}\|\mathbf{r}_{1,l}\|^{2}_{2}$ (89)
$+(\mathbf{q}^{8}_{l,m,n,i,j})^{T}\mathbf{r}_{1,l}+c_{7,3,l,m,n,i,j},$
$(\mathbf{h}^{0,1}_{m,l})^{H}\mathbf{Q}^{6}_{l,m,n}\mathbf{h}^{0,1}_{n,l}\leq{\sum}_{i=1}^{L^{r,1}_{m,l}}{\sum}_{j=1}^{L^{r,1}_{n,l}}\frac{\psi_{9,l,m,n,i,j}}{2}\|\mathbf{r}_{1,l}\|^{2}_{2}$ (90)
$+(\mathbf{q}^{10}_{l,m,n,i,j})^{T}\mathbf{r}_{1,l}+c_{7,4,l,m,n,i,j},$
with $c_{7,3,l,m,n,i,j}$ and $c_{7,4,l,m,n,i,j}$ being constant, $\psi_{6,l,m,n,i,j}=\frac{32\pi^{2}}{\lambda^{2}}|\mathbf{Q}^{10}_{l,m,n}[i,j]|$ and $\psi_{9,l,m,n,i,j}=\frac{32\pi^{2}}{\lambda^{2}}|\mathbf{Q}^{11}_{l,m,n}[i,j]|$.
Furthermore, we can construct an upper-bound for $(\mathbf{h}^{2}_{k,l})^{H}\mathbf{Q}^{4}_{k,l}\mathbf{h}^{2}_{k,l}$ using the same operations as in (61) and (64), as follows
$(\mathbf{h}^{2}_{k,l})^{H}\mathbf{Q}^{4}_{k,l}\mathbf{h}^{2}_{k,l}\leq\frac{\psi_{5,k,l}}{2}\|\mathbf{r}_{1,l}\|^{2}_{2}+(\mathbf{q}^{7}_{k,l})^{T}\mathbf{r}_{1,l}+c_{7,5,k,l},$ (91)
where $c_{7,5,k,l}$ is constant, and $\psi_{5,k,l}$ and $\mathbf{q}^{7}_{k,l}$ are derived using similar arguments as presented in in (61) and (64).
Subsequently, referring to (68), a convex surrogate function for the term $\textrm{Re}\{\mathbf{q}^{6}_{l,m}\mathbf{h}^{0,1}_{m,l}\}$ can be derived as follows
$\textrm{Re}\{\mathbf{q}^{6}_{l,m}\mathbf{h}^{0,1}_{m,l}\}\!\geq\!-\frac{\psi_{7,l,m}}{2}\|\mathbf{r}_{1,l}\|^{2}_{2}+(\mathbf{q}^{9}_{l,m})^{T}\mathbf{r}_{1,l}+c_{7,6,l,m},$ (92)
where $c_{7,6,l,m}$ denotes a constant term, and $\psi_{7,l,m}$ and $\mathbf{q}^{9}_{l,m}$ can be obtained leveraging the results presented in (68).
Substituting the result in (89)-(92) into the original problem (P16), we further obtain
[…]
s.t. $\psi_{10,l}\|\mathbf{r}_{1,l}\|^{2}_{2}+(\mathbf{q}^{12}_{l})^{T}\mathbf{r}_{1,l}+c_{7,8,l}\leq 0,$ (93b)
$\mathbf{r}_{1,l}\in\mathcal{C},$ (93c)
where $c_{7,7,l}$ and $c_{7,8,l}$ are constants, $\psi_{8,l}\triangleq{\sum}_{k=1}^{K_{d}}\frac{\psi_{5,k,l}}{2}+{\sum}_{m=1}^{M_{r}}{\sum}_{n=1}^{M_{r}}{\sum}_{i=1}^{L^{r,1}_{m,l}}{\sum}_{j=1}^{L^{r,1}_{n,l}}\frac{\psi_{6,l,m,n,i,j}}{2}+{\sum}_{m=1}^{M_{r}}\frac{\psi_{7,l,m}}{2}$, $\mathbf{q}^{10}_{l}\triangleq{\sum}_{k=1}^{K_{d}}\mathbf{q}^{7}_{k,l}+{\sum}_{m=1}^{M_{r}}{\sum}_{n=1}^{M_{r}}{\sum}_{i=1}^{L^{r,1}_{m,l}}{\sum}_{j=1}^{L^{r,1}_{n,l}}\mathbf{q}^{8}_{l,m,n,i,j}-{\sum}_{m=1}^{M_{r}}\mathbf{q}^{9}_{l,m}$, $\psi_{10,l}\triangleq{\sum}_{m=1}^{M_{r}}{\sum}_{n=1}^{M_{r}}{\sum}_{i=1}^{L^{r,1}_{m,l}}{\sum}_{j=1}^{L^{r,1}_{n,l}}\frac{\psi_{9,l,m,n,i,j}}{2}$ and $\mathbf{q}^{12}_{l}\!\!\triangleq\!\!{\sum}_{m=1}^{M_{r}}{\sum}_{n=1}^{M_{r}}{\sum}_{i=1}^{L^{r,1}_{m,l}}{\sum}_{j=1}^{L^{r,1}_{n,l}}\mathbf{q}^{11}_{l,m,n,i,j}$. Now, the problem (P17) is convex that can be solved optimally by CVX.
Moreover, the overall algorithm to solve (P1) is specified in Algorithm 1.
[…]
1:  initialize $i=0$;
2:  randomly generate feasible $\{\mathbf{w}^{0}_{m,k},(\mathbf{W}^{r}_{m})^{0}\}$, $\{\mathbf{u}_{p,j}^{0}\}$$\{q_{l}\}$$\{(\mathbf{t}^{0}_{m,n})^{0},(\mathbf{t}^{1}_{m,n})^{0}\}$ and $\{\mathbf{r}_{0,k}^{0},\mathbf{r}_{1,l}^{0}\}$;
3:  repeat
4:     update $\{\omega_{d,k}\}$, $\{\beta_{d,k}\}$, $\{\omega_{u,l}\}$ and $\{\beta_{u,l}\}$, by (32), (31), (34) and (33), respectively;
5:     update $\{\mathbf{w}^{i+1}_{m,k},(\mathbf{W}^{r}_{m})^{i+1}\}$ by solving (P3);
6:     update $\{\mathbf{u}_{p,l}^{i+1}\},\forall l\in\mathcal{K}_{u}$ by equation (49);
7:     update $\{\mathbf{u}_{p,0}^{i+1}\}$ by solving (P8);
8:     update $\{(\mathbf{t}^{0}_{m,n})^{i+1}\}$ by solving (P10);
9:     update $\{(\mathbf{t}^{1}_{m,n})^{i+1}\}$ by solving (P12);
10:     update $\{\mathbf{r}_{0,k}^{i+1}\}$ by equations (85) and (86);
11:     update $\{\mathbf{r}_{1,l}^{i+1}\}$ by solving (P17);
12:     $i++$;
13:  until $convergence$
IV Numerical Results
Figure 2: The simulation setup.
In this section, we present numerical results to validate our proposals. The experimental setup is illustrated in Fig. 2, featuring multiple TBS and RBS units jointly attempting to detect a target amidst clutters while simultaneously respectively serving two DL and two UL users. In the experiment, TBSs, RBSs, DL users, UL users, one target and clutters are randomly located within a circular area with a radius of $70$ meters (m). We employ a geometric channel model [4] for the link concerning each user, wherein the number of transmit and receive channel paths is consistently identical, i.e., $L^{t,0}_{m,k}=L^{r,0}_{m,k}=L^{t,1}_{p,l}=L^{r,1}_{p,l}=L^{t,2}_{k,l}=L^{r,2}_{k,l}\triangleq L,\forall m\in\mathcal{M}_{t},\forall p\in\mathcal{M}_{r},\forall k\in\mathcal{K}_{d},\forall l\in\mathcal{K}_{u}$. Thus, the path-response matrix for each user is diagonal, i.e., the path-response matrix for between the $m$-th TBS and the $k$-th DL user is represented as $\mathbf{\Sigma}^{0}_{m,k}=\textrm{diag}\{\sigma^{0}_{m,k,1},\dots,\sigma^{0}_{m,k,L}\}$, where each $\mathbf{\sigma}^{0}_{m,k,l}$ satisfying $\sigma^{0}_{m,k,l}\sim\mathcal{CN}(0,{(c^{0}_{m,k})^{2}}/{L}),l=1,\dots,L$ [4], and $(c^{0}_{m,k})^{2}=C_{0}d_{m,k}^{-\alpha_{loss}}$, where $C_{0}$ corresponds to the path loss at the reference distance of $1$ m, $d_{m,k}$ is the propagation distance between the $m$-th TBS and the $k$-th DL user, $\alpha_{loss}=2.8$ is the path-loss exponent. The elevation and azimuth of both AoAs and AoDs are assumed to be independent and identically distributed variables following the uniform distribution over $[-\frac{\pi}{2},\frac{\pi}{2}]$. The TBS-target/clutters and RBS-target/clutters are modeled as line-of-sight (LoS) channels. The moving regions for MAs at both the BS and users are set as square areas of size $[-\frac{A}{2},\frac{A}{2}]\times[-\frac{A}{2},\frac{A}{2}]$. In addition, $M_{t}=2$ TBS equip $N_{t}=4$ antennas and $M_{r}=2$ RBS equip $N_{r}=4$ antennas. The transmit power of each TBS is set as $30$dBm. The noise power and the predefined target detection level of RBS are set as $\sigma^{2}_{d,k}=\sigma^{2}_{r}=-80$dBm and $\Gamma_{r}=3$dB, respectively.
Figure 3: Convergence of Alg. 1.
Fig. 3 examines the convergence behaviour of the overall solution in Alg. 1. The obtained sun-rate iterates under varying configurations of $N_{t}$ and $N_{r}$ are illustrated in the plot. As shown in Fig. 3, the proposed algorithm can yield monotonic improvement in sun-rate and generally achieves significant beamforming gain within the first 10 iterations.
Figure 4: The impact of TBS transmit power.
Figure 5: The impact of TBS antenna number.
Figure 6: The impact of RBS antenna number.
Alg. 1 is labeled by “Joint MA”. For comparison in sum-rate enhancement, we consider the following three foundational scenarios: 1) BS MA: All BSs are equipped with the MAs, while all users respectively employ an MA fixing at the original point. 2) User MA: All users respectively employ an MA, while the MAs of all BSs are fixed. 3) FPA: Both all BSs and users are equipped with FPA-based array.
Fig. 6 compares the sum-rate for different schemes versus the transmit power of all TBS. The sum-rate of all schemes increases significantly with the transmit power and the three MA-based schemes significantly improve the sum-rate compared to the FPA-based case. In particular, the “Joint MA” case shows a notable advantage over all other cases in terms of sum-rate, demonstrating performance improvement of approximately 20.0% over the FPA scheme. This is anticipated, as the proposed “Joint MA” scheme exploits the most spatial DoFs, and the other three schemes face performance loss due to less flexibility in channel reconfiguration.
Fig. 6 illustrates the impact of the number of antenna at the TBSs. The left and right subplots corresponding to different numbers of TBSs, respectively. In our test, the number of antenna $N_{t}$ varies from 2 to 9. As reflected by Fig. 6, increasing $N_{t}$ can improve beamforming gain for all schemes. Moreover, the “Joint MA” scheme boosts the system’s sum-rate more significantly compared to all other schemes. Fig. 6 shows the impact of the number of antenna at the RBSs. As can be seen in both left and right subplots, which corresponding to varying quantities of RBSs, sum-rate increases when $N_{r}$ increases. Compared to other foundational scenarios, the deployment of MA in both BS and user can significantly boot the sum-rate.
Figure 7: The impact of the number of paths for DL user.
Figure 8: The impact of the number of paths for UL user.
Figure 9: The impact of the normalized region size.
Fig. 9 explores the influence of the number of paths between TBSs and DL users on the sum-rate. It can be easily observed that a larger number of paths results in a higher achievable sum-rate across all schemes, and our proposed algorithm demonstrates a superior performance over other schemes. Besides, when $K_{d}=6$, all schemes exhibit a higher sum-rate compared to the case when $K_{d}=3$. Fig. 9 compares the sum-rate for different schemes versus the number of paths between RBSs and UL users. It is readily apparent that an increased number of paths leads to a higher achievable sum-rate in all schemes, with the “Joint MA” scheme exhibiting superior performance compared to others. In addition, the $K_{u}=6$ case can achieve better performance than the $K_{u}=3$ case.
Fig. 9 describes the sum-rate versus the normalized region size $A/\lambda$ for MA at the BSs. As can be observed, both the “Joint MA” and “BS MA” schemes experience increases as the normalized region size grows, attributed to the larger moving area of BSs’ MA providing more spatial DoFs. Besides, we observe that the increase of the normalized region size will enlarge the performance gaps between “Joint MA” and “BS MA” schemes. Since both BSs and users are equipped with the MA, the “Joint MA” case can provide more effective channel reconfiguration.
V Conclusions
This paper considers the joint active beamforming and position coefficients design problem in an MA aided networked FD ISAC system that accomplishes radar sensing as well as DL and UL communication capabilities concurrently. We propose an iterative algorithm to jointly optimize BS beamforming, UL users’ power allocation, receiving processors and MA configuration for enabling radar sensing and communication functionalities. Numerical results showcase the effectiveness of our proposed solution and highlight the advantages of deploying MA in the networked FD ISAC system.
-A Derivations of $\nabla h^{R1}_{k,m,n,i,j}$ and $\psi_{1,k,m,n,i,j}$
Firstly, the elements of $\nabla h^{R1}_{k,m,n,i,j}(\mathbf{r}_{0,k,0})=\big{[}\frac{\partial\nabla h^{R1}_{k,m,n,i,j}(\mathbf{r}_{0,k,0})}{\partial x^{d}_{k,0}},\frac{\partial\nabla h^{R1}_{k,m,n,i,j}(\mathbf{r}_{0,k,0})}{\partial y^{d}_{k,0}}\big{]}^{T}$ are given as
$\frac{\partial h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial x^{d}_{k,0}}\!\!\!=\!\!-\!\frac{2\pi}{\lambda}\!|\!\mathbf{Q}^{1}_{k,m,n}\![i,j]\!|\!\big{(}\!\!\!-\!\cos(\!\theta^{r,0}_{m,k,i}\!)\!\sin(\!\phi^{r,0}_{m,k,i}\!)$ (94)
$\!\!+\!\!\cos(\!\theta^{r,0}_{n,k,j}\!)\!\sin(\!\phi^{r,0}_{n,k,j}\!)\!\big{)}\!\!\sin\!\!\big{(}\!\frac{2\pi}{\lambda}\!(\!\mathbf{r}_{0,k,0}\!)^{\!T}\!\!(\!\mathbf{a}^{r,0}_{n,k,j}\!\!\!-\!\mathbf{a}^{r,0}_{m,k,i}\!)\!\!+\!\!\!\angle\!\mathbf{Q}^{1}_{k,m,n}\![i,\!j]\!\big{)},$
$\frac{\partial h^{R1}_{k,m,n,i,j}(\mathbf{r}_{0,k,0})}{\partial y^{d}_{k,0}}=-\!\frac{2\pi}{\lambda}|\mathbf{Q}^{1}_{k,m,n}[i,j]|(-\sin(\theta^{r,0}_{m,k,i})$ (95)
$+\sin(\theta^{r,0}_{n,k,j}))\sin\!\big{(}\!\frac{2\pi}{\lambda}\!(\!\mathbf{r}_{0,k,0}\!)^{T}\!(\!-\!\mathbf{a}^{r,0}_{m,k,i}\!+\!\mathbf{a}^{r,0}_{n,k,j})\!+\!\angle\!\mathbf{Q}^{1}_{k,m,n}[i,j]\big{)}.$
To obtain $\psi_{1,k,m,n,i,j}$, the Hessian matrix of $\nabla^{2}h^{R1}_{k,m,n,i,j}(\mathbf{r}_{0,k,0})$ over $\mathbf{r}_{0,k,0}$ can be formulated as
$\nabla^{2}\!h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)\!\!\triangleq\!\!\!\!\begin{bmatrix}\frac{\partial^{2}h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial x^{d}_{k,0}\partial x^{d}_{k,0}}&\!\!\!\!\!\frac{\partial^{2}h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial x^{d}_{k,0}\partial y^{d}_{k,0}}\\ \frac{\partial^{2}h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial y^{d}_{k,0}\partial x^{d}_{k,0}}&\!\!\!\!\!\frac{\partial^{2}h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial y^{d}_{k,0}\partial y^{d}_{k,0}}\end{bmatrix},$ (96)
where the matrix units are respectively given as follows
$\frac{\partial^{2}\!h^{R1}_{k,m,n,i,j}\!(\!\mathbf{r}_{0,k,0}\!\!)}{\partial x^{d}_{k,0}\partial x^{d}_{k,0}}\!\!\!=\!\!-\!\frac{4\pi^{2}}{\lambda^{2}}|\!\mathbf{Q}^{1}_{k,m,n}[i,j]\!|\!\big{(}\!\!\!-\!\!\cos\!(\!\theta^{r,0}_{m,k,i}\!)\!\sin(\!\phi^{r,0}_{m,k,i}\!)$ (97)
$\!\!+\!\!\cos(\!\theta^{r,0}_{n,k,j}\!)\!\sin(\!\phi^{r,0}_{n,k,j}\!)\!\big{)}^{\!2}\!\!\!\cos\!\!\big{(}\!\frac{2\pi}{\lambda}\!(\!\mathbf{r}_{0,k,0}\!)^{\!T}\!\!(\!\mathbf{a}^{r,0}_{n,k,j}\!\!\!-\!\mathbf{a}^{r,0}_{m,k,i}\!)\!\!+\!\!\!\angle\!\mathbf{Q}^{1}_{k,m,n}\![i,\!j]\!\big{)},$
$\frac{\partial^{2}\!h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial x^{d}_{k,0}\partial y^{d}_{k,0}}\!\!=\!\!\frac{\partial^{2}\!h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial y^{d}_{k,0}\partial x^{d}_{k,0}}\!\!\!=\!\!-\!\frac{4\pi^{2}}{\lambda^{2}}|\mathbf{Q}^{1}_{k,m,n}[i,j]|$ (98)
$\big{(}\!\!\!-\!\cos(\!\theta^{r,0}_{m,k,i}\!)\!\sin(\!\phi^{r,0}_{m,k,i}\!)\!\!+\!\!\cos(\!\theta^{r,0}_{n,k,j}\!)\sin(\!\phi^{r,0}_{n,k,j}\!)\!(-\sin(\theta^{r,0}_{m,k,i})$
$+\sin(\theta^{r,0}_{n,k,j}))\cos\!\!\big{(}\!\frac{2\pi}{\lambda}\!(\!\mathbf{r}_{0,k,0}\!)^{T}\!(\!-\!\mathbf{a}^{r,0}_{m,k,i}\!\!+\!\mathbf{a}^{r,0}_{n,k,j}\!)\!\!+\!\!\angle\!\mathbf{Q}^{1}_{k,m,n}[i,j]\!\big{)}\big{)},$
$\frac{\partial^{2}h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial y^{d}_{k,0}\partial y^{d}_{k,0}}=-\!\frac{4\pi^{2}}{\lambda^{2}}|\mathbf{Q}^{1}_{k,m,n}[i,j]|(-\sin(\theta^{r,0}_{m,k,i})$ (99)
$+\sin(\theta^{r,0}_{n,k,j}))^{2}\cos\!\!\big{(}\!\frac{2\pi}{\lambda}\!(\!\mathbf{r}_{0,k,0}\!)^{T}\!(\!-\!\mathbf{a}^{r,0}_{m,k,i}\!\!+\!\mathbf{a}^{r,0}_{n,k,j}\!)\!\!+\!\!\angle\!\mathbf{Q}^{1}_{k,m,n}[i,j]\!\big{)}.$
And then, we have
$\|\nabla^{2}h^{R1}_{k,m,n,i,j}(\mathbf{r}_{0,k,0})\|_{2}\leq\|\nabla^{2}h^{R1}_{k,m,n,i,j}(\mathbf{r}_{0,k,0})\|_{F}$ (100)
$=\bigg{[}\bigg{(}\frac{\partial^{2}h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial x^{d}_{k,0}\partial x^{d}_{k,0}}\bigg{)}^{2}+\bigg{(}\frac{\partial^{2}\!h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial x^{d}_{k,0}\partial y^{d}_{k,0}}\bigg{)}^{2}$
$+\bigg{(}\frac{\partial^{2}\!h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial y^{d}_{k,0}\partial x^{d}_{k,0}}\bigg{)}^{2}+\bigg{(}\frac{\partial^{2}h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial y^{d}_{k,0}\partial y^{d}_{k,0}}\bigg{)}^{2}\bigg{]}^{\frac{1}{2}}$
$\leq\bigg{[}4\bigg{(}\frac{16\pi^{2}}{\lambda^{2}}|\mathbf{Q}^{4}_{k,m,n}[i,j]\bigg{)}^{2}\bigg{]}^{\frac{1}{2}}.$
Therefore, to guarantee that operation (a) in (80) yields an upper-bound for $h^{R1}_{k,m,n,i,j}$, we can define $\psi_{1,k,m,n,i,j}=\frac{32\pi^{2}}{\lambda^{2}}|\mathbf{Q}^{1}_{k,m,n}[i,j]$ with $\psi_{1,k,m,n,i,j}\mathbf{I}_{2}\succeq\nabla^{2}h^{R1}_{k,m,n,i,j}(\mathbf{r}_{0,k,0})$.