### v3

# Movable Antenna Enhanced Networked Integrated Sensing and Communication System

Abstract
Integrated sensing and communication (ISAC) is a key technology for future 6G networks. Most existing studies focus on monostatic and/or bistatic setups with limited coverage and capabilities. Networked ISAC systems with distributed base stations (BSs) can overcome these limitations. Moreover, movable antenna (MA) architectures offer improved ISAC performance over fixed-position antennas (FPAs) by enabling adaptable antenna movement. In this paper, we utilize the MA to promote communication capability with guaranteed sensing performance via jointly designing beamforming, power allocation, receiving filters and position configuration of transmit/receive MA towards maximizing the sum rate for both downlink (DL) and uplink (UL) users. The optimization problem is highly difficult due to the unique channel model derived from the position coefficient of the MA. To resolve this challenge, via leveraging the cutting-the-edge majorization-minimization (MM) method, we develop an efficient solution that optimizes all variables via convex optimization techniques. Extensive simulation results verify the effectiveness of our proposed algorithms and demonstrate the substantial performance promotion by deploying the MA framework in the networked ISAC system.
I Introduction
Recently, with the rapid increase in devices requiring precision sensing and efficient communication, integrated sensing and communication (ISAC) [1]-[2] technologies have attracted great interest from both industry and academia. In this context, ISAC is envisioned as a promising solution, which aims to deploy both communication and sensing functionalities on one unified hardware platform, allowing the sharing of frequency spectrum, hardware resources, and signal-processing units. Many recent advancements in the joint design of radar sensing and communication are extensively documented in [1]-[2] and their associated references.
However, the conventional fixed-position antenna (FPA) at the base station (BS) and/or mobile user in the ISAC system generally fail to fully exploit the degrees of freedom (DoFs) in the continuous spatial domain, leading to performance loss in both sensing and communication tasks. A novel movable antenna (MA) technology [3]-[5] is expected to overcome the limitation. The MA is considered as a viable approach for enhancing wireless network performance, linked to a radio frequency (RF) chain through a flexible cable. It can reconstruct channel conditions by flexibly adjusting its position in a given spatial area with the aid of a driver component or by other means. The various aspects of wireless systems by applying the MA framework have been extensively studied recently, see [6]-[11] and their references.
I-A Related Works
In the mono-static ISAC system, the full-duplex (FD) operation presents a significant challenge due to self-interference [1], which cannot be overlooked. The work [12] aimed to deal with the challenges associated with the FD operation in mono-static ISAC systems operating at millimeter-Wave (mm-Wave) frequencies, and proposed the beamforming solutions designed to maximize the beamformer power in the sensing direction while simultaneously constraining the beamformer power in the communication directions. The authors of [13] and [14] adopted the Orthogonal Frequency Division Multiplexing (OFDM) waveform to estimate the Direction of Arrival (DoA), range, and relative velocity of the radar targets while maximizing the downlink (DL) communication rate in the mono-static FD ISAC network operating at the mm-Wave frequencies. The paper [15] investigated a mono-static FD ISAC system, in which the BS simultaneously conducts target detection while communicating with multiple DL and uplink (UL) users, sharing the same time-frequency resources. The authors of [16] utilized the FD in the ISAC transceiver to solve the echo-miss problem.
The networked ISAC system with multiple transmitters and/or receivers can significantly improve both the communication and radar sensing performance [1]. In [17], the work investigated a networked ISAC system that considers two types of users: those capable of canceling interference from dedicated sensing signals and those that are not. The paper [18] presented a cell-free ISAC MIMO system, deriving sensing signal-to-noise (SNR) and developing joint beamforming and power allocation for optimal communication and sensing performance. The authors of [19] simultaneously considered the active and passive sensing tasks in a downlink distributed ISAC system, adopting different fusion strategies based on the backhaul capacity and enhancing sensing performance while meeting minimum signal-to-interference-plus-noise (SINR) requirements. The work [20] demonstrated a multi-static ISAC design using network-assisted full-duplex (NAFD) cell-free networks, proposing a deep Q-network for optimizing access point (AP) duplex modes.
[…]
I-B Motivations and Contributions
As shown previously, although several studies [12]-[20] have investigated joint beamforming optimization in the ISAC systems, most have relied on traditional FPAs, limiting their adaptability and performance. Since the MA technology can effectively utilize spatial DoFs in practical networks [4]-[5], extensive literature [21]-[33] has explored the joint beamforming and MA’s position optimization for MA-aided ISAC systems but primarily focused on monostatic and/or bistatic configurations. These settings have limited service coverage and inadequate performance in both sensing and communication within complex environments. However, the networked ISAC is a promising solution for enhancing coverage, spectral efficiency, and reliability [1]-[2]. Moreover, these works [21]-[29], [33], primarily concentrate on DL communication, overlooking the crucial aspect of UL communication which is essential for comprehensive ISAC functionality. This oversight can restrict the practical application of existing systems in dynamic environments where bidirectional communication is vital. Therefore, we are motivated to study an MA-aided networked ISAC system, that incorporates multi-static configurations and accommodates both DL and UL users, aiming to enhance its communication and sensing capabilities. Specifically, the contributions of this paper are elaborated as follows:
  • This paper considers the joint beamforming and antennas’ position design in a networked ISAC system enabled by the MA structure to promote simultaneous DL and UL communication and target sensing functionalities, which has not yet been thoroughly investigated. This novel framework can i) effectively enlarge the service region for both communication and sensing; ii) supply more spatial DoFs by leveraging the MA technology to reconstruct the wireless channel condition; iii) provide comprehensive bidirectional communication (i.e., DL and UL communication) transmission.
  • We investigate the maximization of the sum-rate for both DL and UL users while guaranteeing target detection quality via jointly optimizing the BSs’ beamforming, UL users’ power allocation, receive filter and MAs’ positions. Besides, this paper considers a very generic signal propagation model, which fully takes into account both downlink and uplink communication and reflected radar probing signals. As will be seen, this consideration in modeling significantly complicates the beamforming and antennas’ position design task. To the best of our knowledge, this problem has not been considered in the existing literature, e.g., [21]-[33].
  • The joint beamforming and MA position’s configuration design to improve sum-rate of both DL and UL users is a highly non-convex problem due to the position coefficients of MA. Via utilizing cutting-the-edge optimization frameworks including the weighted minimum mean squared error (WMMSE) [34] and majorization-minimization (MM) [35], we have developed a closed solution for updating MA’s position of the DL users. Furthermore, we successfully develop an effective algorithm that updates the position coefficients of MA by convex optimization techniques.
  • Last but not least, extensive numerical results are provided to verify the effectiveness of our proposed solution, which demonstrates as: 1) our proposed algorithm can converge within several iterations; 2) the proposed algorithm can achieve much better than the benchmark cases (especially its counterpart with the FPA-based array at both BSs and users) in terms of sum-rate of both DL and UL users under various settings. Experiment results validate that the deployment of MA can significantly benefit communication performance in the networked ISAC system.
II System Model and Problem Formulation
Figure 1: A MA-aided networked ISAC system. (a): System model (b): Illustrations of the MAs at the BS and users, respectively.
As shown in Fig. 1111The dotted line represents the link between TBS and RBS. , we consider an MA-enhanced networked ISAC system that comprises $M_{t}$ transmit BSs (TBSs), $M_{r}$ receive BSs (RBSs), $K_{d}$ single MA DL mobile users, $K_{u}$ single MA UL mobile users, one point-like sensing target, and $K_{t}$ clutter sources. Moreover, TBSs and RBSs are respectively equipped with $N_{t}$ and $N_{r}$ MAs. Let $\mathcal{M}_{t}\triangleq\{1,\cdots,M_{t}\}$, $\mathcal{M}_{r}\triangleq\{1,\cdots,M_{r}\}$, $\mathcal{N}_{t}\triangleq\{1,\cdots,N_{t}\}$, $\mathcal{N}_{r}\triangleq\{1,\cdots,N_{r}\}$, $\mathcal{K}_{d}\triangleq\{1,\cdots,K_{d}\}$, $\mathcal{K}_{u}\triangleq\{1,\cdots,K_{u}\}$ and $\mathcal{K}_{t}\triangleq\{1,\cdots,K_{t}\}$ denoted the sets of TBSs, RBSs, TBSs’ MAs, RBSs’ MAs, DL users, UL users and clutter sources, respectively.
Moreover, the MAs are connected to RF chains by flexible cables  [^2] The response times of motor-based mechanisms is the order of milliseconds to seconds [3]., and can move within local region in real time [4]-[5]. Let $\mathbf{t}^{0}_{m,n}=[x^{0}_{m,n},y^{0}_{m,n}]^{T}\in\mathcal{C}$ denote the position of the $n$-th MA of $m$-th TBS. Similarly, the $n$-th MA’ position of $m$-th RBS is given by $\mathbf{t}^{1}_{m,n}=[x^{1}_{m,n},y^{1}_{m,n}]^{T}\in\mathcal{C}$. The positions of MA at the $k$-th DL and $l$-th UL users are respectively written as $\mathbf{r}_{0,k}=[x^{d}_{k},y^{d}_{k}]^{T}\in\mathcal{C}$ and $\mathbf{r}_{1,l}=[x^{u}_{l},y^{u}_{l}]^{T}\in\mathcal{C}$, where $\mathcal{C}$ denotes the given two-dimensional (2D) moving square region with a size of $A\times A$ for both the BSs’ and users’ MAs, while the corresponding reference point’s position is $0,0]^{T}$.
In this network, target detection and communication are performed in a time-division manner [1]. TBSs, with the aid of MA, service DL users and send probing waveform to detect the target, and RBSs, with the assistance of MA, receive the information from UL users and the echo signal from the sensing target. We assumed that the BSs are connected to the central controller via fronthaul links and that cooperative sensing of the target through data-level fusion [17]-[18]. All TBSs and RBSs can achieve perfect synchronization and joint signal processing facilitated by the central controller, and the signals transmitted by the TBSs can be made available at all RBSs. Moreover, the link between BSs (i.e., intra-BS channel) is assumed to be quasi-static. Therefore, the received signal from the direct link between the TBS and the RBS can be perfectly eliminated using prior knowledge obtained through the fronthaul links, i.e., passive optical network (PON), coax, cable [36]. Since the users are located in a complex environment with multiple reflected and scattered paths (containing the target/clutters-users links), the channels of the users are assumed to be multiple paths channel model [5], that can be estimated via least square method [3].
II-A Channel Model
Since the MA’s moving area is usually much smaller than the signal propagation distance, the channels between the BSs/users and the users are assumed to follow the far-field wireless channel model [4]-[5]. Therefore, changing the position of MA has no effect on the angle of departure (AoD), the angle of arrival (AoA), and the amplitude of the complex path coefficient, while only the phase of the complex path coefficient varies with changes in MA’s position.
a) Communication Channel:
1) TBS-DL User Channel: Let $L^{t,0}_{m,k}$ and $L^{r,0}_{m,k}$ denote the total number of transmit and receive channel paths at the $m$-th TBS from the $k$-th DL user, respectively. The elevation and azimuth AoDs of the $i$-th transmit path between TBS $m$ and DL user $k$ are respectively denoted as $\theta^{t,0}_{m,k,i}$ and $\phi^{t,0}_{m,k,i}$, while the corresponding elevation and azimuth AoAs of the $i$-th receive path are represented as $\theta^{r,0}_{m,k,i}$ and $\phi^{r,0}_{m,k,i}$, respectively. In the $i$-th transmit path for $m$-th TBS-$k$-th DL user link, the propagation distance difference between the $n$-th MA and its reference point can be given as
$\rho^{t,0}_{m,k,n,i}(\mathbf{t}^{0}_{m,n})=({\mathbf{t}_{m,n}^{{0}}})^{T}\mathbf{a}^{t,0}_{m,k,i}$ (1)
$=x^{0}_{m,n}\cos(\theta^{t,0}_{m,k,i})\sin(\phi^{t,0}_{m,k,i})+y^{0}_{m,n}\sin(\theta^{t,0}_{m,k,i}),$
and its phase difference can be obtained as $2\pi\rho^{t,0}_{m,k,n,i}(\mathbf{t}^{0}_{m,n})/\lambda$, where $\lambda$ denotes the carrier wavelength.
Thus, the field-response vector (FRV) of the transmit channel paths between the $k$-th DL user and the $m$-th TBS’ $n$-th MA is given by
$\mathbf{h}^{0}_{m,k,n}\!\!=\!\![e^{j\frac{2\pi}{\lambda}\rho^{t,0}_{m,k,n,1}},e^{j\frac{2\pi}{\lambda}\rho^{t,0}_{m,k,n,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{t,0}_{m,k,n,L^{t,0}_{m,k}}}]^{T},$ (2)
and then the corresponding field response matrix (FRM) of all $N_{t}$ MAs can be written as
$\mathbf{H}^{0}_{m,k}(\mathbf{t}^{0}_{m})=[\mathbf{h}^{0}_{m,k,1},\mathbf{h}^{0}_{m,k,2},\cdots,\mathbf{h}^{0}_{m,k,N_{t}}]\in\mathbb{C}^{L^{t,0}_{m,k}\times N_{t}},$ (3)
where $\mathbf{t}^{0}_{m}\triangleq\{\mathbf{t}^{0}_{m,n}\}$. The FRV of the receive channel paths between DL user $k$ and TBS $m$ is defined as
$\mathbf{h}^{0,0}_{m,k}(\mathbf{r}_{0,k})=[e^{j\frac{2\pi}{\lambda}\rho^{r,0}_{m,k,1}},e^{j\frac{2\pi}{\lambda}\rho^{r,0}_{m,k,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{r,0}_{m,k,L^{r,0}_{m,k}}}]^{T},$ (4)
where $\rho^{r,0}_{m,k,i}(\mathbf{r}_{0,k})\!\triangleq\!\mathbf{r}_{0,k}^{T}\mathbf{a}^{r,0}_{m,k,i}\!=\!x^{d}_{k}\cos(\theta^{r,0}_{m,k,i})\sin(\phi^{r,0}_{m,k,i})+y^{d}_{k}\sin(\theta^{r,0}_{m,k,i})$. Therefore, the channel vector between the $m$-th TBS and the $k$-th DL user is given by
$\mathbf{h}_{d,m,k}^{H}(\mathbf{t}^{0}_{m},\mathbf{r}_{0,k})=\mathbf{h}^{0,0}_{m,k}(\mathbf{r}_{0,k})^{H}\mathbf{\Sigma}^{0}_{m,k}\mathbf{H}^{0}_{m,k}(\mathbf{t}^{0}_{m})\in\mathbb{C}^{1\times N_{t}},$ (5)
where the matrix $\mathbf{\Sigma}^{0}_{m,k}$ is denoted as the response of all transmit and receive paths from TBS $m$ to DL user $k$.
2) RBS-UL User Channel: Let $L^{t,1}_{m,k}$ and $L^{r,1}_{m,k}$ denote the total number of transmit and receive channel paths at RBS $m$ from UL user $k$, respectively. The elevation and azimuth AoDs of the $i$-th transmit path between RBS $m$ and UL user $k$ are respectively denoted as $\theta^{t,1}_{m,k,i}$ and $\phi^{t,1}_{m,k,i}$, while the corresponding elevation and azimuth AoAs of the $i$-th receive path are represented as $\theta^{r,1}_{m,k,i}$ and $\phi^{r,1}_{m,k,i}$, respectively. In the $i$-th transmit path for $m$-th RBS-$k$-th UL user link, the phase difference of the $n$-th MA and its reference point is given by
$2\pi\rho^{t,1}_{m,k,n,i}(\mathbf{t}^{1}_{m,n})/\lambda=2\pi({\mathbf{t}_{m,n}^{1}})^{T}\mathbf{a}^{t,1}_{m,k,i}/\lambda$ (6)
$=2\pi\big(x^{1}_{m,n}\cos(\theta^{t,1}_{m,k,i})\sin(\phi^{t,1}_{m,k,i})+y^{1}_{m,n}\sin(\theta^{t,1}_{m,k,i})\big)/\lambda.$
Thus, the FRV of the transmit channel paths between the $k$-th UL user and the $m$-th RBS’ $n$-th MA is given by
$\mathbf{h}^{1}_{m,k,n}\!\!=\!\![e^{j\frac{2\pi}{\lambda}\rho^{t,1}_{m,k,n,1}},e^{j\frac{2\pi}{\lambda}\rho^{t,1}_{m,k,n,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{t,1}_{m,k,n,L^{t,1}_{m,k}}}]^{T},$ (7)
and then the corresponding FRM of all $N_{r}$ MAs is written as
$\mathbf{H}^{1}_{m,k}(\mathbf{t}^{0}_{m})=[\mathbf{h}^{1}_{m,k,1},\mathbf{h}^{1}_{m,k,2},\cdots,\mathbf{h}^{1}_{m,k,N_{r}}]\in\mathbb{C}^{L^{t,1}_{m,k}\times N_{r}},$ (8)
where $\mathbf{t}^{1}_{m}\triangleq\{\mathbf{t}^{1}_{m,n}\}$. The FRV of the receive channel paths between UL user $k$ and RBS $m$ is given as
$\mathbf{h}^{0,1}_{m,k}(\mathbf{r}_{1,k})=[e^{j\frac{2\pi}{\lambda}\rho^{r,1}_{m,k,1}},e^{j\frac{2\pi}{\lambda}\rho^{r,1}_{m,k,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{r,1}_{m,k,L^{r,1}_{m,k}}}]^{T},$ (9)
where $\rho^{r,1}_{m,k,i}(\mathbf{r}_{1,k})\!\!\triangleq\!\!\mathbf{r}_{1,k}^{T}\mathbf{a}^{r,1}_{m,k,i}\!=\!x^{u}_{k}\cos(\theta^{r,1}_{m,k,i})\sin(\phi^{r,1}_{m,k,i})+y^{u}_{k}\sin(\theta^{r,1}_{m,k,i})$. Therefore, the channel vector between the $m$-th RBS and the $k$-th UL user is represented as
$\mathbf{h}_{u,m,k}^{H}(\mathbf{t}^{1}_{m},\mathbf{r}_{1,k})\!=\!\mathbf{h}^{0,1}_{m,k}(\mathbf{r}_{1,k})^{H}\mathbf{\Sigma}^{1}_{m,k}\mathbf{H}^{1}_{m,k}(\mathbf{t}^{1}_{m})\in\mathbb{C}^{1\times N_{r}},$ (10)
with the matrix $\mathbf{\Sigma}^{1}_{m,k}$ denoting as the response of all transmit and receive paths from the $m$-th RBS to the $k$-th UL user.
3) DL User-UL User Channel: $L^{t,2}_{k,l}$ and $L^{r,2}_{k,l}$ are respectively denoted as the total number of transmit and receive channel paths between DL user $k$ and UL user $l$. Denote the elevation and azimuth AoDs of the $i$-th transmit path between UL user $l$ and DL user $k$ as $\theta^{t,2}_{l,k,i}$ and $\phi^{t,2}_{l,k,i}$, respectively, while the corresponding elevation and azimuth AoAs for the $i$-th receive path are represented as $\theta^{r,2}_{l,k,i}$ and $\phi^{r,2}_{l,k,i}$, respectively.
The FRVs of the transmit and receive channel paths between DL user $k$ and UL user $l$ are respectively expressed as
$\mathbf{h}^{2}_{k,l}(\mathbf{r}_{1,l})=[e^{j\frac{2\pi}{\lambda}\rho^{t,2}_{k,l,1}},e^{j\frac{2\pi}{\lambda}\rho^{t,2}_{k,l,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{t,2}_{k,l,L^{t,2}_{k,l}}}]^{T},$ (11)
$\mathbf{h}^{3}_{k,l}(\mathbf{r}_{0,k})=[e^{j\frac{2\pi}{\lambda}\rho^{r,2}_{k,l,1}},e^{j\frac{2\pi}{\lambda}\rho^{r,2}_{k,l,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{r,2}_{k,l,L^{r,2}_{k,l}}}]^{T},$ (12)
where $\rho^{t,2}_{k,l,i}(\mathbf{r}_{1,l})\triangleq\mathbf{r}_{1,l}^{T}\mathbf{a}^{t,2}_{k,l,i}=x^{u}_{l}\cos(\theta^{t,2}_{k,l,i})\sin(\phi^{t,2}_{k,l,i})+y^{u}_{l}\sin(\theta^{t,2}_{k,l,i})$ and $\rho^{r,2}_{k,l,i}(\mathbf{r}_{0,k})\triangleq\mathbf{r}_{0,k}^{T}\mathbf{a}^{r,2}_{k,l,i}=y^{d}_{k}\sin(\theta^{r,2}_{k,l,i})+x^{d}_{k}\cos(\theta^{r,2}_{k,l,i})\sin(\phi^{r,2}_{k,l,i})$. And, the channel between the $k$-th DL user and the $l$-th UL user can be written as
${h}_{du,k,l}^{*}(\mathbf{r}_{0,k},\mathbf{r}_{1,l})=(\mathbf{h}^{3}_{k,l}(\mathbf{r}_{0,k}))^{H}\mathbf{\Sigma}^{2}_{k,l}\mathbf{h}^{2}_{k,l}(\mathbf{r}_{1,l}),$ (13)
where the matrix $\mathbf{\Sigma}^{2}_{k,l}$ is denoted as the response of all transmit and receive paths from the $l$-th UL user to the $k$-th DL user.
b) Radar Channel: The steering vector of the $m$-th TBS towards the sensing target and/or clutters is represented by
$\mathbf{g}^{t}_{m,j}=[e^{j\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{m,1})^{T}\mathbf{a}^{4}_{m,j}},e^{j\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{m,2})^{T}\mathbf{a}^{4}_{m,j}},$ (14)
$\cdots,e^{j\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{m,N_{t}})^{T}\mathbf{a}^{4}_{m,j}}],\forall j\in\mathcal{J}\triangleq\{0\}\cup\mathcal{K}_{t},$
where index 0 represents the sensing target, $\mathbf{a}^{4}_{m,j}\triangleq[\cos(\theta^{t,4}_{m,j})\sin(\phi^{t,4}_{m,j});\sin(\theta^{t,4}_{m,j})]^{T}$, and $\theta^{t,4}_{m,j}$ and $\phi^{t,4}_{m,j}$ denote the elevation and azimuth AoDs between TBS $m$ and target/clutters, respectively. Similarly, the steering vector of RBS $p$ towards the sensing target/clutters can also be written as
$\mathbf{g}^{r}_{p,j}=[e^{j\frac{2\pi}{\lambda}(\mathbf{t}^{1}_{p,1})^{T}\mathbf{a}^{5}_{p,j}},e^{j\frac{2\pi}{\lambda}(\mathbf{t}^{1}_{p,2})^{T}\mathbf{a}^{5}_{p,j}},$ (15)
$\cdots,e^{j\frac{2\pi}{\lambda}(\mathbf{t}^{1}_{p,N_{r}})^{T}\mathbf{a}^{5}_{p,j}}],\forall j\in\mathcal{J},$
where $\mathbf{a}^{5}_{p,j}\triangleq[\cos(\theta^{r,4}_{p,j})\sin(\phi^{r,4}_{p,j});\sin(\theta^{r,4}_{p,j})]^{T}$, and $\theta^{r,4}_{p,j}$ and $\phi^{r,4}_{p,j}$ denote respectively the elevation and azimuth AoAs between RBS $p$ and target/clutters. Therefore, the $m$-th TBS-target/clutters and the $p$-th RBS-target/clutters channels are respectively given as
$\mathbf{g}_{t,m,j}=\beta^{t}_{m,j}\mathbf{g}^{t}_{m,j},\ \mathbf{g}_{r,p,j}=\beta^{r}_{p,j}\mathbf{g}^{r}_{p,j},$ (16)
where $\{\beta^{t}_{m,j}\}$ and $\{\beta^{r}_{p,j}\}$ are complex fading coefficients and are assumed to be known [1].
II-B Signal Model
Denote $\mathbf{s}\triangleq[s_{1},\cdots,s_{k},\cdots,s_{K_{d}}]^{T}\in\mathbb{C}^{K_{d}\times 1}$ and $\mathbf{s}_{r}\in\mathbb{C}^{N_{t}\times 1}$ as the communication symbol transmitted to DL users and radar probing signal from TBS, respectively. For simplicity, it is assumed that $\mathbb{E}\{\mathbf{s}\mathbf{s}^{H}\}=\mathbf{I}$, $\mathbb{E}\{\mathbf{s}_{r}\mathbf{s}_{r}^{H}\}=\mathbf{I}$ and $\mathbb{E}\{\mathbf{s}\mathbf{s}_{r}^{H}\}=\mathbf{0}$. The transmit signal of the $m$-th TBS can be written as  [^3] The joint communication and sensing signal model effectively exploits the available DoF [37].
$\mathbf{x}_{m}={\sum}_{k=1}^{K_{d}}\mathbf{w}_{m,k}s_{k}+\mathbf{W}^{r}_{m}\mathbf{s}_{r},\forall m\in\mathcal{M}_{t},$ (17)
where $\mathbf{w}_{m,k}\in\mathbb{C}^{N_{t}\times 1}$ and $\mathbf{W}^{r}_{m}\in\mathbb{C}^{N_{t}\times N_{t}}$ represent the transmit beamformer of the $m$-th TBS for the $k$-th DL user’s communication signal and the probing signal, respectively.
The $l$-th UL user’s uplink signal can be written as
$x^{u}_{l}=\sqrt{q_{l}}s_{l}^{u},\ \forall l\in\mathcal{K}_{u},$ (18)
where $s_{l}^{u}$ and $q_{l}$ denote the information symbol and transmission power of UL user $l$, respectively, and $s_{l}^{u}$ are mutually uncorrelated and each has zero mean and unit variance.
The received signal at DL user $k$ can be written as
$y_{D,k}=\underbrace{{\sum}_{m=1}^{M_{t}}\mathbf{h}_{d,m,k}^{H}\mathbf{x}_{m}}\limits_{\textrm{DL users signals}}+\underbrace{{\sum}_{l=1}^{K_{u}}{h}_{du,k,l}x^{u}_{l}}\limits_{\textrm{UL users signals}}+n_{d,k}$ (19)
$={\sum}_{i=1}^{K_{d}}\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{i}s_{i}+\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{W}}^{r}\mathbf{s}^{r}+{\sum}_{l=1}^{K_{u}}{h}_{du,k,l}x^{u}_{l}+n_{d,k},$
where $\hat{\mathbf{h}}_{d,k}\triangleq[\mathbf{h}_{d,1,k}^{T},\mathbf{h}_{d,2,k}^{T},\cdots,\mathbf{h}_{d,M_{t},k}^{T}]^{T}\in\mathbb{C}^{M_{t}N_{t}\times 1}$, $\hat{\mathbf{w}}_{k}\triangleq[\mathbf{w}_{1,k}^{T},\mathbf{w}_{2,k}^{T},\cdots,\mathbf{w}_{M_{t},k}^{T}]^{T}\in\mathbb{C}^{M_{t}N_{t}\times 1}$, $\bar{\mathbf{W}}_{r}\triangleq[(\mathbf{W}^{r}_{1})^{T},(\mathbf{W}^{r}_{2})^{T},\cdots,(\mathbf{W}^{r}_{M_{t}})^{T}]^{T}\in\mathbb{C}^{M_{t}N_{t}\times N_{t}}$, $\hat{\mathbf{W}}_{r}\triangleq\mathrm{blkdiag}(\bar{\mathbf{W}}_{r})$ and $n_{d,k}\!\!\sim\!\!\mathcal{C\!N}(0,\sigma_{d,k}^{2})$ is denoted as the complex additive white Gaussian noise (AWGN) at DL user $k$.
II-C Problem Formulation
a) Communication Rate: The received signal at the $p$-th RBS can be expressed as
$\mathbf{y}_{r,p}$ $=\underbrace{{\sum}_{j=0}^{J}\alpha_{t,j}\mathbf{g}_{r,p,j}({\sum}_{m=1}^{M_{t}}\mathbf{g}_{t,m,j}^{H}\mathbf{x}_{m})}\limits_{\textrm{sensing echoes}}$ (20)
$+\underbrace{{\sum}_{l=1}^{K_{u}}\mathbf{h}_{u,p,l}x^{u}_{l}}\limits_{\textrm{UL users signals}}+\mathbf{n}_{r,p},$
where $\mathbf{n}_{r,p}\sim\mathcal{CN}(0,\sigma_{r}^{2}\mathbf{I})$ is the AWGN at the $p$-th RBS, $\alpha_{t,j}$ represents the target RCS between TBS and RBS and $\alpha_{t,j}\sim\mathcal{CN}(0,\sigma_{t,j}^{2})$ [38].
To obtain the information of UL users and target separately, RBS $p$ will adopt $K_{u}+1$ linear filters $\mathbf{u}_{p,i}\in\mathbb{C}^{M_{r}\times 1}$, $\forall i\in\mathcal{J}_{u}\triangleq\{0\}\cup\mathcal{K}_{u}$, to post-process the received signal, where $\mathbf{u}_{p,0}$ corresponds to the radar sensing filter. Therefore, the output information of the $i$-th filter at the $p$-th RBS can be written as
$\hat{y}_{r,p,i}=\mathbf{u}_{p,i}^{H}\mathbf{y}_{r,p},\forall i\in\mathcal{J}_{u},$ (21)
and then the information from all $M_{r}$ RBS collected at the central controller is given by
$\hat{y}^{u}_{r,i}={\sum}_{p=1}^{M_{r}}\hat{y}_{r,p,i}={\sum}_{l=1}^{K_{u}}\hat{\mathbf{u}}_{i}^{H}\hat{\mathbf{h}}_{u,l}x^{u}_{l}$ (22)
$+{\sum}_{j=0}^{K_{t}}\alpha_{j}\hat{\mathbf{u}}_{i}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}({\sum}_{k=1}^{K_{d}}\hat{\mathbf{w}}_{k}s_{k}+\hat{\mathbf{W}}_{r}\hat{\mathbf{s}}^{r})+\hat{\mathbf{u}}_{i}^{H}\mathbf{n}_{r},$
where
$\hat{\mathbf{u}}_{i}\triangleq[\mathbf{u}_{1,i}^{T},\mathbf{u}_{2,i}^{T},\cdots,\mathbf{u}_{M_{r},i}^{T}]^{T}\in\mathbb{C}^{M_{r}N_{r}\times 1},$ (23)
$\hat{\mathbf{h}}_{u,l}\triangleq[\mathbf{h}_{u,1,l}^{T},\mathbf{h}_{u,2,l}^{T},\cdots,\mathbf{h}_{u,M_{r},l}^{T}]^{T}\in\mathbb{C}^{M_{r}N_{r}\times 1},$
$\hat{\mathbf{g}}_{t,j}\triangleq[\mathbf{g}_{t,1,j}^{T},\mathbf{g}_{t,2,j}^{T},\cdots,\mathbf{g}_{t,M_{t},j}^{T}]^{T}\in\mathbb{C}^{M_{t}N_{t}\times 1},$
$\hat{\mathbf{g}}_{r,j}\triangleq[\mathbf{g}_{r,1,j}^{T},\mathbf{g}_{r,2,j}^{T},\cdots,\mathbf{g}_{r,M_{r},j}^{T}]^{T}\in\mathbb{C}^{M_{r}N_{r}\times 1},$
$\mathbf{n}_{r}\triangleq[\mathbf{n}_{r,1}^{T},\mathbf{n}_{r,2}^{T},\cdots,\mathbf{n}_{r,M_{r}}^{T}]^{T}\in\mathbb{C}^{M_{r}N_{r}\times 1}.$
The achievable rate of the $k$-th DL user and the $l$-th UL user can be obtained as
$\textrm{R}_{d,k}=\textrm{log}(1+\textrm{SINR}_{d,k}),\forall k\in\mathcal{K}_{d},$ (24)
$\textrm{R}_{u,l}=\textrm{log}(1+\textrm{SINR}_{u,l}),\forall l\in\mathcal{K}_{u},$ (25)
respectively, where
$\textrm{SINR}_{d,k}=$ (26)
$\frac{|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{k}|^{2}}{{\sum}_{i\neq k}^{K_{d}}|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{i}|^{2}+\|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{W}}^{r}\|^{2}_{2}+{\sum}_{l=1}^{K_{u}}q_{l}|{h}_{du,k,l}|^{2}+\sigma_{d,k}^{2}},$
and the SINR formulation of UL user is written in (27), as shown on the top of this page.
$\textrm{SINR}_{u,l}=\frac{q_{l}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,l}|^{2}}{\sum_{i\neq l}^{K_{u}}q_{i}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,i}|^{2}+\sum_{j=0}^{K_{t}}\sum_{k=1}^{K_{d}}\sigma_{t,j}^{2}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}|^{2}+\sum_{j=0}^{K_{t}}\sigma_{t,j}^{2}\|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{2}+\sigma_{r}^{2}\|\hat{\mathbf{u}}_{l}^{H}\|^{2}_{2}}$ (27)
b) Sensing SINR: In the target detection problem, since the detection probability for a given false alarm probability monotonically increases with the sensing SINR, we select the sensing SINR as the metric [39]-[40]. The radar SINR [39]-[40] for sensing can be derived as (28), as shown on the top of this page.
$\textrm{SINR}_{t}=\frac{\sum_{k=1}^{K_{d}}\sigma_{t,0}^{2}|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,0}\hat{\mathbf{g}}_{t,0}^{H}\hat{\mathbf{w}}_{k}|^{2}+\sigma_{t,0}^{2}\|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,0}\hat{\mathbf{g}}_{t,0}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{F}}{\sum_{i=1}^{K_{u}}q_{i}|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{h}}_{u,i}|^{2}+\sum_{j=1}^{K_{t}}\sum_{k=1}^{K_{d}}\sigma_{t,j}^{2}|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}|^{2}+\sum_{j=1}^{K_{t}}\sigma_{t,j}^{2}\|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{2}+\sigma_{r}^{2}\|\hat{\mathbf{u}}_{0}^{H}\|^{2}_{2}}$ (28)
c) Objective Function and Constraints: Our goal is to maximize the sum-rate of all DL and UL users via jointly optimizing the transmit beamformer $\{\mathbf{w}_{m,k},\mathbf{W}^{r}_{m}\}$, the linear filters $\{\mathbf{u}_{p,i}\}$, the DL users’ uplink transmit power $\{q_{l}\}$, the positions of all MAs $\{\mathbf{t}^{0}_{m,n},\mathbf{t}^{1}_{p,j},\mathbf{r}_{0,k},\mathbf{r}_{1,l}\}$. The weighted sum-rate optimization problem can be formulated as
$\textrm{(P0)}:$ $\mathop{\textrm{max}}\limits_{\{\mathbf{w}_{m,k},\mathbf{W}^{r}_{m}\},\{\mathbf{u}_{p,i}\},\{q_{l}\},\atop\{\mathbf{t}^{0}_{m,n},\mathbf{t}^{1}_{p,j},\mathbf{r}_{0,k},\mathbf{r}_{1,l}\}}\ \!\!\!{\sum}_{k=1}^{K_{d}}\mu_{D,k}\textrm{R}_{d,k}\!\!+\!\!{\sum}_{l=1}^{K_{u}}\mu_{U,l}\textrm{R}_{u,l}$ (29a)
s.t. $\textrm{SINR}_{t}\geq\Gamma_{r},$ (29b)
${\sum}_{k=1}^{K_{d}}\|\mathbf{w}_{m,k}\|_{2}^{2}+\|\mathbf{W}^{r}_{m}\|_{F}^{2}\leq{P}^{BS}_{m},\forall m\in\mathcal{M}_{t},$ (29c)
$0\leq q_{l}\leq P_{u,l},\forall l\in\mathcal{K}_{u},$ (29d)
$\|\mathbf{t}^{0}_{m,n}\!-\!\mathbf{t}^{0}_{m,i}\|_{2}\!\geq\!D_{t},\forall n,i\!\in\!\mathcal{N}_{t},n\!\neq\!i,\forall m\!\in\!\mathcal{M}_{t},$ (29e)
$\|\mathbf{t}^{1}_{p,j}\!-\!\mathbf{t}^{1}_{p,s}\|_{2}\!\geq\!D_{t},\forall j,s\in\mathcal{N}_{r},j\neq s,\forall p\in\mathcal{M}_{r},$ (29f)
$\mathbf{t}^{0}_{m,n}\in\mathcal{C},\ \mathbf{t}^{1}_{p,j}\in\mathcal{C},\ \mathbf{r}_{0,k}\in\mathcal{C},\ \mathbf{r}_{1,l}\in\mathcal{C},$ (29g)
where $\mu_{D,k}\in(0,1)$ and $\mu_{U,l}\in(0,1)$, and ${\sum}_{k=1}^{K_{d}}\mu_{D,k}+{\sum}_{l=1}^{K_{u}}\mu_{U,l}=1$  [^4] Since the weight factors are treated as constants in the optimization process, we will omit them in the following., $\Gamma_{r}$ denotes the minimum target sensing SINR threshold, $P^{BS}_{m}$ and $P_{u,l}$ represent the maximum transmission power of the $m$-th TBS and the $l$-th UL user, respectively, and $D_{t}$ is the minimum distance between MAs of BSs to avoid the coupling effect. Obviously, due to the non-convex objective function and constraints, the optimization problem (P0) is highly challenging.
III Algorithm Design
III-A Problem Reformulation
To make the problem (P0) more tractable, we first employ the WMMSE method [34] to equivalently transform the objective functions $\textrm{R}_{d,k}$ and $\textrm{R}_{u,l}$ into (31) and (32), respectively, at the top of the next page, where $\{\omega_{d,k}\}$, $\{\beta_{d,k}\}$, $\{\omega_{u,l}\}$ and $\{\beta_{u,l}\}$ are the introduced auxiliary variables. Therefore, the original problem (P0) can be reexpressed as
$\textrm{(P1)}:$ $\mathop{\textrm{max}}\limits_{\{\mathbf{w}_{m,k},\mathbf{W}^{r}_{m}\},\{\mathbf{u}_{p,i}\},\{q_{l}\},\atop\{\mathbf{t}^{0}_{m,n},\mathbf{t}^{1}_{p,j},\mathbf{r}_{0,k},\mathbf{r}_{1,l}\}}\ \!\!{\sum}_{k=1}^{K_{d}}\tilde{\mathrm{R}}_{d,k}\!+\!{\sum}_{l=1}^{K_{u}}\tilde{\mathrm{R}}_{u,l}$ (30a)
s.t. $\textrm{SINR}_{t}\geq\Gamma_{r},$ (30b)
${\sum}_{k=1}^{K_{d}}\|\mathbf{w}_{m,k}\|_{2}^{2}+\|\mathbf{W}^{r}_{m}\|_{F}^{2}\leq\textrm{P}^{BS}_{m},\forall m\in\mathcal{M}_{t},$ (30c)
$0\leq q_{l}\leq P_{u,l},\forall l\in\mathcal{K}_{u},$ (30d)
$\|\mathbf{t}^{0}_{m,n}\!-\!\mathbf{t}^{0}_{m,i}\|_{2}\!\geq\!D_{t},\forall n,i\!\in\!\mathcal{N}_{t},n\!\neq\!i,\forall m\!\in\!\mathcal{M}_{t},$ (30e)
$\|\mathbf{t}^{1}_{p,j}\!-\!\mathbf{t}^{1}_{p,s}\|_{2}\!\geq\!D_{t},\forall j,s\in\mathcal{N}_{r},j\neq s,\forall p\in\mathcal{M}_{r},$ (30f)
$\mathbf{t}^{0}_{m,n}\in\mathcal{C},\ \mathbf{t}^{1}_{p,j}\in\mathcal{C},\ \mathbf{r}_{0,k}\in\mathcal{C},\ \mathbf{r}_{1,l}\in\mathcal{C},$ (30g)
Next, we will adopt block coordinate ascent (BCA) method [41] to efficiently solve problem (P1) by updating different blocks of variables.
$\tilde{\mathrm{R}}_{d,k}\triangleq\textrm{log}(\omega_{d,k})-\omega_{d,k}\big(1-2\textrm{Re}\{\beta_{d,k}^{\ast}\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{k}\}+|\beta_{d,k}|^{2}({\sum}_{i=1}^{K_{d}}|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{i}|^{2}+\|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{W}}^{r}\|^{2}_{2}+{\sum}_{l=1}^{K_{u}}q_{l}|{h}_{du,k,l}|^{2}+\sigma_{d,k}^{2})\big)+1$ (31)
$\tilde{\mathrm{R}}_{u,l}\triangleq\textrm{log}(\omega_{u,l})-\omega_{u,l}\bigg(1-2\textrm{Re}\{\beta_{u,l}^{\ast}\sqrt{q_{l}}\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,l}\}$ (32)
$+|\beta_{u,l}|^{2}\big({\sum}_{i=1}^{K_{u}}q_{i}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,i}|^{2}+{\sum}_{j=0}^{K_{t}}\sigma_{t,j}^{2}({\sum}_{k=1}^{K_{d}}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}|^{2}+\|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{2})+\sigma_{r}^{2}\|\hat{\mathbf{u}}_{l}^{H}\|^{2}_{2}\big)\bigg)+1$
III-B Optimizing auxiliary variables
With other variables being fixed, the analytical solutions of the auxiliary variables, i.e., $\{\omega_{d,k}\}$, $\{\beta_{d,k}\}$, $\{\omega_{u,l}\}$ and $\{\beta_{u,l}\}$, can be obtained from the derivation of WMMSE transformation proven in [34], as shown in (33)-(36).
$\beta^{\star}_{d,k}={\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{k}}/({{\sum}_{i=1}^{K_{d}}|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{i}|^{2}+\|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{W}}^{r}\|^{2}_{2}+{\sum}_{l=1}^{K_{u}}q_{l}|{h}_{du,k,l}|^{2}+\sigma_{d,k}^{2}}),$ (33)
$\omega^{\star}_{d,k}=\big(|1-\beta^{\ast}_{d,k}\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{k}|^{2}+|\beta^{\ast}_{d,k}|^{2}({\sum}_{i\neq k}^{K_{d}}|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{i}|^{2}+\|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{W}}^{r}\|^{2}_{2}+{\sum}_{l=1}^{K_{u}}q_{l}|{h}_{du,k,l}|^{2}+\sigma_{d,k}^{2})\big)^{-1},$ (34)
$\beta^{\star}_{u,l}={\sqrt{q_{l}}\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,l}}/({{\sum}_{i=1}^{K_{u}}q_{i}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,i}|^{2}+{\sum}_{j=0}^{K_{t}}\sigma_{t,j}^{2}({\sum}_{k=1}^{K_{d}}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}|^{2}+\|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{2})+\sigma_{r}^{2}\|\hat{\mathbf{u}}_{l}^{H}\|^{2}_{2}}),$ (35)
$\omega^{\star}_{u,l}=\bigg(|1-\beta^{\ast}_{u,l}\sqrt{q_{l}}\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,l}|^{2}+|\beta^{\ast}_{u,l}|^{2}\big({\sum}_{j=0}^{K_{t}}\sigma_{t,j}^{2}({\sum}_{k=1}^{K_{d}}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}|^{2}+\|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{2})+\sigma_{r}^{2}\|\hat{\mathbf{u}}_{l}^{H}\|^{2}_{2}\big)\bigg)^{-1}$ (36)
III-C Updating The TBS Beamformer
In this subsection, we investigate the optimization of all $M_{t}$ TBS beamformer $\{\mathbf{w}_{m,k},\mathbf{W}^{r}_{m}\}$ with other variables being given. Firstly, the functions $\tilde{\mathrm{R}}_{d,k}$ and $\tilde{\mathrm{R}}_{u,l}$, which are respectively defined in (31) and (32), are rewritten as
$-\!\sum_{k=1}^{K_{d}}\!\tilde{\mathrm{R}}_{\!d\!,k}\!\!=\!\!\!\sum_{k=1}^{K_{d}}\!(\!\hat{\mathbf{w}}^{H}_{k}\!\!\mathbf{A}_{1}\!\hat{\mathbf{w}}_{k}\!\!-\!\!2\textrm{Re}\{\!\mathbf{a}^{H}_{1,k}\hat{\mathbf{w}}_{k}\!\}\!)\!\!+\!\!\!\sum_{i=1}^{M_{t}}\!(\!\bar{\mathbf{w}}^{r}_{i}\!)^{\!H\!}\!\mathbf{A}_{\!2\!,i}\!\bar{\mathbf{w}}^{r}_{i}\!\!-\!\!c_{\!1\!,\!1},$ (37)
$-\sum_{l=1}^{K_{u}}\tilde{\mathrm{R}}_{u,l}\!=\!\sum_{k=1}^{K_{d}}\!\hat{\mathbf{w}}^{H}_{k}\mathbf{A}_{3}\hat{\mathbf{w}}_{k}\!+\!\sum_{i=1}^{M_{t}}\!(\bar{\mathbf{w}}^{r}_{i})^{H}\mathbf{A}_{4,i}\bar{\mathbf{w}}^{r}_{i}\!-\!c_{1,2},$ (38)
where $c_{1,1}$ and $c_{1,2}$ are constant terms and the newly introduced parameters in (37) and (38) being defined as follows
$\mathbf{a}_{1,k}\triangleq\omega_{d,k}\beta_{d,k}\hat{\mathbf{h}}_{d,k},\bar{\mathbf{w}}^{r}_{i}\triangleq\textrm{vec}({\mathbf{W}^{r}_{i}}),$ (39)
$\mathbf{A}_{1}\triangleq{\sum}_{k=1}^{K_{d}}\omega_{d,k}|\beta_{d,k}|^{2}\hat{\mathbf{h}}_{d,k}\hat{\mathbf{h}}_{d,k}^{H},\mathbf{A}_{2,i}\triangleq\mathbf{I}\otimes{\mathbf{h}}_{d,i,k}{\mathbf{h}}_{d,i,k}^{H},$
$\mathbf{A}_{3}\triangleq{\sum}_{l=1}^{K_{u}}{\sum}_{j=0}^{K_{t}}\beta_{u,l}|\beta_{u,l}|^{2}\sigma_{t,j}^{2}\hat{\mathbf{g}}_{t,j}\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{l}\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H},$
$\mathbf{A}_{4,i}\!\triangleq\!\!{\sum}_{l=1}^{K_{u}}{\sum}_{j=0}^{K_{t}}\beta_{u,l}|\beta_{u,l}|^{2}\big(\sigma_{t,j}^{2}\mathbf{I}\!\otimes\!({\mathbf{g}}_{t,i,j}\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{l}\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}{\mathbf{g}}_{t,i,j}^{H})\big).$
Similarly, the constraint (30b) can be rewritten as
$\textrm{SINR}_{t}\geq\Gamma_{r}\Longleftrightarrow{\sum}_{k=1}^{K_{d}}\hat{\mathbf{w}}_{k}^{H}\mathbf{A}_{11}\hat{\mathbf{w}}_{k}+{\sum}_{i=1}^{M_{t}}(\bar{\mathbf{w}}^{r}_{i})^{H}\mathbf{A}_{12,i}\bar{\mathbf{w}}^{r}_{i}$
$+c_{1,3}-{\sum}_{k=1}^{K_{d}}\hat{\mathbf{w}}_{k}^{H}\mathbf{A}_{9}\hat{\mathbf{w}}_{k}-{\sum}_{i=1}^{M_{t}}(\bar{\mathbf{w}}^{r}_{i})^{H}\mathbf{A}_{10,i}\bar{\mathbf{w}}^{r}_{i}\leq 0,$ (40)
where $c_{1,3}$ is constant and the above newly introduced coefficients are defined as
$\mathbf{A}_{9}\triangleq\sigma^{2}_{t,0}\hat{\mathbf{g}}_{t,0}\hat{\mathbf{g}}_{r,0}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,0}\hat{\mathbf{g}}_{t,0}^{H}/\Gamma_{r},$ (41)
$\mathbf{A}_{10,i}\triangleq\mathbf{I}\otimes\sigma^{2}_{t,0}{\mathbf{g}}_{t,i,0}\hat{\mathbf{g}}_{r,0}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,0}{\mathbf{g}}_{t,i,0}^{H}/\Gamma_{r},$
$\mathbf{A}_{11}\triangleq{\sum}_{j=1}^{K_{t}}\sigma^{2}_{t,j}\hat{\mathbf{g}}_{t,j}\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H},$
$\mathbf{A}_{12,i}\triangleq{\sum}_{j=1}^{K_{t}}\mathbf{I}\otimes\sigma^{2}_{t,j}{\mathbf{g}}_{t,i,j}\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,j}{\mathbf{g}}_{t,i,j}^{H}.$
Based on the above equivalent transformation, the problem of updating $\{\hat{\mathbf{w}}_{k},\bar{\mathbf{w}}^{r}_{i}\}$ can be formulated as
$\textrm{(P2)}:$ $\mathop{\textrm{min}}\limits_{\{\hat{\mathbf{w}}_{k},\bar{\mathbf{w}}_{r,i}\}}\ {\sum}_{k=1}^{K_{d}}(\hat{\mathbf{w}}^{H}_{k}\mathbf{A}_{5}\hat{\mathbf{w}}_{k}-2\textrm{Re}\{\mathbf{a}^{H}_{1,k}\hat{\mathbf{w}}_{k}\})$ (42a)
$+{\sum}_{i=1}^{M_{t}}(\bar{\mathbf{w}}^{r}_{i})^{H}\mathbf{A}_{6,i}\bar{\mathbf{w}}^{r}_{i}-\hat{c}_{1,2}$
s.t. ${\sum}_{k=1}^{K_{d}}\hat{\mathbf{w}}_{k}^{H}\mathbf{A}_{11}\hat{\mathbf{w}}_{k}\!+\!\!{\sum}_{i=1}^{M_{t}}(\bar{\mathbf{w}}^{r}_{i})^{H}\!\mathbf{A}_{12,i}\bar{\mathbf{w}}^{r}_{i}\!+\!c_{1,3}$ (42b)
- ${\sum}_{k=1}^{K_{d}}\hat{\mathbf{w}}_{k}^{H}\mathbf{A}_{9}\hat{\mathbf{w}}_{k}-{\sum}_{i=1}^{M_{t}}(\bar{\mathbf{w}}^{r}_{i})^{H}\mathbf{A}_{10,i}\bar{\mathbf{w}}^{r}_{i}\leq 0,$
${\sum}_{k=1}^{K_{d}}\|\hat{\mathbf{w}}_{k}\mathbf{R}_{i}\|_{2}^{2}+\|\bar{\mathbf{w}}^{r}_{i}\|_{2}^{2}\leq\textrm{P}^{BS}_{i},\forall i\in\mathcal{M}_{t},$ (42c)
where $\mathbf{A}_{5}\triangleq\mathbf{A}_{1}+\mathbf{A}_{3}$, $\mathbf{A}_{6,i}\triangleq\mathbf{A}_{2,i}+\mathbf{A}_{4,i}$, $\hat{c}_{1,2}\triangleq c_{1,2}+c_{1,1}$, $\mathbf{R}_{i}\triangleq\textrm{blkdiag}([\mathbf{0},\cdots,\mathbf{0}_{i-1,i-1},\mathbf{I}_{i,i},\mathbf{0}_{i+1,i+1},\cdots,\mathbf{0}])\in\mathbb{C}^{M_{t}N_{t}\times M_{t}N_{t}}$, $\mathbf{0}\in\mathbb{C}^{N_{t}\times N_{t}}$, $\mathbf{I}\in\mathbb{C}^{N_{t}\times N_{t}}$.
It is obvious that the non-convex constraint (45b) makes the problem (P2) intractable. Inspired by the MM framework [35], we respectively linearize the quadratic terms $\hat{\mathbf{w}}_{k}^{H}\mathbf{A}_{9}\hat{\mathbf{w}}_{k}$ and $\bar{\mathbf{w}}^{H}_{r,i}\mathbf{A}_{10,i}\bar{\mathbf{w}}_{r,i}$ as follows
$\hat{\mathbf{w}}_{k}^{H}\mathbf{A}_{9}\hat{\mathbf{w}}_{k}\geq 2\textrm{Re}\{\hat{\mathbf{w}}_{k,0}^{H}\mathbf{A}_{9}(\hat{\mathbf{w}}_{k}-\hat{\mathbf{w}}_{k,0})\}+\hat{\mathbf{w}}_{k,0}^{H}\mathbf{A}_{9}\hat{\mathbf{w}}_{k,0}$
$=2\textrm{Re}\{\hat{\mathbf{w}}_{k,0}^{H}\mathbf{A}_{9}\hat{\mathbf{w}}_{k}\}-(\hat{\mathbf{w}}_{k,0}^{H}\mathbf{A}_{9}\hat{\mathbf{w}}_{k,0})^{\ast},$ (43)
$(\bar{\mathbf{w}}^{r}_{i})^{H}\mathbf{A}_{10,i}\bar{\mathbf{w}}^{r}_{i}\!\!\geq\!\!2\textrm{Re}\{\!(\bar{\mathbf{w}}^{r}_{i,0})^{H}\!\mathbf{A}_{10,i}(\!\bar{\mathbf{w}}^{r}_{i}\!\!-\!\!\bar{\mathbf{w}}^{r}_{i,0}\!)\!\}\!\!+\!\!(\bar{\mathbf{w}}^{r}_{i,0})^{H}\!\mathbf{A}_{10,i}\bar{\mathbf{w}}^{r}_{i,0}$
$=2\textrm{Re}\{(\bar{\mathbf{w}}^{r}_{i,0})^{H}\mathbf{A}_{10,i}\bar{\mathbf{w}}^{r}_{i}\}-((\bar{\mathbf{w}}^{r}_{i,0})^{H}\mathbf{A}_{10,i}\bar{\mathbf{w}}^{r}_{i,0})^{\ast},$ (44)
where $\hat{\mathbf{w}}_{k,0}$ and $\bar{\mathbf{w}}^{r}_{i,0}$ are feasible solutions obtained in the last iteration. Therefore, by replacing the non-convex terms of constraint (45b) by (43) and (44), the problem (P2) can be rewritten as
$\textrm{(P3)}:$ $\mathop{\textrm{min}}\limits_{\{\hat{\mathbf{w}}_{k},\bar{\mathbf{w}}_{r,i}\}}\ {\sum}_{k=1}^{K_{d}}(\hat{\mathbf{w}}^{H}_{k}\mathbf{A}_{5}\hat{\mathbf{w}}_{k}-2\textrm{Re}\{\mathbf{a}^{H}_{1,k}\hat{\mathbf{w}}_{k}\})$ (45a)
$+{\sum}_{i=1}^{M_{t}}(\bar{\mathbf{w}}^{r}_{i})^{H}\mathbf{A}_{6,i}\bar{\mathbf{w}}^{r}_{i}-\hat{c}_{1,2}$
s.t. ${\sum}_{k=1}^{K_{d}}\hat{\mathbf{w}}_{k}^{H}\mathbf{A}_{11}\hat{\mathbf{w}}_{k}\!+\!\!{\sum}_{i=1}^{M_{t}}(\bar{\mathbf{w}}^{r}_{i})^{H}\!\mathbf{A}_{12,i}\bar{\mathbf{w}}^{r}_{i}\!+\!\hat{c}_{1,3}$ (45b)
- $\!{\sum}_{k=1}^{K_{d}}2\textrm{Re}\{\mathbf{a}_{2,k}^{H}\hat{\mathbf{w}}_{k}\}\!-\!{\sum}_{i=1}^{M_{t}}2\textrm{Re}\{\mathbf{a}_{3,i}^{H}\bar{\mathbf{w}}^{r}_{i}\}\!\leq\!0,$
${\sum}_{k=1}^{K_{d}}\|\hat{\mathbf{w}}_{k}\mathbf{R}_{i}\|_{2}^{2}+\|\bar{\mathbf{w}}^{r}_{i}\|_{2}^{2}\leq\textrm{P}^{BS}_{i},\forall i\in\mathcal{M}_{t},$ (45c)
where $\mathbf{a}_{2,k}\triangleq\mathbf{A}_{9}^{H}\hat{\mathbf{w}}_{k,0}$, $\mathbf{a}_{3,i}\triangleq\mathbf{A}_{10,i}^{H}\bar{\mathbf{w}}^{r}_{i,0}$, $\hat{c}_{1,3}\triangleq\!-\!{\sum}_{k=1}^{K_{d}}\!(\hat{\mathbf{w}}_{k,0}^{H}\mathbf{A}_{9}\hat{\mathbf{w}}_{k,0})^{\ast}\!-\!{\sum}_{i=1}^{M_{t}}\!((\bar{\mathbf{w}}^{r}_{i,0})^{H}\mathbf{A}_{10,i}\bar{\mathbf{w}}^{r}_{i,0})^{\ast}+{c}_{1,3}$. The problem (P3) is a typical second order cone program (SOCP) and can be efficiently solved by standard numerical solvers, such as CVX [42].
III-D Optimizing The UL User Transmission Power
In this subsection, we present the method for optimizing the UL users’ transmission power allocation. With other variables being fixed, the optimization problem of updating $\{q_{l}\}$ is formulated as
$\textrm{(P4)}:$ $\mathop{\textrm{min}}\limits_{\{q_{l}\}}\ {\sum}_{l=1}^{K_{u}}({b}_{1,l}q_{l}-b_{2,l}\sqrt{q_{l}})+c_{2,1}$ (46a)
s.t. ${\sum}_{l=1}^{K_{u}}b_{3,l}q_{l}-c_{2,2}\leq 0,$ (46b)
$0\leq q_{l}\leq P_{u,l},\forall l\in\mathcal{K}_{u},$ (46c)
where $c_{2,1}$ and $c_{2,2}$ are constant terms, and the new parameters in problem (P4) are defined as
$b_{2,l}\triangleq 2\textrm{Re}\{\omega_{u,l}\beta_{u,l}^{\ast}\hat{\mathbf{u}}_{i}^{H}\hat{\mathbf{h}}_{u,l}\},b_{3,l}\triangleq|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{h}}_{u,l}|,$ (47)
$b_{1,l}\!\triangleq\!\!{\sum}_{i=1}^{K_{u}}\omega_{u,l}|\beta_{u,l}|^{2}|\hat{\mathbf{u}}_{i}^{H}\hat{\mathbf{h}}_{u,l}|^{2}\!\!+\!\!\!{\sum}_{k=1}^{K_{d}}\omega_{d,k}|\beta_{d,k}|^{2}|{h}_{du,k,l}|^{2}.$
The problem (P4) is an SOCP and can be directly solved by CVX.
III-E Optimizing The UL User Receiver Filter
With the given other variables, the UL user receive filters $\{\hat{\mathbf{u}}_{l}\}$ optimization reduces to solving the following problem
$\textrm{(P5)}:$ $\mathop{\textrm{min}}\limits_{\{\hat{\mathbf{u}}_{l}\}}\ {\sum}_{l=1}^{K_{u}}\{\hat{\mathbf{u}}_{l}^{H}\mathbf{D}_{l}\hat{\mathbf{u}}_{l}-2\textrm{Re}\{\hat{\mathbf{u}}_{l}^{H}\mathbf{d}_{1,l}\}\}+c_{3,1}$ (48a)
where the above newly introduced coefficients are given as
$\mathbf{d}_{1,l}\triangleq\omega_{u,l}\beta_{u,l}^{\ast}\sqrt{q_{l}}\hat{\mathbf{h}}_{u,l},\mathbf{D}_{1}\triangleq\omega_{u,l}|\beta_{u,l}|^{2}\big({\sum}_{i=1}^{K_{u}}q_{i}\hat{\mathbf{h}}_{u,i}\hat{\mathbf{h}}_{u,i}^{H}$ (49)
$\!+\!\!\!\!{\sum}_{j\!=\!0}^{K_{t}}\!\!\sigma_{t,j}^{2}\!(\!{\sum}_{k\!=\!1}^{K_{d}}\!\hat{\mathbf{g}}_{\!r,j}\!\hat{\mathbf{g}}_{\!t,j}^{H}\!\hat{\mathbf{w}}_{k}\!\hat{\mathbf{w}}_{k}^{H}\!\hat{\mathbf{g}}_{\!t,j}\hat{\mathbf{g}}_{\!r,j}^{H}\!\!+\!\!\hat{\mathbf{g}}_{\!r,j}\!\hat{\mathbf{g}}_{\!t,j}^{H}\!\hat{\mathbf{W}}_{\!r}\!\hat{\mathbf{W}}_{\!r}^{H}\!\hat{\mathbf{g}}_{\!t,j}\hat{\mathbf{g}}_{\!r,j}^{H}\!\!+\!\!\sigma_{r}^{2}\mathbf{I}\!\big).$
Since the variables $\{\hat{\mathbf{u}}_{l}\}$ are not intricately coupled in (48a), the problem (P5) can be decomposed into $K_{u}$ independent subproblems, which each subproblem can be written as
$\textrm{(P6${}_{l}$)}:$ $\mathop{\textrm{min}}\limits_{\hat{\mathbf{u}}_{l}}\ \hat{\mathbf{u}}_{l}^{H}\mathbf{D}_{l}\hat{\mathbf{u}}_{l}-2\textrm{Re}\{\hat{\mathbf{u}}_{l}^{H}\mathbf{d}_{1,l}\}$ (50a)
The subproblem (P6l) is a typical unconstrained convex quadratic problem. And then, by setting problem (P6l)’s derivative to zero, the optimal solution can be obtained as
$\hat{\mathbf{u}}_{l}^{\star}=\mathbf{D}_{l}^{-1}\mathbf{d}_{1,l},\forall l\in\mathcal{K}_{u}.$ (51)
III-F Optimizing The Sensing Filter
Given fixed other variables, the problem with respect to (w.r.t.) $\hat{\mathbf{u}}_{0}$ is reduced to a typical feasibility check problem, i.e., Phase-I problem [43], which is given as
$\textrm{(P7)}:$ $\mathop{\textrm{Find}}\limits_{\hat{\mathbf{u}}_{0}}\ \hat{\mathbf{u}}_{0}$ (52a)
$\textrm{s.t.}\ \frac{\hat{\mathbf{u}}_{0}^{H}\mathbf{E}_{1}\hat{\mathbf{u}}_{0}}{\hat{\mathbf{u}}_{0}^{H}\mathbf{E}_{2}\hat{\mathbf{u}}_{0}}\geq\Gamma,$ (52b)
where the newly defined coefficients are given as
$\mathbf{E}_{1}\!\!\triangleq\!\!\sigma_{t,0}^{2}(\!{\sum}_{k=1}^{K_{d}}\!\hat{\mathbf{g}}_{r,0}\hat{\mathbf{g}}_{t,0}^{H}\!\hat{\mathbf{w}}_{k}\!\hat{\mathbf{w}}_{k}^{H}\!\hat{\mathbf{g}}_{t,0}\hat{\mathbf{g}}_{r,0}^{H}\!\!+\!\!\hat{\mathbf{g}}_{r,0}\hat{\mathbf{g}}_{t,0}^{H}\hat{\mathbf{W}}_{r}\!\hat{\mathbf{W}}_{r}^{H}\hat{\mathbf{g}}_{t,0}\hat{\mathbf{g}}_{r,0}^{H}),$
$\mathbf{E}_{2}\!\triangleq\!\!{\sum}_{i=1}^{K_{u}}q_{i}\hat{\mathbf{h}}_{u,i}\hat{\mathbf{h}}_{u,i}^{H}\!\!+\!\!{\sum}_{j=1}^{K_{t}}\!\sigma_{t,j}^{2}\!(\!{\sum}_{k=1}^{K_{d}}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}\hat{\mathbf{g}}_{t,j}\hat{\mathbf{g}}_{r,j}^{H}$
$+\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\!\hat{\mathbf{W}}_{r}^{H}\hat{\mathbf{g}}_{t,j}\hat{\mathbf{g}}_{r,j}^{H})+\sigma_{r}^{2}\mathbf{I}.$ (53)
Since the variable $\hat{\mathbf{u}}_{0}$ only appears in the constraint (52b), we can turn to consider a sensing SINR maximization problem, which is formulated as
$\textrm{(P8)}:$ $\mathop{\textrm{max}}\limits_{\hat{\mathbf{u}}_{0}}\ \frac{\hat{\mathbf{u}}_{0}^{H}\mathbf{E}_{1}\hat{\mathbf{u}}_{0}}{\hat{\mathbf{u}}_{0}^{H}\mathbf{E}_{2}\hat{\mathbf{u}}_{0}}$ (54a)
The above problem (P8) is a typical Rayleigh quotient maximization. By applying the Rayleigh-Ritz theorem, the optimal $\hat{\mathbf{u}}_{0}$ can be obtained by aligning with the eigenvector associated with the largest eigenvalue of the matrix $\mathbf{E}_{2}^{-1}\mathbf{E}_{1}$ [44].
III-G Optimizing The TBS MAs’ Position
In this subsection, when other variables are given, we discuss updating the position of $m$-th TBS’ $n$-th MA, i.e., $\mathbf{t}^{0}_{m,n}$  [^5] To enhance readability, we have omitted the subscripts, i.e., $m$ and $n$, from the variables and/or coefficients. Specifically, we rewrite: $\mathbf{t}^{0}_{m,n}$ as $\mathbf{t}^{0}$, $\mathbf{h}^{0}_{m,k,n}$ as $\mathbf{h}^{0}_{k}$, and $\mathbf{g}_{t,m,j}[n]$ as ${g}_{t,j}$.. Firstly, the function $\tilde{\mathrm{R}}_{d,k}$, which is defined in (31), can be rewritten as
$\!-\!\tilde{\mathrm{R}}_{d,k}\!\!=\!\!(\mathbf{h}^{0}_{k})^{H}\mathbf{{F}}^{1}_{k}\mathbf{h}^{0}_{k}\!\!+\!\!\textrm{Re}\{(\mathbf{f}^{1}_{k})^{H}\mathbf{h}^{0}_{k}\}\!\!-\!\!\textrm{Re}\{(\mathbf{f}^{2}_{k})^{H}\mathbf{h}^{0}_{k}\}\!\!+\!\!c_{4,1,k},$ (55)
where $c_{4,1,k}$ is a constant and the above new coefficients are given as follows
$\mathbf{\bar{F}}^{1}_{m,n,k,i}\triangleq(\mathbf{w}_{m,i}\mathbf{w}_{n,i}^{H})^{T}\otimes((\mathbf{\Sigma}^{0}_{n,k})^{H}\mathbf{h}^{0,0}_{n,k}(\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{\Sigma}^{0}_{n,k}),$ (56)
$\tilde{\mathbf{h}}^{0}_{m,k}\triangleq\textrm{vec}({\mathbf{H}^{0}_{m,k}}),\mathbf{f}_{m,k,j}^{15}\triangleq{\sum}_{j\neq m}^{M_{t}}(\mathbf{\bar{F}}^{1}_{m,j,k,i})^{H}\tilde{\mathbf{h}}^{0}_{j,k},$
$\mathbf{F}^{2}_{m,k}\triangleq(\bar{\mathbf{w}}^{r}_{m}(\bar{\mathbf{w}}^{r}_{m})^{H})^{T}\otimes((\mathbf{\Sigma}^{0}_{m,k})^{H}\mathbf{h}^{0,0}_{m,k}(\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{\Sigma}^{0}_{m,k}),$
$\mathbf{f}^{16}_{m,k}\!\!\triangleq\!\!2\omega_{d,k}\beta_{d,k}(\mathbf{w}_{m,k}^{T}\!\!\otimes\!\!(\!(\mathbf{h}^{0,0}_{m,k})^{H}\!\mathbf{\Sigma}^{0}_{m,k})\!)^{H},\mathbf{F}^{4}_{m,k}\!\!\triangleq\!\omega_{d,k}\!|\!\beta_{d,k}\!|^{2}\!\mathbf{F}^{2}_{m,k},$
$\mathbf{F}^{3}_{m,k}\!\!\triangleq\!\!\!{\sum}_{i=1}^{K_{d}}\omega_{d,k}|\!\beta_{d,k}\!|^{2}\mathbf{\bar{F}}^{1}_{m,m,k,i},\mathbf{f}^{17}_{m,k}\!\!\triangleq\!\!\!{\sum}_{i=1}^{K_{d}}\!2\omega_{d,k}|\!\beta_{d,k}\!|^{2}\!\mathbf{f}_{m,k,i}^{15},$
$\mathbf{F}^{5}_{m,k}\triangleq\mathbf{F}^{3}_{m,k}+\mathbf{F}^{4}_{m,k},\mathbf{f}^{2}_{k}\triangleq\mathbf{f}^{16}_{m,k}[(n-1)L^{t,0}_{m,k}+1:nL^{t,0}_{m,k}],$
${\mathbf{F}}^{1}_{m,k,i,j}\!\triangleq\!\mathbf{F}^{5}_{m,k}[(i-1)L^{t,0}_{m,k}\!+\!1\!:\!iL^{t,0}_{m,k},(j-1)L^{t,0}_{m,k}\!+\!1\!:\!jL^{t,0}_{m,k}],$
${\mathbf{F}}^{1}_{k}\!\triangleq\!\mathbf{F}^{5}_{m,k}[(n-1)L^{t,0}_{m,k}\!+\!1\!:\!nL^{t,0}_{m,k},(n-1)L^{t,0}_{m,k}\!+\!1\!:\!nL^{t,0}_{m,k}],$
$\mathbf{f}^{18}_{m,k,n}\triangleq 2({\sum}_{i\neq n}^{N_{t}}({\mathbf{\bar{F}}}^{1}_{m,k,i,n})^{H}\mathbf{h}^{0}_{m,k,i}),$
$\mathbf{f}^{1}_{k}\triangleq\mathbf{f}^{18}_{m,k,n}+\mathbf{f}^{17}_{m,k}[(n-1)L^{t,0}_{m,k}+1:nL^{t,0}_{m,k}].$
Furthermore, the function $\tilde{\mathrm{R}}_{u,l}$ defined in (32) is also rewritten as follows
${\sum}_{l=1}^{K_{u}}\tilde{\mathrm{R}}_{u,l}\!={\sum}_{j=0}^{K_{t}}-\textrm{Re}\{(f^{3}_{j})^{\ast}{g}_{t,j}\}+c_{4,2},$ (57)
where $c_{4,2}$ is a constant term and the new parameters in (57) are defined as follows
$\mathbf{F}^{7}_{l,j}\triangleq(\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{l}\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j})^{T}\otimes(\hat{\mathbf{W}}_{r}\hat{\mathbf{W}}_{r}^{H}),$ (58)
$\mathbf{F}^{6}_{l,j,k}\triangleq(\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{l}\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j})^{T}\otimes(\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}),$
$\mathbf{F}^{8}_{l,j}\triangleq{\sum}_{k=1}^{K_{d}}\mathbf{F}^{6}_{l,j,k}+\mathbf{F}^{7}_{l,j},\mathbf{F}^{9}_{j}\triangleq{\sum}_{l=1}^{K_{u}}\omega_{u,l}|\beta_{u,l}|^{2}\sigma_{t,j}^{2}\mathbf{F}^{8}_{l,j},$
$\bar{\mathbf{F}}^{9}_{j,m,n}\triangleq\mathbf{F}^{9}_{j}[(m-1)N_{t}+1:mN_{t},(n-1)N_{t}+1:nN_{t}],$
$f^{20}_{m,j,n}\triangleq{\sum}_{i\neq n}^{N_{t}}(\bar{\mathbf{F}}^{9}_{j,m,m}[i,n])^{\ast}\mathbf{g}_{t,m,j}[i],$
$\mathbf{f}^{19}_{m,j}\triangleq{\sum}_{i\neq m}^{M_{t}}2(\bar{\mathbf{F}}^{9}_{j,m,n})^{H}\mathbf{g}_{t,i,j},f^{3}_{j}\triangleq f^{20}_{m,j,n}+\mathbf{f}^{19}_{m,j}[n].$
Similarly, the sensing SINR constraint (30b) is rewritten as
$\textrm{SINR}_{t}\geq\Gamma_{r}\Longleftrightarrow$ (59)
${\sum}_{i=1}^{K_{t}}\textrm{Re}\{(f^{4}_{i})^{\ast}{g}_{t,i}\}-\textrm{Re}\{(f^{5})^{\ast}\mathbf{g}_{t,0}\}+c_{4,3}\leq 0,$
where $c_{4,3}$ is constant and the newly introduced coefficients are defined as follows
$\mathbf{F}^{10}_{k}\triangleq\sigma_{t,0}^{2}(\hat{\mathbf{g}}_{r,0}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,0})^{T}\otimes(\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}),$ (60)
$\mathbf{F}^{11}\triangleq\sigma_{t,0}^{2}(\hat{\mathbf{g}}_{r,0}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,0})^{T}\otimes(\hat{\mathbf{W}}_{r}\hat{\mathbf{W}}_{r}^{H}),$
$\mathbf{F}^{13}_{k,j}\triangleq\sigma_{t,j}^{2}(\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,j})^{T}\otimes(\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}),$
$\mathbf{F}^{14}_{j}\triangleq\sigma_{t,j}^{2}(\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,j})^{T}\otimes(\hat{\mathbf{W}}_{r}\hat{\mathbf{W}}_{r}^{H}),$
$\mathbf{F}^{12}\triangleq({\sum}_{k=1}^{K_{d}}\mathbf{F}^{10}_{k}+\mathbf{F}^{11})/\Gamma,\mathbf{F}^{15}_{j}\triangleq{\sum}_{k=1}^{K_{d}}\mathbf{F}^{13}_{k,j}+\mathbf{F}^{14}_{j},$
$\bar{\mathbf{F}}^{15}_{j,m,n}\triangleq\mathbf{F}^{15}_{j}[(m-1)N_{t}+1:mN_{t},(n-1)N_{t}+1:nN_{t}],$
$f^{22}_{j,m,n}\triangleq{\sum}_{z\neq n}^{N_{t}}2(\bar{\mathbf{F}}^{15}_{j,m,m}[z,n])^{\ast}\mathbf{g}_{t,m,j}[z],$
$\bar{\mathbf{F}}^{12}_{m,n}\triangleq\mathbf{F}^{12}[(m-1)N_{t}+1:mN_{t},(n-1)N_{t}+1:nN_{t}],$
${f}^{24}_{m,n}\triangleq{\sum}_{i\neq n}^{N_{t}}2(\bar{\mathbf{F}}^{12}_{m,m}[i,n])^{\ast}\mathbf{g}_{t,m,0}[i],$
$\mathbf{f}^{21}_{j,m}\triangleq{\sum}_{i\neq m}^{M_{t}}2(\bar{\mathbf{F}}^{15}_{j,i,m})^{H}\mathbf{g}_{t,i,j},\mathbf{f}^{23}_{m}\triangleq{\sum}_{n\neq m}^{M_{t}}2(\bar{\mathbf{F}}^{12}_{m,n})^{H}\mathbf{g}_{t,n,0},$
$f^{4}_{j}\triangleq f^{22}_{j,m,n}+\mathbf{f}^{21}_{j,m}[n],{f}^{5}\triangleq{f}^{24}_{m,n}+\mathbf{f}^{23}_{m}[n].$
Based on the above equivalent transformation, the subproblem w.r.t $\mathbf{t}^{0}$ is formulated as
$\textrm{(P9)}:$ $\mathop{\textrm{min}}\limits_{\mathbf{t}^{0}}{\sum}_{k=1}^{K_{d}}\big((\mathbf{h}^{0}_{k}(\mathbf{t}^{0}))^{H}\mathbf{{F}}^{1}_{k}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})$
$+\textrm{Re}\{(\mathbf{f}^{1}_{k})^{H}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})\}-\textrm{Re}\{(\mathbf{f}^{2}_{k})^{H}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})\}\big)$
$+{\sum}_{j=0}^{K_{t}}\textrm{Re}\{({f}^{3}_{j})^{\ast}{g}_{t,j}(\mathbf{t}^{0})\}+{c}_{4,4}$ (61a)
s.t. ${\sum}_{i=1}^{K_{t}}\textrm{Re}\{(f^{4}_{i})^{\ast}{g}_{t,i}(\mathbf{t}^{0})\}$ (61b)
$-\textrm{Re}\{(f^{5})^{\ast}{g}_{t,0}(\mathbf{t}^{0})\}+c_{4,3}\leq 0,$
$\|\mathbf{t}^{0}-\mathbf{t}^{0}_{m,i}\|_{2}\geq D_{t},\forall n,i\in\mathcal{N}_{t},n\neq i,$ (61c)
$\mathbf{t}^{0}\in\mathcal{C},$ (61d)
where ${c}_{4,4}$ is a constant. Obviously, the current form of problem (P9) is difficult to obtain the optimal $\mathbf{t}^{0}$ due to the non-convex objective function and constraints. In the following, we adopt the MM methodology [35] to resolve the above difficulty. Firstly, for any positive semidefinite matrix $\mathbf{F}$, we have the following inequality
$\mathbf{h}^{H}(\mathbf{t}^{0})\mathbf{F}\mathbf{h}(\mathbf{t}^{0})$ (62)
$=(\mathbf{h}\!-\!\mathbf{h}_{0})^{H}\mathbf{F}(\mathbf{h}\!-\!\mathbf{h}_{0})+2\mathrm{Re}\{\mathbf{h}_{0}^{H}\mathbf{F}(\mathbf{h}\!-\!\mathbf{h}_{0})\}+\mathbf{h}_{0}^{H}\mathbf{F}\mathbf{h}_{0}$
$\leq\lambda_{max}(\mathbf{F})\|\mathbf{h}-\mathbf{h}_{0}\|^{2}_{2}+2\mathrm{Re}\{\mathbf{h}_{0}^{H}\mathbf{F}(\mathbf{h}-\mathbf{h}_{0})\}+\mathbf{h}_{0}^{H}\mathbf{F}\mathbf{h}_{0}$
$=\lambda_{max}(\mathbf{F})\|\mathbf{h}\|^{2}_{2}+2\mathrm{Re}\{(\mathbf{F}\mathbf{h}_{0}-\lambda_{max}(\mathbf{F})\mathbf{h}_{0})^{H}\mathbf{h}\}+c_{F},$
with $\lambda_{max}(\mathbf{F})$ being the maximal eigenvalue of the matrix $\mathbf{F}$, $c_{F}$ is a constant, $\mathbf{h}_{0}\triangleq\mathbf{h}(\mathbf{t}^{0}_{0})$, $\mathbf{t}^{0}_{0}=[x^{0}_{0},y^{0}_{0}]^{T}$ is the latest value of $\mathbf{t}^{0}$.
Next, by leveraging the arguments presented in (62), we can construct a tight upper-bound for $(\mathbf{h}^{0}_{k}(\mathbf{t}^{0}))^{H}\mathbf{{F}}^{1}_{k}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})$, which can be formulated in (63), where $\mathbf{h}^{0}_{k,0}\triangleq\mathbf{h}^{0}_{k}(\mathbf{t}^{0}_{0})$, $\|\mathbf{h}^{0}_{k}\|^{2}_{2}=\|\mathbf{h}^{0}_{k,0}\|^{2}_{2}=L^{t,0}_{m,k}$, $c_{4,5,k}\triangleq 2\lambda_{max}(\mathbf{{F}}^{1}_{k})L^{t,0}_{m,k}-((\mathbf{h}^{0}_{k,0})^{H}\mathbf{{F}}^{1}_{k}\mathbf{h}^{0}_{k,0})^{\ast}$, $\mathbf{f}^{6}_{k}\!\triangleq\!2\big((\mathbf{{F}}^{1}_{k})^{H}\mathbf{h}^{0}_{k,0}\!-\!\lambda_{max}(\mathbf{{F}}^{1}_{k})\mathbf{h}^{0}_{k,0}\big)$ and $\mathbf{f}^{7}_{k}\!\triangleq\!\mathbf{f}^{6}_{k}\!+\!\mathbf{f}^{1}_{k}$.
$(\mathbf{h}^{0}_{k}(\mathbf{t}^{0}))^{H}\mathbf{{F}}^{1}_{k}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})+\textrm{Re}\{(\mathbf{f}^{1}_{k})^{H}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})\}$ (63)
$\leq\lambda_{max}(\mathbf{{F}}^{1}_{k})\|\mathbf{h}^{0}_{k}\|^{2}_{2}-2\textrm{Re}\{\lambda_{max}(\mathbf{{F}}^{1}_{k})(\mathbf{h}^{0}_{k,0})^{H}\mathbf{h}^{0}_{k}\}+\lambda_{max}(\mathbf{{F}}^{1}_{k})\|\mathbf{h}^{0}_{k,0}\|^{2}_{2}$
$+2\textrm{Re }\{(\mathbf{h}^{0}_{k,0})^{H}\mathbf{{F}}^{1}_{k}\mathbf{h}^{0}_{k}\}-2\textrm{Re }\{(\mathbf{h}^{0}_{k,0})^{H}\mathbf{{F}}^{1}_{k}\mathbf{h}^{0}_{k,0}\}+(\mathbf{h}^{0}_{k,0})^{H}\mathbf{{F}}^{1}_{k}\mathbf{h}^{0}_{k,0}+\textrm{Re}\{(\mathbf{f}^{1}_{k})^{H}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})\}$
$=\textrm{Re}\{(\mathbf{f}^{6}_{k})^{H}\mathbf{h}^{0}_{k}\}+c_{4,5,k}+\textrm{Re}\{(\mathbf{f}^{1}_{k})^{H}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})\}=\textrm{Re}\{(\mathbf{f}^{7}_{k})^{H}\mathbf{h}^{0}_{k}\}+c_{4,5,k}$
Furthermore, to convexity the challenging terms $\textrm{Re}\{(\mathbf{f}^{7}_{k})^{H}\mathbf{h}^{0}_{k}\}$, $\textrm{Re}\{(\mathbf{f}^{2}_{k})^{H}\mathbf{h}^{0}_{k}\}$ and $\textrm{Re}\{({f}^{3}_{j})^{\ast}{g}_{t,j}(\mathbf{t}^{0})\}$, we introduce the following results that are proved in [4].
Lemma 1.
Assume that $\mathbf{f}\in\mathbb{C}^{L\times 1}$, $\mathbf{h}(\mathbf{t})=[e^{j\!\frac{2\pi}{\lambda}\!(\cos(\!\theta_{1}\!)\sin(\!\phi_{1}\!)x\!+\!\sin(\theta_{1})y)},\cdots,e^{j\!\frac{2\pi}{\lambda}\!(\cos(\!\theta_{L}\!)\sin(\!\phi_{L}\!)x\!+\!\sin(\theta_{L})y)}]\in\mathbb{C}^{L\times 1}$ and $\mathbf{t}=[x,y]^{T}$. The convex lower and upper bounds of the function $h_{t}(\mathbf{t})\triangleq\textrm{Re}\{\mathbf{f}^{H}\mathbf{h}(\mathbf{t})\}=\sum_{l=1}^{L}|\mathbf{f}[l]|\cos(\angle\mathbf{h}(\mathbf{t})[l]-\angle\mathbf{f}[l])$ ( the variable $\mathbf{t}$ only exists in the phase part of $h_{t}(\mathbf{t})$ ) can be written as
$h_{t}(\mathbf{t})\!\geq\!h_{t}(\mathbf{t}_{0})\!\!+\!\nabla h_{t}^{T}(\mathbf{t}_{0})(\mathbf{t}\!-\!\mathbf{t}_{0})\!-\!{\delta}/{2}(\mathbf{t}\!-\!\mathbf{t}_{0})^{T}(\mathbf{t}\!-\!\mathbf{t}_{0}),$ (64)
$h_{t}(\mathbf{t})\!\leq\!h_{t}(\mathbf{t}_{0})\!\!+\!\nabla h_{t}^{T}(\mathbf{t}_{0})(\mathbf{t}\!-\!\mathbf{t}_{0})\!+\!{\delta}/{2}(\mathbf{t}\!-\!\mathbf{t}_{0})^{T}(\mathbf{t}\!-\!\mathbf{t}_{0}),$ (65)
respectively, where $\mathbf{t}_{0}$ is obtained from the last iteration, the symbols |||\ | and $\angle$ denote the operations of obtaining the amplitude and phase, respectively, $\delta$ is a positive real number satisfying $\delta=\frac{8\pi^{2}}{\lambda^{2}}{\sum}_{l=1}^{L}|\mathbf{f}[l]|$, $\nabla h_{t}(\mathbf{t})=\big[\frac{\partial h_{t}(\mathbf{t})}{\partial x},\frac{\partial h_{t}(\mathbf{t})}{\partial y}\big]^{T}\in\mathbb{C}^{2\times 1}$ denotes the gradient vector of $h_{t}(\mathbf{t})$ over $\mathbf{t}$.
Firstly, by using the inequality (65) in Lemma 1, we can obtain a convex upper-bound of $\textrm{Re}\{(\mathbf{f}^{7}_{k})^{H}\mathbf{h}^{0}_{k}\}$, showing as
$\textrm{Re}\{(\mathbf{f}^{7}_{k})^{H}\mathbf{h}^{0}_{k}\}\triangleq h^{1}_{k}(\mathbf{t}^{0})$ (66)
$={\sum}_{l=1}^{L^{t,0}_{m,k}}|\mathbf{f}^{7}_{k}[l]|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{0})^{T}\mathbf{a}^{t,0}_{m,k,l}-\angle\mathbf{f}^{7}_{k}[l]\big)$
$\leq h^{1}_{k}(\mathbf{t}^{0}_{0})+\nabla^{T}h^{1}_{k}(\mathbf{t}^{0}_{0})(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})+0.5{\delta_{1,k}}(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})^{T}(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})$
$=0.5{\delta_{1,k}}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{8}_{k})^{T}\mathbf{t}^{0}+c_{4,6,k},$
where $c_{4,6,k}$ is a constant term, $\delta_{1,k}=\frac{8\pi^{2}}{\lambda^{2}}{\sum}_{l=1}^{L^{t,0}_{m,k}}|\mathbf{f}^{7}_{k}[l]|$, $\mathbf{f}^{8}_{k}\triangleq\nabla h^{1}_{k}(\mathbf{t}^{0}_{0})-\delta_{1,k}\mathbf{t}^{0}_{0}$, and the elements of the gradient vector $\nabla h^{1}_{k}(\mathbf{t}^{0}_{0})=\big[\frac{\partial h^{1}_{k}(\mathbf{t}^{0}_{0})}{\partial x^{0}_{0}},\frac{\partial h^{1}_{k}(\mathbf{t}^{0}_{0})}{\partial y^{0}_{0}}\big]^{T}$ are respectively given as  [^6] In the following, when applying the equations (64) and/or (65), the derivation details of the gradient vector can refer to (67) and (68).
$\frac{\partial h^{1}_{k}(\mathbf{t}^{0}_{0})}{\partial x^{0}_{0}}=-{\sum}_{l=1}^{L^{t,0}_{m,k}}\frac{2\pi}{\lambda}|\mathbf{f}^{7}_{k}[l]|\cos(\theta^{t,0}_{m,k,l})$ (67)
$\sin(\phi^{t,0}_{m,k,l})\sin\big(\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{0})^{T}\mathbf{a}^{t,0}_{m,k,l}-\angle\mathbf{f}^{7}_{k}[l]\big),$
$\frac{\partial h^{1}_{k}(\mathbf{t}^{0}_{0})}{\partial y^{0}_{0}}=-{\sum}_{l=1}^{L^{t,0}_{m,k}}\frac{2\pi}{\lambda}|\mathbf{f}^{7}_{k}[l]|\sin(\theta^{t,0}_{m,k,l})$ (68)
$\sin\big(\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{0})^{T}\mathbf{a}^{t,0}_{m,k,l}-\angle\mathbf{f}^{7}_{k}[l]\big).$
Furthermore, the non-convex term $\textrm{Re}\{(\mathbf{f}^{2}_{k})^{H}\mathbf{h}^{0}_{k}\}$ can be lower-bounded by the inequality (64), which is given as
$\textrm{Re}\{(\mathbf{f}^{2}_{k})^{H}\mathbf{h}^{0}_{k}\}\triangleq h^{2}_{k}(\mathbf{t}^{0})$ (69)
$={\sum}_{l=1}^{L^{t,0}_{m,k}}|\mathbf{f}^{2}_{k}[l]|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{0})^{T}\mathbf{a}^{t,0}_{m,k,l}-\angle\mathbf{f}^{2}_{k}[l]\big)$
$\geq h^{2}_{k}(\mathbf{t}^{0}_{0})+\nabla^{T}h^{2}_{k}(\mathbf{t}^{0}_{0})(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})-0.5{\bar{\delta}_{1,k}}(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})^{T}(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})$
$=-0.5{\bar{\delta}_{1,k}}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{{f}}^{9}_{k})^{T}\mathbf{t}^{0}+{c}_{4,7,k},$
where ${c}_{4,7,k}$ is a constant term, $\bar{\delta}_{1,k}=\frac{8\pi^{2}}{\lambda^{2}}{\sum}_{l=1}^{L^{t,0}_{m,k}}|\mathbf{f}^{2}_{k}[l]|$, $\mathbf{{f}}^{9}_{k}\triangleq\nabla h^{2}_{k}(\mathbf{t}^{0}_{0})+\bar{\delta}_{1,k}\mathbf{t}^{0}_{0}$, $\nabla h^{2}_{k}(\mathbf{t}^{0}_{0})$ is the gradient vector.
Next, the term $\textrm{Re}\{({f}^{3}_{j})^{\ast}{g}_{t,j}(\mathbf{t}^{0})\}$ be convexified by leveraging (65) as follows
$\textrm{Re}\{({f}^{10}_{j})^{\ast}{g}_{t,j}(\mathbf{t}^{0})\}\triangleq g^{1}_{j}(\mathbf{t}^{0})$ (70)
$=|{f}^{3}_{j}||{g}_{t,j}|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{0})^{T}\mathbf{a}^{4}_{m,j}-\angle{f}^{3}_{j}\big)$
$\leq g^{1}_{j}(\mathbf{t}^{0}_{0})+\nabla^{T}g^{1}_{j}(\mathbf{t}^{0}_{0})(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})+0.5{\delta_{2,j}}(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})^{T}(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})$
$=0.5{\delta_{2,j}}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{10}_{j})^{T}\mathbf{t}^{0}+c_{4,8,j},$
with $c_{4,8,j}$ being constant, $\delta_{2,j}\triangleq\frac{8\pi^{2}}{\lambda^{2}}|{f}^{3}_{j}||\mathbf{g}_{t,j}|$, $\mathbf{f}^{10}_{j}\triangleq\nabla g^{1}_{j}(\mathbf{t}^{0}_{0})-\delta_{2,j}\mathbf{t}^{0}_{0}$, $\nabla g^{1}_{j}(\mathbf{t}^{0}_{0})$ is the gradient vector over $\mathbf{t}^{0}_{0}$.
Based on the above transformation, the upper-bound of the objective function (61a) is given as
${\sum}_{k=1}^{K_{d}}\big((\mathbf{h}^{0}_{k}(\mathbf{t}^{0}))^{H}\mathbf{{F}}_{1,k}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})+\textrm{Re}\{(\mathbf{f}^{1}_{k})^{H}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})\}$ (71)
$-\textrm{Re}\{(\mathbf{f}^{2}_{k})^{H}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})\}\big)+{\sum}_{j=0}^{K_{t}}\textrm{Re}\{({f}^{3}_{j})^{\ast}{g}_{t,j}(\mathbf{t}^{0})\}+\hat{c}_{4,2}$
$\leq{\sum}_{k=1}^{K_{d}}\big(0.5{\delta_{1,k}}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{8}_{k})^{T}\mathbf{t}^{0}+c_{4,6,k}$
$-(-{0.5\bar{\delta}_{1,k}}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{{f}}^{9}_{k})^{T}\mathbf{t}^{0}+{c}_{4,7,k})\big)$
$+{\sum}_{j=0}^{K_{t}}({0.5\delta_{2,j}}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{10}_{j})^{T}\mathbf{t}^{0}+c_{4,8,j})+{c}_{4,4}$
$={\delta_{3}}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{11})^{T}\mathbf{t}^{0}+c_{4,9}$
where $c_{4,9}$ is a constant term, $\delta_{3}\triangleq\sum_{k=1}^{K_{d}}({\delta_{1,k}}/{2}+{\bar{\delta}_{1,k}}/{2})+\sum_{j=0}^{K_{t}}{\delta_{2,j}}/{2}$ and $\mathbf{f}^{11}\triangleq\sum_{k=1}^{K_{d}}(\mathbf{f}^{8}_{k}-\mathbf{{f}}^{9}_{k})+\sum_{j=0}^{K_{t}}\mathbf{f}^{10}_{j}$.
In the following, the term $\textrm{Re}\{(f^{13}_{i})^{\ast}{g}_{t,i}(\mathbf{t}^{0})\}$ of the sensing SINR constraint (61b) can be convexified via (65) as follows
$\textrm{Re}\{(f^{4}_{i})^{\ast}\mathbf{g}_{t,i}(\mathbf{t}^{0})\}\triangleq g^{2}_{i}(\mathbf{t}^{0})$ (72)
$=|f^{4}_{i}||\mathbf{g}_{t,i}|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{0})^{T}\mathbf{a}^{4}_{m,i}-\angle f^{4}_{i}\big)$
$\leq g^{2}_{i}(\mathbf{t}^{0}_{0})+\nabla^{T}g^{2}_{i}(\mathbf{t}^{0}_{0})(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})+0.5{\delta_{4,i}}(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})^{T}(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})$
$={0.5\delta_{4,i}}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{12}_{i})^{T}\mathbf{t}^{0}+c_{4,10,i},$
with $c_{4,10,i}$ being constant, $\delta_{4,i}\triangleq\frac{8\pi^{2}}{\lambda^{2}}|f^{4}_{i}||\mathbf{g}_{t,i}|$, $\mathbf{f}^{12}_{i}\triangleq\nabla g^{2}_{i}(\mathbf{t}^{0}_{0})-\delta_{4,i}\mathbf{t}^{0}_{0}$, $\nabla g^{2}_{i}(\mathbf{t}^{0}_{0})$ is the gradient vector.
And then, we proceed to cope with the non-convex term $\textrm{Re}\{(f^{5})^{\ast}{g}_{t,0}(\mathbf{t}^{0})\}$ of the sensing SINR constraint (61b). This time, we require a convex lower bound for $\textrm{Re}\{(f^{5})^{\ast}{g}_{t,0}(\mathbf{t}^{0})\}$ by applying (64) in Lemma 1, which is given as follows
$\textrm{Re}\{(f^{5})^{\ast}{g}_{t,0}(\mathbf{t}^{0})\}\triangleq g^{3}(\mathbf{t}^{0})$ (73)
$=|f^{5}||\mathbf{g}_{t,0}|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{0})^{T}\mathbf{a}^{4}_{m,0}-\angle f^{5}\big)$
$\geq g^{3}(\mathbf{t}^{0}_{0})+\nabla^{T}g^{3}(\mathbf{t}^{0}_{0})(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})-{0.5\delta_{6}}(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})^{T}(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})$
$=-{0.5\delta_{6}}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{13})^{T}\mathbf{t}^{0}+c_{4,11},$
where $c_{4,11}$ is constant, $\delta_{6}\triangleq\frac{8\pi^{2}}{\lambda^{2}}|f^{5}||\mathbf{g}_{t,0}|$, $\mathbf{f}^{13}\triangleq\nabla g^{3}(\mathbf{t}^{0}_{0})+\delta_{6}\mathbf{t}^{0}_{0}$, $\nabla g^{3}(\mathbf{t}^{0}_{0})$ is the gradient vector.
Therefore, a convex upper-bound of the SINR constraint (61b) can be written as
${\sum}_{i=1}^{K_{t}}\textrm{Re}\{(f^{4}_{i})^{\ast}{g}_{t,i}(\mathbf{t}^{0})\}-\textrm{Re}\{(f^{5})^{\ast}{g}_{t,0}(\mathbf{t}^{0})\}+c_{4,3}$ (74)
$\leq{\sum}_{i=1}^{K_{t}}({\delta_{4,i}}/{2}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{12}_{i})^{T}\mathbf{t}^{0}+c_{4,10,i})$
$-(-{\delta_{6}}/{2}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{13})^{T}\mathbf{t}^{0}+c_{4,11})+c_{4,3}$
$=\delta_{7}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{14})^{T}\mathbf{t}^{0}+c_{4,12},$
where $c_{4,12}$ is constant term, $\delta_{7}\triangleq\sum_{j=1}^{K_{t}}{\delta_{4,j}}/{2}+{\delta_{6}}/{2}$ and $\mathbf{f}^{14}\triangleq\sum_{j=1}^{K_{t}}\mathbf{f}^{12}_{j}-\mathbf{f}^{13}$.
To make (P9) tractable, we still need to convexify the constraint (61c). Inspired by the MM framework [35], the constraint (61c) can be linearized as follows
$\|\mathbf{t}^{0}-\mathbf{t}^{0}_{m,i}\|_{2}\geq\frac{(\mathbf{t}^{0}_{0}-\mathbf{t}^{0}_{m,i})^{T}}{\|\mathbf{t}^{0}_{0}-\mathbf{t}^{0}_{m,i}\|_{2}}(\mathbf{t}^{0}-\mathbf{t}^{0}_{m,i}).$ (75)
Based on the above MM transformation, the update of $\mathbf{t}^{0}$ can be performed by solving a convex surrogate of the original challenging problem (P9), as follows
$\textrm{(P10)}:$ $\mathop{\textrm{min}}\limits_{\mathbf{t}^{0}}\delta_{3}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{11})^{T}\mathbf{t}^{0}_{m,n}+c_{4,9}$ (76a)
s.t. $\delta_{7}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{14})^{T}\mathbf{t}^{0}+c_{4,12}\leq 0,$ (76b)
$\frac{(\mathbf{t}^{0}_{0}-\mathbf{t}^{0}_{m,i})^{T}}{\|\mathbf{t}^{0}_{0}\!\!-\!\mathbf{t}^{0}_{m,i}\|_{2}}\!(\mathbf{t}^{0}\!-\!\mathbf{t}^{0}_{m,i})\geq D_{t},\forall i\in\mathcal{N}_{t},n\neq i,$ (76c)
$\mathbf{t}^{0}\in\mathcal{C},$ (76d)
Clearly, problem (P10) is convex and can be numerically solved.
III-H Optimizing The RBS MAs’ Position
In this subsection, with other variables being given, we investigate the update of the position of the $m$-th RBS’ $n$-th MA, i.e., $\mathbf{t}^{1}_{m,n}$  [^7] To improve readability, we have removed the subscripts, i.e., $m$ and $n$, from the variables and/or coefficients. Specifically, we represent: $\mathbf{t}^{1}_{m,n}$ as $\mathbf{t}^{1}$, $\mathbf{h}^{1}_{m,l,n}$ as $\mathbf{h}^{1}_{l}$, and $\mathbf{g}_{r,m,j}[n]$ as ${g}_{r,j}$. .
Firstly, the function $\tilde{\mathrm{R}}_{u,l}$ given in (32) is rewritten as
$-{\sum}_{l=1}^{K_{u}}\tilde{\mathrm{R}}_{u,l}={\sum}_{l=1}^{K_{u}}\big((\mathbf{h}^{1}_{l})^{H}\mathbf{D}^{1}_{l}\mathbf{h}^{1}_{l}+\textrm{Re}\{(\mathbf{d}^{1}_{l})^{H}\mathbf{h}^{1}_{l}\}$ (77)
$-\textrm{Re}\{(\mathbf{{d}}^{2}_{l})^{H}\mathbf{h}^{1}_{l}\}\big)+{\sum}_{j=0}^{K_{t}}\textrm{Re}\{(d^{3}_{j})^{\ast}\mathbf{g}_{r,j}\}+c_{5,1},$
with $c_{5,1}$ being a constant and the parameters in (77) defining in following
$\mathbf{\bar{D}}^{1}_{m,n,l,i}\!\!\triangleq\!\omega_{u,l}|\beta_{u,l}|^{2}\!(\mathbf{u}_{m,l}\mathbf{u}_{n,l}^{H})^{T}\!\!\!\otimes\!\!(q_{i}(\mathbf{\Sigma}^{1}_{n,i})^{H}\!\mathbf{h}^{0,1}_{n,i}\!\mathbf{h}^{0,1}_{m,i}\mathbf{\Sigma}^{1}_{m,i}),$ (78)
$\hat{\mathbf{h}}^{1}_{m,l}\triangleq\textrm{vec}(\mathbf{H}^{1}_{m,l}(\mathbf{t}^{1}_{m})),\mathbf{d}^{19}_{m,l,i}\triangleq 2({\sum}_{j\neq m}^{M_{t}}(\mathbf{\bar{D}}^{1}_{m,j,l,i})^{H}\hat{\mathbf{h}}^{1}_{j,i}),$
$\mathbf{d}^{20}_{m,l}\triangleq 2\omega_{u,l}\beta_{u,l}\sqrt{q_{l}}\mathbf{u}_{m,l}^{H}\otimes((\mathbf{h}^{0,1}_{m,l})^{H}\mathbf{\Sigma}^{1}_{m,l}),$
$\mathbf{\bar{D}}^{2}_{j,l}\triangleq{\sum}_{k=1}^{K_{d}}[(\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}\hat{\mathbf{g}}_{t,j})^{T}\otimes\sigma^{2}_{t,j}(\hat{\mathbf{u}}_{l}\hat{\mathbf{u}}_{l}^{H})],$
$\mathbf{D}^{4}_{j,l}\triangleq\omega_{u,l}|\beta_{u,l}|^{2}(\mathbf{\bar{D}}^{2}_{j,l}+\mathbf{D}^{3}_{j,l}),$
$\mathbf{D}^{3}_{j,l}\triangleq(\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}^{r}(\hat{\mathbf{W}}^{r})^{H}\hat{\mathbf{g}}_{t,j})^{T}\otimes\sigma^{2}_{t,j}(\hat{\mathbf{u}}_{l}\hat{\mathbf{u}}_{l}^{H}),$
$\bar{\mathbf{D}}^{4}_{j,l,m,n}\triangleq\mathbf{D}^{4}_{j,l}[(m-1)N_{r}+1:mN_{r},(n-1)N_{r}+1:nN_{r}],$
$\mathbf{d}^{21}_{j,l,m}\triangleq 2({\sum}_{i\neq m}^{M_{r}}(\bar{\mathbf{D}}^{4}_{j,l,i,m})^{H}\mathbf{g}_{r,i,j}),\mathbf{D}^{5}_{m,i}\triangleq{\sum}_{l=1}^{K_{u}}\mathbf{\bar{D}}^{1}_{m,m,l,i},$
$\mathbf{d}^{22}_{m,i}\!\triangleq\!\!{\sum}_{l=1}^{K_{u}}\mathbf{d}^{19}_{m,l,i},\mathbf{D}^{6}_{j,m}\!\triangleq\!\!{\sum}_{l=1}^{K_{u}}\bar{\mathbf{D}}^{4}_{j,l,m,m},\mathbf{d}^{23}_{j,m}\!\triangleq\!\!{\sum}_{l=1}^{K_{u}}\mathbf{d}^{21}_{j,l,m},$
$\mathbf{D}^{1}_{l}\triangleq\mathbf{D}^{5}_{m,l}[(n-1)L^{t,1}_{m,l}+1:nL^{t,1}_{m,l},(j-1)L^{t,1}_{m,l}+1:jL^{t,1}_{m,l}],$
$\mathbf{d}^{24}_{m,l,n}\triangleq 2({\sum}_{j\neq n}^{N_{r}}(\mathbf{D}^{7}_{m,l,n,j})^{H}\mathbf{h}^{1}_{m,l,n}),$
$\mathbf{d}^{1}_{l}\!\triangleq\!\mathbf{d}^{24}_{m,l,n}\!\!+\!\mathbf{d}^{22}_{m,l}[(n\!-\!1)L^{t,1}_{m,l}\!+\!1\!:\!nL^{t,1}_{m,l}],$
$d^{3}_{j}\!\triangleq\!2({\sum}_{i\neq n}^{N_{r}}(\mathbf{D}^{6}_{j,m}[i,n])^{\ast}{\mathbf{g}}_{r,m,j}[i])\!+\!\mathbf{d}^{23}_{j,m}[n],$
$\mathbf{{d}}^{2}_{l}\!\!\triangleq\!\mathbf{d}^{20}_{m,l}[(n\!-\!1)L^{t,1}_{m,l}+1\!:\!nL^{t,1}_{m,l}].$
And the sensing SINR constraint (30b) can be rewritten as
$\textrm{SINR}_{t}\geq\Gamma_{r}\!\!\!\Longleftrightarrow\!\!\!{\sum}_{l=1}^{K_{u}}\big((\mathbf{h}^{1}_{l})^{H}\mathbf{D}^{2}_{l}\mathbf{h}^{1}_{l}\!+\!\!\textrm{Re}\{(\mathbf{d}^{4}_{l})^{H}\mathbf{h}^{1}_{l}\}\big)$ (79)
$+{\sum}_{j=1}^{K_{t}}\big(\textrm{Re}\{(d^{5}_{j})^{\ast}\mathbf{g}_{r,j}\}\big)-\textrm{Re}\{(d^{6})^{\ast}{g}_{r,0}\}+c_{5,2}\leq 0,$
where $c_{5,2}$ is a constant term and the above newly introduced coefficients are given in
$\mathbf{D}^{10}\triangleq\big({\sum}_{k=1}^{K_{d}}[(\hat{\mathbf{g}}_{t,0}^{H}\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}\hat{\mathbf{g}}_{t,0})^{T}\otimes\sigma^{2}_{t,0}(\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H})]$ (80)
$+(\hat{\mathbf{g}}_{t,0}^{H}\hat{\mathbf{W}}^{r}(\hat{\mathbf{W}}^{r})^{H}\hat{\mathbf{g}}_{t,0})^{T}\otimes\sigma^{2}_{t,0}(\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H})\big)/\Gamma_{t},$
$\bar{\mathbf{D}}^{10}_{n,m}\triangleq\mathbf{D}^{10}[(n-1)N_{r}+1:nN_{r},(m-1)N_{r}+1:mN_{r}],$
$\mathbf{D}^{11}_{l,m,n}\triangleq(\mathbf{u}_{m,0}\mathbf{u}_{n,0}^{H})^{T}\otimes(q_{l}(\mathbf{\Sigma}^{1}_{n,l})^{H}\mathbf{h}^{0,1}_{n,l}\mathbf{h}^{0,1}_{m,l}\Sigma^{1}_{m,l}),$
$\mathbf{d}^{25}_{l,m}\triangleq 2({\sum}_{n\neq m}^{M_{r}}(\mathbf{D}^{11}_{l,m,n})^{H}\hat{\mathbf{h}}^{1}_{n,l}),$
$\mathbf{D}^{14}_{j}\triangleq{\sum}_{k=1}^{K_{d}}[(\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}\hat{\mathbf{g}}_{t,j})^{T}\otimes\sigma^{2}_{t,j}(\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H})]$
$+(\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}^{r}(\hat{\mathbf{W}}^{r})^{H}\hat{\mathbf{g}}_{t,j})^{T}\otimes\sigma^{2}_{t,j}(\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}),$
$\bar{\mathbf{D}}^{14}_{j,i,z}\triangleq\mathbf{D}^{14}_{j}[(i-1)N_{r}+1:iN_{r},(z-1)N_{r}+1:zN_{r}],$
$\mathbf{d}^{26}_{j,m}\triangleq 2({\sum}_{n\neq m}^{M_{r}}({\mathbf{g}}_{r,n,j}^{H}\bar{\mathbf{D}}^{14}_{j,n,m})),\mathbf{d}^{27}_{m}\triangleq 2({\sum}_{n\neq m}^{M_{r}}({\mathbf{g}}_{r,n,0}^{H}\bar{\mathbf{D}}^{10}_{n,m})),$
${\mathbf{D}}^{2}_{l}\triangleq\mathbf{D}^{11}_{l,m,m}[(n-1)L^{t,1}_{m,l}+1:nL^{t,1}_{m,l},(i-1)L^{t,1}_{m,l}+1:iL^{t,1}_{m,l}],$
$\mathbf{d}^{28}_{l,m,n}\triangleq 2({\sum}_{j\neq n}^{N_{r}}({\mathbf{D}}^{15}_{l,m,j,n})^{H}\mathbf{h}^{1}_{m,l,j}),$
$\mathbf{d}^{4}_{l}\!\triangleq\!\mathbf{d}^{28}_{l,m,n}\!+\!\mathbf{d}^{25}_{l,m}[(n\!-\!1)L^{t,1}_{m,l}\!+\!1\!:\!nL^{t,1}_{m,l}],$
$d^{5}_{j}\!\triangleq\!2({\sum}_{i\neq n}^{N_{r}}(\bar{\mathbf{D}}^{14}_{j,m,m}[i,n])^{\ast}{\mathbf{g}}_{r,m,j}[i])\!+\!\mathbf{d}^{26}_{j,m}[n],$
$d^{6}\!\triangleq\!2({\sum}_{i\neq n}^{N_{r}}(\bar{\mathbf{D}}^{10}_{m,m}[i,n])^{\ast}{\mathbf{g}}_{r,m,0}[i])\!+\!\mathbf{d}^{27}_{m}[n].$
Therefore, the variable $\mathbf{t}^{1}$ optimization reduces to solving the following problem
$\textrm{(P11)}:$ $\mathop{\textrm{min}}\limits_{\mathbf{t}^{1}}{\sum}_{l=1}^{K_{u}}\big((\mathbf{h}^{1}_{l})^{H}\mathbf{D}^{1}_{l}\mathbf{h}^{1}_{l}+\textrm{Re}\{(\mathbf{d}^{1}_{l})^{H}\mathbf{h}^{1}_{l}\}$ (81a)
$-\textrm{Re}\{(\mathbf{{d}}^{2}_{l})^{H}\mathbf{h}^{1}_{l}\}\big)+{\sum}_{j=0}^{K_{t}}\textrm{Re}\{(d^{3}_{j})^{\ast}\mathbf{g}_{r,j}\}+c_{5,1}$
s.t. ${\sum}_{l=1}^{K_{u}}\big((\mathbf{h}^{1}_{l})^{H}\mathbf{D}^{2}_{l}\mathbf{h}^{1}_{l}+\textrm{Re}\{(\mathbf{d}^{4}_{l})^{H}\mathbf{h}^{1}_{l}\}\big)$ (81b)
$\!+\!{\sum}_{j=1}^{K_{t}}\!\big(\!\textrm{Re}\{\!(d^{5}_{j})^{\ast}\!{g}_{r,j}\}\!\big)-\textrm{Re}\{(d^{6})^{\ast}\mathbf{g}_{r,0}\}+c_{5,2}\leq 0,$
$\|\mathbf{t}^{1}-\mathbf{t}^{1}_{m,i}\|_{2}\geq D_{t},\forall n,i\in\mathcal{N}_{t},n\neq i,$ (81c)
$\mathbf{t}^{1}\in\mathcal{C}.$ (81d)
In the following, we will obtain the upper-bound of the objective function (81a) via convexify it. By leveraging the arguments presented in (62), a tight upper-bound of $(\mathbf{h}^{1}_{l})^{H}\mathbf{D}^{1}_{l}\mathbf{h}^{1}_{l}+\textrm{Re}\{(\mathbf{d}^{1}_{l})^{H}\mathbf{h}^{1}_{l}\}$ can be shown as follows
$(\mathbf{h}^{1}_{l})^{H}\mathbf{D}^{1}_{l}\mathbf{h}^{1}_{l}+\textrm{Re}\{(\mathbf{d}^{1}_{l})^{H}\mathbf{h}^{1}_{l}\}$ (82)
$\leq\lambda_{max}(\mathbf{D}^{1}_{l})(\|\mathbf{h}^{1}_{l}\|^{2}_{2}+\|\mathbf{h}^{1}_{l,0}\|^{2}_{2})-2\textrm{Re}\{\lambda_{max}(\mathbf{D}^{1}_{l})(\mathbf{h}^{1}_{l,0})^{H}\mathbf{h}^{1}_{l}\}$
$\!+\!2\textrm{Re }\{(\mathbf{h}^{1}_{l,0})^{H}\mathbf{D}^{1}_{l}(\mathbf{h}^{1}_{l}\!-\!\mathbf{h}^{1}_{l,0})\}+(\mathbf{h}^{1}_{l,0})^{H}\mathbf{D}^{1}_{l}\mathbf{h}^{1}_{l,0}\!+\!\textrm{Re}\{(\mathbf{d}^{1}_{l})^{H}\mathbf{h}^{1}_{l}\}$
$=\!\textrm{Re}\{(\mathbf{d}^{7}_{l})^{H}\mathbf{h}^{1}_{l}\}\!+\!c_{5,3,l}\!+\!\textrm{Re}\{(\mathbf{d}^{1}_{l})^{H}\!\mathbf{h}^{1}_{l}\}\!=\!\textrm{Re}\{(\mathbf{d}^{8}_{l})^{H}\!\mathbf{h}^{1}_{l}\}\!\!+\!\!c_{5,3,l},$
where $\mathbf{t}^{1}_{0}$ is the feasible solution obtained in the last iteration, $\mathbf{h}^{1}_{l,0}\!\triangleq\!\mathbf{h}^{1}_{l}(\mathbf{t}^{1}_{0})$, $\|\mathbf{h}^{1}_{l}\|^{2}_{2}\!=\!\|\mathbf{h}^{1}_{l,0}\|^{2}_{2}\!=\!L^{t,1}_{m,l}$, $c_{5,3,l}\!\triangleq\!2\lambda_{max}(\mathbf{D}^{1}_{l})L^{t,1}_{m,l}-((\mathbf{h}^{1}_{l,0})^{H}\mathbf{D}^{1}_{l}\mathbf{h}^{1}_{l,0})^{\ast}$, $\mathbf{d}^{7}_{l}\!\triangleq\!2((\mathbf{D}^{1}_{l})^{H}\mathbf{h}^{1}_{l,0}\!-\!\lambda_{max}(\mathbf{D}^{1}_{l})\mathbf{h}^{1}_{l,0})$ and $\mathbf{d}^{8}_{l}\!\triangleq\!\mathbf{d}^{7}_{l}\!+\!\mathbf{d}^{1}_{l}$.
Furthermore, by using the inequality (65) in Lemma 1, we can obtain a convex upper-bound of $\textrm{Re}\{(\mathbf{d}^{8}_{l})^{H}\mathbf{h}^{1}_{l}\}$ as follows
$\textrm{Re}\{(\mathbf{d}^{8}_{l})^{H}\mathbf{h}^{1}_{l}\}\triangleq h^{3}_{l}(\mathbf{t}^{1})$ (83)
$={\sum}_{i=1}^{L^{t,1}_{m,l}}|\mathbf{d}^{8}_{l}[i]|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{1})^{T}\mathbf{a}^{r,1}_{m,l,i}-\angle\mathbf{d}^{8}_{l}[i]\big)$
$\leq h^{3}_{l}(\mathbf{t}^{1}_{0})+\nabla^{T}h^{3}_{l}(\mathbf{t}^{1}_{0})(\mathbf{t}^{1}-\mathbf{t}^{1}_{0})+{0.5\tau_{1,l}}(\mathbf{t}^{1}-\mathbf{t}^{1}_{0})^{T}(\mathbf{t}^{1}-\mathbf{t}^{1}_{0})$
$={0.5\tau_{1,l}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{9}_{l})^{T}\mathbf{t}^{1}+c_{5,4,l}$
where $c_{5,4,l}$ is a constant term, $\tau_{1,l}=\frac{8\pi^{2}}{\lambda^{2}}{\sum}_{i=1}^{L^{t,1}_{m,l}}|\mathbf{d}^{8}_{l}[i]|$, $\mathbf{d}^{9}_{l}\triangleq\nabla h^{3}_{l}(\mathbf{t}^{1}_{0})-\tau_{1,l}\mathbf{t}^{1}_{0}$, $\nabla h^{3}_{l}(\mathbf{t}^{1}_{0})$ is the gradient vector.
Furthermore, by leveraging the inequality (64), the lower-bound of $\textrm{Re}\{(\mathbf{d}^{2}_{l})^{H}\mathbf{h}^{1}_{l}\}$ is given as
$\textrm{Re}\{(\mathbf{d}^{2}_{l})^{H}\mathbf{h}^{1}_{l}\}\triangleq h^{4}_{l}(\mathbf{t}^{1})$ (84)
$={\sum}_{i=1}^{L^{t,1}_{m,l}}|\mathbf{d}^{2}_{l}[i]|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{1})^{T}\mathbf{a}^{r,1}_{m,l,i}-\angle\mathbf{d}^{2}_{l}[i]\big)$
$\geq h^{4}_{l}(\mathbf{t}^{1}_{0})+\nabla^{T}h^{4}_{l}(\mathbf{t}^{1}_{0})(\mathbf{t}^{1}-\mathbf{t}^{1}_{0})-{0.5\bar{\tau}_{1,l}}(\mathbf{t}^{1}-\mathbf{t}^{1}_{0})^{T}(\mathbf{t}^{1}-\mathbf{t}^{1}_{0})$
$=-{0.5\bar{\tau}_{1,l}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{10}_{l})^{T}\mathbf{t}^{1}+c_{5,5,l},$
with $c_{5,5,l}$ being constant term, $\bar{\tau}_{1,l}=\frac{8\pi^{2}}{\lambda^{2}}{\sum}_{i=1}^{L^{t,1}_{m,l}}|\mathbf{d}^{2}_{l}[i]|$, $\mathbf{d}^{10}_{l}\triangleq\bar{\tau}_{1,l}\mathbf{t}^{1}_{0}+h^{4}_{l}(\mathbf{t}^{1}_{0})$, $\nabla h^{4}_{l}(\mathbf{t}^{1}_{0})$ is the gradient vector.
Next, the term $\textrm{Re}\{(d^{3}_{j})^{\ast}{g}_{r,j}\}$ can be convexified by leveraging (65) as follows
$\textrm{Re}\{(d^{3}_{j})^{\ast}{g}_{r,j}\}\triangleq g^{4}_{j}(\mathbf{t}^{1})$ (85)
$=|d^{3}_{j}||\mathbf{g}_{r,j}|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{1})^{T}\mathbf{a}^{5}_{m,j}-\angle d^{3}_{j}\big)$
$\leq g^{4}_{j}(\mathbf{t}^{1}_{0})\!+\!\nabla^{T}g^{4}_{j}(\mathbf{t}^{1}_{0})(\mathbf{t}^{1}\!-\!\mathbf{t}^{1}_{0})\!+\!{0.5{\tau}_{2,j}}(\mathbf{t}^{1}\!-\!\mathbf{t}^{1}_{0})^{T}(\mathbf{t}^{1}\!-\!\mathbf{t}^{1}_{0})$
$={0.5\tau_{2,j}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{11}_{j})^{T}\mathbf{t}^{1}+c_{5,6,j},$
with $c_{5,6,j}$ being constant, $\tau_{2,j}\triangleq\frac{8\pi^{2}}{\lambda^{2}}|d^{3}_{j}||\mathbf{g}_{r,j}|$, $\mathbf{d}^{11}_{j}\triangleq\nabla g^{4}_{j}(\mathbf{t}^{1}_{0})-\tau_{2,j}\mathbf{t}^{1}_{0}$, $\nabla g^{4}_{j}(\mathbf{t}^{1}_{0})$ is the gradient vector.
Based on the above transformation, the upper-bound of the objective function (81a) is given as
${\sum}_{l=1}^{K_{u}}\big((\mathbf{h}^{1}_{l})^{H}\mathbf{D}^{1}_{l}\mathbf{h}^{1}_{l}+\textrm{Re}\{(\mathbf{d}^{1}_{l})^{H}\mathbf{h}^{1}_{l}\}$ (86)
$-\textrm{Re}\{(\mathbf{d}^{2}_{l})^{H}\mathbf{h}^{1}_{l}\}\big)+{\sum}_{j=0}^{K_{t}}\textrm{Re}\{(d^{3}_{j})^{\ast}\mathbf{g}_{r,j}\}+c_{5,1}$
$\leq{\sum}_{l=1}^{K_{u}}\big({0.5\tau_{1,l}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{9}_{l})^{T}\mathbf{t}^{1}+c_{5,4,l}$
$-(-{0.5\bar{\tau}_{1,l}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{10}_{l})^{T}\mathbf{t}^{1}+\bar{c}_{5,5,l})\big)$
$+{\sum}_{j=0}^{K_{t}}(0.5{\tau_{2,j}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{11}_{j})^{T}\mathbf{t}^{1}+c_{5,6,j})+c_{5,1}$
$={\tau_{3}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{12})^{T}\mathbf{t}^{1}+c_{5,7},$
where $c_{5,7}$ is a constant term, $\tau_{3}\triangleq{\sum}_{l=1}^{K_{u}}\big({\tau_{1,l}}/{2}+{\bar{\tau}_{1,l}}/{2}\big)+{\sum}_{j=0}^{K_{t}}({\tau_{2,j}}/{2})$ and $\mathbf{d}^{12}\triangleq{\sum}_{l=1}^{K_{u}}(\mathbf{d}^{9}_{l}-\mathbf{d}^{10}_{l})+{\sum}_{j=0}^{K_{t}}\mathbf{d}^{11}_{j}$.
Next, we will construct the upper-bound of the sensing SINR constraint (81b). Firstly, by leveraging the arguments presented in (62) again, a tight upper-bound of $(\mathbf{h}^{1}_{l})^{H}\mathbf{D}^{2}_{l}\mathbf{h}^{1}_{l}+\textrm{Re}\{(\mathbf{d}^{4}_{l})^{H}\mathbf{h}^{1}_{l}\}$, can be formulated as
$(\mathbf{h}^{1}_{l})^{H}\mathbf{D}^{2}_{l}\mathbf{h}^{1}_{l}+\textrm{Re}\{(\mathbf{d}^{4}_{l})^{H}\mathbf{h}^{1}_{l}\}$ (87)
$\leq\lambda_{max}(\mathbf{D}^{2}_{l})(\|\mathbf{h}^{1}_{l}\|^{2}_{2}+\|\mathbf{h}^{1}_{l,0}\|^{2}_{2})-2\textrm{Re}\{\lambda_{max}(\mathbf{D}^{2}_{l})(\mathbf{h}^{1}_{l0})^{H}\mathbf{h}^{1}_{l}\}$
$+2\textrm{Re}\{(\mathbf{h}^{1}_{l,0})^{H}\mathbf{D}^{2}_{l}(\mathbf{h}^{1}_{l}\!-\!\mathbf{h}^{1}_{l,0})\}\!\!+\!\!(\mathbf{h}^{1}_{l,0})^{H}\mathbf{D}^{2}_{l}\mathbf{h}^{1}_{l,0}\!\!+\!\!\textrm{Re}\{(\mathbf{d}^{4}_{l})^{H}\mathbf{h}^{1}_{l}\}$
$=\!\textrm{Re}\{(\mathbf{d}^{13}_{l})^{H}\!\mathbf{h}^{1}_{l}\}\!\!+\!\!c_{5,8,l}\!+\!\textrm{Re}\{(\mathbf{d}^{4}_{l})^{H}\!\mathbf{h}^{1}_{l}\}\!\!=\!\textrm{Re}\{(\mathbf{d}^{14}_{l})^{H}\!\mathbf{h}^{1}_{l}\}\!\!+\!\!c_{5,8,l},$
where $c_{5,8,l}$ is constant, $\mathbf{d}^{13}_{l}\!\!\triangleq\!\!2((\mathbf{D}^{2}_{l})^{H}\mathbf{h}^{1}_{l,0}\!-\!\lambda_{max}(\mathbf{D}^{2}_{l})\mathbf{h}^{1}_{l,0})$ and $\mathbf{d}^{14}_{l}\!\triangleq\!\mathbf{d}^{13}_{l}\!+\!\mathbf{d}^{2}_{l}$.
Furthermore, by using the inequality (65) in Lemma 1, a convex upper-bound of $\textrm{Re}\{(\mathbf{d}^{14}_{l})^{H}\mathbf{h}^{1}_{l}\}$ is given as follows
$\textrm{Re}\{(\mathbf{d}^{14}_{l})^{H}\mathbf{h}^{1}_{l}\}\triangleq h^{5}_{l}(\mathbf{t}^{1})$ (88)
$={\sum}_{i=1}^{L^{t,1}_{m,l}}|\mathbf{d}^{14}_{l}[i]|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{1})^{T}\mathbf{a}^{r,1}_{m,l,i}-\angle\mathbf{d}^{14}_{l}[i]\big)$
$\leq h^{5}_{l}(\mathbf{t}^{1}_{0})+\nabla^{T}h^{5}_{l}(\mathbf{t}^{1}_{0})(\mathbf{t}^{1}-\mathbf{t}^{1}_{0})+{0.5\tau_{4,l}}(\mathbf{t}^{1}-\mathbf{t}^{1}_{0})^{T}(\mathbf{t}^{1}-\mathbf{t}^{1}_{0})$
$={0.5\tau_{4,l}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{15}_{l})^{T}\mathbf{t}^{1}+c_{5,9,l},$
where $c_{5,9,l}$ is a constant term, $\tau_{4,l}=\frac{8\pi^{2}}{\lambda^{2}}{\sum}_{i=1}^{L^{t,1}_{m,l}}|\mathbf{d}^{14}_{l}[i]|$, $\mathbf{d}^{15}_{l}\triangleq\nabla h^{5}_{l}(\mathbf{t}^{1}_{0})-\tau_{4,l}\mathbf{t}^{1}_{0}$, $\nabla h^{5}_{l}(\mathbf{t}^{1}_{0})$ is the gradient vector.
Similarly, the term $\textrm{Re}\{(d^{5}_{j})^{\ast}{g}_{r,j}\}$ can be convexified by leveraging (65) as follows
$\textrm{Re}\{(d^{5}_{j})^{\ast}{g}_{r,j}\}\triangleq g^{5}_{j}(\mathbf{t}^{1})$ (89)
$=|d^{5}_{j}||\mathbf{g}_{r,j}|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{1})^{T}\mathbf{a}^{5}_{m,j}-\angle d^{5}_{j}\big)$
$\leq{0.5\tau_{5,j}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{16}_{j})^{T}\mathbf{t}^{1}+c_{5,10,j},$
with $c_{5,10,j}$ being constants, $\tau_{5,j}\triangleq\frac{8\pi^{2}}{\lambda^{2}}|d^{5}_{j}||{g}_{r,j}|$, $\mathbf{d}^{16}_{j}\triangleq\nabla g^{5}_{j}(\mathbf{t}^{1}_{0})-\tau_{5,j}\mathbf{t}^{1}_{0}$, $\nabla g^{5}_{j}(\mathbf{t}^{1}_{0})$ is the gradient vector.
And then, a convex lower bound for $\text{Re}\{(d^{6})^{\ast}{g}_{r,0}\}$ by applying (64) in Lemma 1 is given as follows
$\text{Re}\{(d^{6})^{\ast}{g}_{r,0}(\mathbf{t}^{1})\}\triangleq g^{6}(\mathbf{t}^{1})$ (90)
$=|d^{6}||\mathbf{g}_{r,0}|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{1})^{T}\mathbf{a}^{5}_{m,0}-\angle d^{6}\big)$
$\geq-{0.5\tau_{6}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{17})^{T}\mathbf{t}^{1}_{m,n}+c_{5,11},$
where $c_{5,11}$ is constant, $\tau_{6}\triangleq\frac{8\pi^{2}}{\lambda^{2}}|d^{6}||\mathbf{g}_{r,0}|$, $\mathbf{d}^{17}\triangleq\nabla g^{6}(\mathbf{t}^{1})+\tau_{6}\mathbf{t}^{1}_{0}$, $\nabla g^{6}(\mathbf{t}^{0})$ is the the gradient vector.
Therefore, a convex upper-bound of the sensing SINR constraint (81b) can be expressed as
${\sum}_{l=1}^{K_{u}}\big((\mathbf{h}^{1}_{l})^{H}\mathbf{D}^{2}_{l}\mathbf{h}^{1}_{l}+\textrm{Re}\{(\mathbf{d}^{4}_{l})^{H}\mathbf{h}^{1}_{l}\}\big)$ (91)
$+{\sum}_{j=1}^{K_{t}}\big(\textrm{Re}\{(d^{5}_{j})^{\ast}{g}_{r,j}\}\big)-\textrm{Re}\{(d^{6})^{\ast}\mathbf{g}_{r,0}\}+c_{5,2}$
$\leq{\sum}_{l=1}^{K_{u}}\big({0.5\tau_{4,l}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{15}_{l})^{T}\mathbf{t}^{1}+c_{5,9,l}+c_{5,8,l}\big)$
$+{\sum}_{j=1}^{K_{t}}({0.5\tau_{5,j}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{16}_{j})^{T}\mathbf{t}^{1}+c_{5,10,j})$
$-(-{0.5\tau_{6}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{17})^{T}\mathbf{t}^{1}_{m,n}+c_{5,11})+c_{5,2},$
$=\tau_{7}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{18})^{T}\mathbf{t}^{1}+c_{5,12},$
where $c_{5,12}$ is constant, $\tau_{7}\triangleq{\sum}_{l=1}^{K_{u}}({\tau_{4,l}}/{2})+{\sum}_{j=1}^{K_{t}}({\tau_{5,j}}/{2})-(-{\tau_{6}}/{2})$, $\mathbf{d}^{18}\triangleq{\sum}_{l=1}^{K_{u}}\mathbf{d}^{15}_{l}+{\sum}_{j=1}^{K_{t}}\mathbf{d}^{16}_{j}-(\mathbf{d}^{17})$.
Inspired by the MM framework, the constraint (81c) can be linearized as follows
$\|\mathbf{t}^{1}-\mathbf{t}^{1}_{m,i}\|_{2}\geq\frac{(\mathbf{t}^{1}_{0}-\mathbf{t}^{1}_{m,i})^{T}}{\|\mathbf{t}^{1}_{0}-\mathbf{t}^{1}_{m,i}\|_{2}}(\mathbf{t}^{1}-\mathbf{t}^{1}_{m,i}).$ (92)
Therefore, the optimization problem (P11) is rewritten as
$\textrm{(P12)}:$ $\mathop{\textrm{min}}\limits_{\mathbf{t}^{1}}{\tau_{3}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{12})^{T}\mathbf{t}^{1}+c_{5,7}$ (93a)
s.t. $\tau_{7}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{18})^{T}\mathbf{t}^{1}+c_{5,12}\leq 0,$ (93b)
$\frac{(\mathbf{t}^{1}_{0}-\mathbf{t}^{1}_{m,i})^{T}}{\|\mathbf{t}^{1}_{0}-\mathbf{t}^{1}_{m,i}\|_{2}}(\mathbf{t}^{1}-\mathbf{t}^{1}_{m,i})\geq\!\!D_{t},\!\forall\!n,\!i\!\in\!\mathcal{N}_{r},n\!\neq\!i,$ (93c)
$\mathbf{t}^{1}\in\mathcal{C}.$ (93d)
Obviously, the problem (P12) is convex and its optimal solution can be obtained by CVX.
III-I Optimizing DL User MA’s Position
In this subsection, we will discuss the optimization of the MA position $\mathbf{r}_{0,k}$ for the $k$-th DL user. With given other variables, the problem w.r.t. $\mathbf{r}_{0,k}$ is formulated as
$\textrm{(P13)}:$ $\mathop{\textrm{min}}\limits_{\mathbf{r}^{0}_{k}\in\mathcal{C}}{\sum}_{m=1}^{M_{t}}{\sum}_{n=1}^{M_{t}}(\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{Q}^{1}_{k,m,n}\mathbf{h}^{0,0}_{n,k}$ (94a)
$+{\sum}_{m=1}^{M_{t}}((\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{Q}^{2}_{k,m}\mathbf{h}^{0,0}_{m,k}\!-\!\!\textrm{ Re }\{(\mathbf{q}^{1}_{k,m})^{H}\mathbf{h}^{0,0}_{m,k}\})$
$+{\sum}_{l=1}^{K_{u}}(\mathbf{h}^{3}_{k,l})^{H}\mathbf{Q}^{3}_{k,l}\mathbf{h}^{3}_{k,l}+c_{6,1,k},$
with $c_{6,1,k}$ being constant and other new coefficients are given as follows
$\mathbf{q}^{1}_{k,m}\triangleq 2\omega_{d,k}\beta_{d,k}^{\ast}\mathbf{\Sigma}^{0}_{m,k}\mathbf{H}^{0}_{m,k}\mathbf{w}_{m,k},$ (95)
$\mathbf{Q}^{3}_{k,l}\triangleq\omega_{d,k}|\beta_{d,k}|^{2}q_{l}\mathbf{\Sigma}^{2}_{k,l}\mathbf{h}^{2}_{k,l}(\mathbf{h}^{2}_{k,l})^{H}(\mathbf{\Sigma}^{2}_{k,l})^{H},$
$\mathbf{Q}^{2}_{k,m}\triangleq\omega_{d,k}|\beta_{d,k}|^{2}\mathbf{\Sigma}^{0}_{m,k}\mathbf{H}^{0}_{m,k}\mathbf{W}^{r}_{m}\mathbf{W}^{r}_{m}(\mathbf{H}^{0}_{m,k})^{H}(\mathbf{\Sigma}^{0}_{m,k})^{H},$
$\mathbf{Q}^{1}_{k,m,n}\!\triangleq\!\!\!{\sum}_{i=1}^{K_{d}}\!\omega_{d,k}|\beta_{d,k}|^{2}\mathbf{\Sigma}^{0}_{m,k}\mathbf{H}^{0}_{m,k}\mathbf{w}_{m,i}\mathbf{w}_{n,i}(\mathbf{H}^{0}_{n,k})^{\!H}\!\!(\mathbf{\Sigma}^{0}_{n,k})^{H}.$
Obviously, obtaining the optimal solution to the problem (P13) is challenging due to the non-convex nature of the objective function (94a). Therefore, we still employ the MM methodology to convexify it. Due to space constraints, we omit the transformation process.
The update of $\mathbf{r}_{0,k}$ is achieved by solving a convex surrogate of the original problem (P13), as follows
$\textrm{(P14)}:$ $\mathop{\textrm{min}}\limits_{\mathbf{r}_{0,k}\in\mathcal{C}}{\psi_{7,k}}\|\mathbf{r}_{0,k}\|^{2}_{2}+(\mathbf{q}^{2}_{k})^{T}\mathbf{r}_{0,k}+c_{6,2,k}$ (96a)
The problem (P14) is convex and can be solved by CVX.
Since the units of $\mathbf{r}_{0,k}$, i.e., $x^{d}_{k}$ and $y^{d}_{k}$, are not coupled, the subproblem w.r.t. $x^{d}_{k}$ can be written as
$\textrm{(P15)}:$ $\mathop{\textrm{min}}\limits_{x^{d}_{k}}{\psi_{7,k}}(x^{d}_{k})^{2}+\mathbf{q}^{2}_{k}[1]x^{d}_{k}$ (97a)
s.t. $-{A}/{2}\leq x^{d}_{k}\leq{A}/{2},$ (97b)
Since $\psi_{7,k}>0$, we can obtain the optimal solution $(x^{d}_{k})^{\star}$ by directly determining the position of the axis of symmetry of the function (97a), which are given as
$(x^{d}_{k})^{\star}\!=\!\left\{\begin{aligned} -{A}/{2},\ &\textrm{if}\ -{\mathbf{q}^{2}_{k}[1]}/({2\psi_{7,k}})<-{A}/{2},\\ -{\mathbf{q}^{2}_{k}[1]}/({2\psi_{7,k}}),\ &\textrm{if}\ -{A}/{2}\leq-{\mathbf{q}^{2}_{k}[1]}/({2\psi_{7,k}})\leq{A}/{2},\\ {A}/{2},\ &\textrm{if}\ -{A}/{2}<-{\mathbf{q}^{2}_{k}[1]}/({2\psi_{7,k}}).\end{aligned}\right$. (98)
Similarly, the optimal solution $(y^{d}_{k})^{\star}$ can be directly formulated as
$(y^{d}_{k})^{\star}\!=\!\left\{\begin{aligned} -{A}/{2},\ &\textrm{if}\ -{\mathbf{q}^{2}_{k}[2]}/({2\psi_{7,k}})<-{A}/{2},\\ -{\mathbf{q}^{2}_{k}[1]}/({2\psi_{7,k}}),\ &\textrm{if}\ -{A}/{2}\leq-{\mathbf{q}^{2}_{k}[2]}/({2\psi_{7,k}})\leq{A}/{2},\\ {A}/{2},\ &\textrm{if}\ -{A}/{2}<-{\mathbf{q}^{2}_{k}[2]}/({2\psi_{7,k}}).\end{aligned}\right$. (99)
III-J Optimizing UL User MA’s Position
In this subsection, we will investigate the update of the MA’s position $\mathbf{r}_{1,l}$} for the $l$-th UL user. Firstly, the functions $\tilde{\mathrm{R}}_{d,k}$ and $\tilde{\mathrm{R}}_{u,l}$ can be respectively rewritten as
${\sum}_{k=1}^{K_{d}}\tilde{\mathrm{R}}_{d,k}\!\triangleq\!{\sum}_{k=1}^{K_{d}}\!-\!(\mathbf{h}^{2}_{k,l}(\mathbf{r}_{1,l}))^{H}\mathbf{Q}^{4}_{k,l}\mathbf{h}^{2}_{k,l}(\mathbf{r}_{1,l})\!+\!c_{7,1,l},$ (100)
${\sum}_{l=1}^{K_{u}}\tilde{\mathrm{R}}_{u,l}\!\triangleq\!\!{\sum}_{m=1}^{M_{r}}{\sum}_{n=1}^{M_{r}}(\mathbf{h}^{0,1}_{m,l}(\mathbf{r}_{1,l}))^{H}\mathbf{Q}^{5}_{l,m,n}\mathbf{h}^{0,1}_{n,l}(\mathbf{r}_{1,l})$ (101)
$-{\sum}_{m=1}^{M_{r}}\textrm{Re}\{\mathbf{q}^{3}_{l,m}\mathbf{h}^{0,1}_{m,l}(\mathbf{r}_{1,l})\}+c_{7,2,l},$
with $c_{7,1,l}$ and $c_{7,2,l}$ being constant terms, and we define
$\mathbf{q}^{3}_{l,m}\triangleq 2\omega_{d,k}\beta_{d,k}\mathbf{\Sigma}^{1}_{m,l}\mathbf{H}^{1}_{m,i}\mathbf{u}_{m,l},$ (102)
$\mathbf{Q}^{4}_{k,l}\triangleq\omega_{d,k}|\beta_{d,k}|^{2}q_{l}(\mathbf{\Sigma}^{2}_{k,l})^{H}\mathbf{h}^{3}_{k,l}(\mathbf{h}^{3}_{k,l})^{H}\mathbf{\Sigma}^{2}_{k,l},$
${\mathbf{Q}}^{5}_{i,m,n}\!\triangleq\!\!{\sum}_{l=1}^{K_{u}}\omega_{d,k}|\beta_{d,k}|^{2}\mathbf{\Sigma}^{1}_{m,i}\mathbf{H}^{1}_{m,i}\mathbf{u}_{m,l}\mathbf{u}_{n,l}^{H}(\mathbf{H}^{1}_{n,i})^{H}\!(\mathbf{\Sigma}^{1}_{n,i})^{H}.$
Furthermore, the sensing SINR constraint (30b) can be equivalent rewritten as
$\textrm{SINR}_{t}\geq\Gamma_{r}\Longleftrightarrow$ (103)
${\sum}_{m=1}^{M_{r}}{\sum}_{n=1}^{M_{r}}(\mathbf{h}^{0,1}_{m,l}(\mathbf{r}_{1,l}))^{H}\mathbf{Q}^{6}_{l,m,n}\mathbf{h}^{0,1}_{n,l}(\mathbf{r}_{1,l})+c_{7,3,l},$
where $c_{7,3,l}$ is constant and $\mathbf{Q}^{6}_{l,m,n}\triangleq\mathbf{\Sigma}^{1}_{m,l}\mathbf{H}^{1}_{m,l}\mathbf{u}_{m,0}\mathbf{u}_{n,0}^{H}(\mathbf{H}^{1}_{n,l})^{H}(\mathbf{\Sigma}^{1}_{n,l})^{H}$.
Therefore, the problem w.r.t. $\mathbf{r}_{1,l}$ can be rewritten as
$\textrm{(P16)}:$ $\mathop{\textrm{min}}\limits_{\mathbf{r}_{1,l}}\!{\sum}_{k=1}^{K_{d}}\!(\mathbf{h}^{2}_{k,l})^{H}\mathbf{Q}^{4}_{k,l}\mathbf{h}^{2}_{k,l}\!\!-\!\!\!{\sum}_{m=1}^{M_{r}}\textrm{Re}\{\mathbf{q}^{3}_{l,m}\mathbf{h}^{0,1}_{m,l}\}$
$+{\sum}_{m=1}^{M_{r}}{\sum}_{n=1}^{M_{r}}(\mathbf{h}^{0,1}_{m,l})^{H}\mathbf{Q}^{5}_{l,m,n}\mathbf{h}^{0,1}_{n,l}+\hat{c}_{7,1,l}$ (104a)
s.t. $\!{\sum}_{m=1}^{M_{r}}{\sum}_{n=1}^{M_{r}}(\mathbf{h}^{0,1}_{m,l})^{H}\mathbf{Q}^{6}_{l,m,n}\mathbf{h}^{0,1}_{n,l}\!+\!c_{7,3,l}\!\leq\!0,$ (104b)
$\mathbf{r}_{1,l}\in\mathcal{C},$ (104c)
with $\hat{c}_{7,1,l}$ being a constant term.
Unfortunately, this problem is still difficult to solve due to the non-convex objective (104a) and constraint (104b) w.r.t $\mathbf{r}_{1,l}$. To address these issues, we will construct the convex surrogate functions of (104a) and (104b) by applying the MM method, which the derivation details are omitted for the space constraints, and then recast the problem (P11) into an explicit form of $\mathbf{r}_{1,l}$, which is given as
$\textrm{(P17)}:$ $\mathop{\textrm{min}}\limits_{\mathbf{r}_{1,l}}\ \psi_{8,l}\|\mathbf{r}_{1,l}\|^{2}_{2}+(\mathbf{q}^{4}_{l})^{T}\mathbf{r}_{1,l}+c_{7,4,l}$ (105a)
s.t. $\psi_{10,l}\|\mathbf{r}_{1,l}\|^{2}_{2}+(\mathbf{q}^{5}_{l})^{T}\mathbf{r}_{1,l}+c_{7,5,l}\leq 0,$ (105b)
$\mathbf{r}_{1,l}\in\mathcal{C}.$ (105c)
Now, the problem (P17) is convex, which can be solved optimally by CVX.
Moreover, the overall algorithm to solve (P1) is specified in Algorithm 1. $\eta=\{\{\mathbf{w}_{m,k},\mathbf{W}^{r}_{m}\},\{\mathbf{u}_{p,i}\},\{q_{l}\},\{\mathbf{t}^{0}_{m,n},\mathbf{t}^{1}_{p,j},\mathbf{r}_{0,k},\mathbf{r}_{1,l}\}\}$ denote the collection of all variables. And let $g(\eta)$ represent the sum-rate for both downlink and uplink users.
Algorithm 1 Overall Algorithm to Solve $(\mathrm{P1})$
1: initialize $i=0$;
2: randomly generate feasible $\{\mathbf{w}^{0}_{m,k},(\mathbf{W}^{r}_{m})^{0}\}$, $\{\mathbf{u}_{p,j}^{0}\}$, $\{q_{l}\}$, $\{(\mathbf{t}^{0}_{m,n})^{0},(\mathbf{t}^{1}_{m,n})^{0}\}$ and $\{\mathbf{r}_{0,k}^{0},\mathbf{r}_{1,l}^{0}\}$;
3:repeat
4:  update $\{\omega_{d,k}\}$, $\{\beta_{d,k}\}$, $\{\omega_{u,l}\}$ and $\{\beta_{u,l}\}$, by (34), (33), (36) and (35), respectively;
5:  update $\{\mathbf{w}^{i+1}_{m,k},(\mathbf{W}^{r}_{m})^{i+1}\}$ by solving (P3);
6:  update $\{\mathbf{u}_{p,l}^{i+1}\},\forall l\in\mathcal{K}_{u}$ by equation (51);
7:  update $\{\mathbf{u}_{p,0}^{i+1}\}$ by solving (P8);
8:  update $\{(\mathbf{t}^{0}_{m,n})^{i+1}\}$ by solving (P10);
9:  update $\{(\mathbf{t}^{1}_{m,n})^{i+1}\}$ by solving (P12);
10:  update $\{\mathbf{r}_{0,k}^{i+1}\}$ by equations (98) and (99);
11:  update $\{\mathbf{r}_{1,l}^{i+1}\}$ by solving (P17);
12:  $i++$;
13:until $\bigg|\frac{g(\eta^{i+1})-g(\eta^{i})}{g(\eta^{i})}\bigg|\leq\epsilon,$
III-K Complexity
In the following, we will discuss the complexity of our proposed algorithms. According to the complexity analysis in [45], the complexity of solving SOCP problems (P3), (P10), (P12) and (P17) are $\mathcal{O}(C_{1}M_{t}^{3.5}N_{t}^{6})$, $\mathcal{O}(C_{2}M_{t}N_{t}^{1.5})$, $\mathcal{O}(C_{3}M_{r}N_{r}^{1.5})$ and $\mathcal{O}(76C_{4})$, respectively, where $C_{1}$, $C_{2}$, $C_{3}$ and $C_{4}$ are respectively the number of iterations of solving problem (P3), (P10), (P12) and (P17) by the MM method. The complexity of solving (P4) is $\mathcal{O}(K_{u}^{3.5})$. Therefore, the total computational complexity of Algorithm 1 is approximately given as $\mathcal{O}(C_{5}(C_{1}M_{t}^{3.5}N_{t}^{6}+K_{u}^{3.5}))$ with $C_{5}$ represented as the number of iterations to solve problem (P1).
IV Numerical Results
Figure 2: The simulation setup.
In this section, we present numerical results to validate our proposals. The experimental setup is illustrated in Fig. 2, featuring multiple TBS and RBS units jointly attempting to detect a target amidst clutters while simultaneously respectively serving three DL and three UL users. In the experiment, TBSs, RBSs, DL users, UL users, one target, and clutters are randomly located within a circular area with a radius of $70$ meters (m). We employ a geometric channel model [5] for the link concerning each user, wherein the number of transmit and receive channel paths is consistently identical, i.e., $L^{t,0}_{m,k}=L^{r,0}_{m,k}=L^{t,1}_{p,l}=L^{r,1}_{p,l}=L^{t,2}_{k,l}=L^{r,2}_{k,l}\triangleq L=6,\forall m\in\mathcal{M}_{t},\forall p\in\mathcal{M}_{r},\forall k\in\mathcal{K}_{d},\forall l\in\mathcal{K}_{u}$.Thus, the path-response matrix for each user is diagonal, i.e., the path-response matrix for between the $m$-th TBS and the $k$-th DL user is represented as $\mathbf{\Sigma}^{0}_{m,k}=\textrm{diag}\{\sigma^{0}_{m,k,1},\dots,\sigma^{0}_{m,k,L}\}$, where each $\mathbf{\sigma}^{0}_{m,k,l}$ satisfying $\sigma^{0}_{m,k,l}\sim\mathcal{CN}(0,{(c^{0}_{m,k})^{2}}/{L}),l=1,\dots,L$ [5], and $(c^{0}_{m,k})^{2}=C_{0}d_{m,k}^{-\alpha_{loss}}$, where $C_{0}$ corresponds to the path loss at the reference distance of $1$ m, $d_{m,k}$ is the propagation distance between the $m$-th TBS and the $k$-th DL user, $\alpha_{loss}=2.8$ is the path-loss exponent. The elevation and azimuth of both AoAs and AoDs are assumed to be independent and identically distributed variables following the uniform distribution over $\frac{\pi}{2},\frac{\pi}{2}]$. The TBS-target/clutters and RBS-target/clutters are modeled as line-of-sight (LoS) channels. The moving regions for MAs at both the BS and users are set as square areas of size $\frac{A}{2},\frac{A}{2}]\times[-\frac{A}{2},\frac{A}{2}]$. In addition, $M_{t}=2$ TBS equip $N_{t}=4$ MAs and $M_{r}=2$ RBS equip $N_{r}=4$ MAs. The transmit power of each TBS is set as 3030dBm. The noise power and the predefined target detection level of RBS are set as $\sigma^{2}_{d,k}=\sigma^{2}_{r}=-80$dBm and $\Gamma_{r}=3$dB, respectively.
Figure 3: Convergence of Alg. 1.
Fig. 3 examines the convergence behavior of the overall solution in Alg. 1. The obtained sum-rate iterates under varying configurations of $N_{t}$ and $N_{r}$ are illustrated in the plot. As shown in Fig. 3, the proposed algorithm can yield monotonic improvement in sum-rate and generally achieves significant beamforming gain within the first 10 iterations.
Figure 4: The impact of TBS transmit power.
Alg. 1 is labeled by “Joint MA”. For comparison in sum-rate enhancement, we consider the following four foundational scenarios: 1) BS MA: All BSs are equipped with the MAs, while all users respectively employ an MA with fixed; 2) User MA: All users respectively employ an MA, while the MAs of all BSs are fixed; 3) Rand. MA: The MAs of all BSs and users have been fixed; 4) FPA: Both BSs and users are equipped with the FPA-based array.
Fig. 4 compares the sum-rate for different schemes versus the transmit power of all TBS. The sum-rate of all schemes increases significantly with the transmit power and the three MA-based schemes, i.e., “Joint MA”, “BS MA”, and “User MA” cases, significantly improve the sum-rate compared to both “FPA” and “Rand. MA” cases. For the “Joint MA” case, when the transmit power increases from 10 dBm to 35 dBm, the sum-rate increases by about 67%. In particular, the “Joint MA” case shows a notable advantage over all other cases regarding sum-rate. Specifically, under the 20 dBm case, the “Joint MA” case outperforms “FPA” and “Rand. MA” cases by approximately 30% and 35%, respectively. This is anticipated, as the proposed “Joint MA” scheme exploits the most spatial DoFs, and the other four schemes face performance loss due to less flexibility in channel reconfiguration.
Figure 5: The impact of TBS antenna number.
Figure 6: The impact of RBS antenna number.
Fig. 5 illustrates the impact of the number of antenna at the TBSs. The upper and lower subplots correspond to different numbers of TBSs, respectively. In our test, the number of antenna $N_{t}$ varies from 2 to 9. As reflected by Fig. 5, increasing $N_{t}$ can improve beamforming gain for all schemes. Moreover, the “Joint MA” scheme boosts the system’s sum-rate more significantly compared to all other schemes. For the “Joint MA” case, when the transmit antenna number increases from 2 to 9, the sum-rate improves by approximately 25% for the $M_{t}=2$” scenario and by about 33% for the $M_{t}=4$” scenario.
Fig. 6 shows the impact of the number of antenna at the RBSs. As can be seen in both upper and lower subplots, which correspond to varying quantities of RBSs, sum-rate increases when $N_{r}$ increases. Compared to other benchmarks, the deployment of MA in both BS and user can significantly boost the sum-rate.
Figure 7: The impact of the number of paths between TBS and DL user.
Figure 8: The impact of the number of paths between RBS and UL user.
Fig. 7 explores the influence of the number of paths between TBSs and DL users on the sum-rate. It can be easily observed that a larger number of paths results in a higher achievable sum-rate across all schemes, and our proposed algorithm demonstrates a superior performance over other schemes. Besides, when $K_{d}=6$, all schemes exhibit a higher sum-rate compared to the case when $K_{d}=3$. Fig. 8 compares the sum-rate for different schemes versus the number of paths between RBSs and UL users. It is readily apparent that an increased number of paths leads to a higher achievable sum-rate in all schemes, with the “Joint MA” scheme exhibiting superior performance compared to others. In addition, the $K_{u}=6$ case can achieve better performance than the $K_{u}=3$ case.
Figure 9: The impact of the sensing SINR performance.
The impact of varying sensing SINR requirements is illustrated in Fig. 9. A clear trade-off between communication performance and target sensing can be observed. In this experiment, the sensing SINR ranges from 0 dB to $5$ dB. As shown in the figure, increasing the sensing SINR results in a decrease in the sum-rate for all cases.
V Conclusions
In this paper, we considered a novel MA-aided networked ISAC system that accomplishes radar sensing as well as DL and UL communication capabilities concurrently. Aiming to maximize the sum-rate among all DL and UL users, we investigated a joint active beamforming and antenna position coefficients design problem. Despite the non-convex nature of the considered problem with highly coupled optimization variables, we propose an iterative algorithm to jointly optimize BS beamforming, UL users’ power allocation, receiving processors and MAs’ position configuration for enabling radar sensing and communication functionalities. Numerical results showcase the effectiveness of our proposed solution, highlight the advantages of deploying MA in the networked ISAC system, and offer useful engineering insights. Firstly, while ensuring the sensing performance, MA can enhance communication performance compared to the case with the FPA-based array. Additionally, in comparison to benchmark schemes, our proposed algorithm significantly reduces the required transmit power or the number of antennas needed to achieve the desired level of communication performance.